### v4

# 2cm2cm1cm1cm

Abstract
Quadratic Hamiltonians are important in quantum field theory and quantum statistical mechanics. Their general studies, which go back to the sixties, are relatively incomplete for the fermionic case studied here. Following Berezin, they are quadratic in the fermionic field and in this way well-defined self-adjoint operators acting on the fermionic Fock space. We analyze their diagonalization by applying a novel elliptic operator-valued differential equations on the one-particle Hilbert space studied in a companion paper. This allows for their ($\mathrm{N}$–) diagonalization under much weaker assumptions than before. Last but not least, in 1994 Bach, Lieb and Solovej defined them to be generators of strongly continuous unitary groups of Bogoliubov transformations. This is shown to be an equivalent definition, as soon as the vacuum state belongs to the domain of definition of these Hamiltonians. This second outcome is demonstrated to be reminiscent to the celebrated Shale-Stinespring condition on Bogoliubov transformations.
Keywords: quadratic operators, Bogoliubov transformation, Shale-Stinespring condition, elliptic flow, Brocket-Wegner flow.
2020 AMS Subject Classification: 81Q10, 34A34, 47D06
1 Introduction
Quadratic Hamiltonians have been used for decades in theoretical physics, starting with Bogoliubov’s theory of superfluidity [1] and the BCS theory of superconductivity [2, 3, 4, 5] for bosonic and fermionic systems, respectively. They can formally be defined as formal series:
$\mathrm{H}_{0}\doteq\sum_{k,l\in\mathbb{N}}\{\Upsilon_{0}\}_{k,l}a_{k}^{\ast}a_{l}+\{D_{0}\}_{k,l}a_{k}^{\ast}a_{l}^{\ast}+\epsilon\overline{\{D_{0}\}}_{k,l}a_{k}a_{l}+E_{0}\mathbf{1\ ,}$ (1)
[…]
In this paper, we only study the fermionic case. Indeed, after 1967, to our knowledge, there is surprisingly no new general mathematical result on fermionic quadratic Hamiltonians defined by (1). By contrast, driven by the mathematical justification of Bogoliubov’s theory or the Brockett-Wegner flow, new studies on the bosonic case have been developed in recent years, see [6, 8, 9, 7], even if bosonic quadratic Hamiltonians were ditched also for decades, from the sixties [10, 11, 12] until the year 2007 with the paper [7].
Making sense of the formal definition (1) as a self-adjoint operator acting on the fermionic Fock space $\mathcal{F}$ over an infinite dimensional Hilbert space $\mathfrak{h}$ can be a non-trivial task, in general. This issue is already studied in [11, Theorem 6.1], but the corresponding proof is not completely rigorous. Another proof is given by [13, Proposition 2.1], but under the assumption of bounded operators $\Upsilon_{0}$, which is a too strong condition in practice. In Proposition 3.4, we show that $\mathrm{H}_{0}$ can be well-defined as a self-adjoint operator on $\mathcal{F}$ even for unbounded self-adjoint operators $\Upsilon_{0}$, provided $D_{0}$ is always a Hilbert-Schmidt operator. This necessary outcome uses (a priori) nontrivial arguments, which are, however, relatively direct to obtain by translating the results of [7, Theorem 5.3] (itself inspired by [11, Theorem 6.1], following [14]) on bosonic quadratic operators to the fermionic case.
The definition of fermionic quadratic operators via formal series like (1) refers to Berezin’s approach [11]. An alternative method to define fermionic quadratic operators has been used by Bach, Lieb and Solovej in [15]. Instead of (1), they are defined to be the generators of strongly continuous unitary groups of Bogoliubov transformations. In this definition, quadratic operators are automatically self-adjoint, thanks to Stone’s theorem. By contrast, an explicit form of such quadratic Hamiltonians, as given by (1), becomes an open problem. In fact, Bach, Lieb and Solovej’s viewpoint is very general and it is even not clear whether the (possibly unbounded) generator $h=h^{\ast}$ of the unitary group defining the Bogoliubov transformation in their definition can be written as a block operator matrix in infinite dimension, allowing to speak about $\Upsilon_{0}$ and $D_{0}$. This is a necessary assumption one has to use. Up to this condition, we show in Theorem 2.15 as well as in Corollaries 2.14 and 2.17 that Berezin’s approach and Bach, Lieb and Solovej’s viewpoint are equivalent to each other, as soon as the vacuum state belongs to the domain of definition of quadratic Hamiltonians. It corresponds to take a Hilbert-Schmidt operator $D_{0}$ in (1), which is a condition already used to prove the self-adjointness of operators defined by (1). This is one of the two main results of our paper.
By Theorems 2.15 and 2.18 as well as Corollary 2.17, the Hilbert-Schmidt condition on $D_{0}$ can be seen as a Shale-Stinespring condition for generators of strongly continuous unitary groups of Bogoliubov transformations. Recall that the celebrated Shale-Stinespring condition is a sufficient and necessary condition to implement Bogoliubov transformations on the underlying indexing Hilbert space of an abstract CAR $C^{\ast}$–algebras via unitary transformations in the Fock space representation, see [16, Theorems 3.1 and 6.2] or [17, Theorem 7].
Bogoliubov transformations are not only useful to define quadratic Hamiltonians within Bach, Lieb and Solovej’s approach. By definition, they transform creation and annihilation operators in other creation and annihilation operators, leading in theoretical physics to quasi-particle interpretations. Implementable Bogoliubov transformations on Fock spaces are well-known in relatively simple cases to ($\mathrm{N}$–) diagonalize quadratic Hamiltonians $\mathrm{H}_{0}$ of the form (1). The $\mathrm{N}$–diagonalization of $\mathrm{H}_{0}$ refers to the construction of a unitary operator $\mathrm{U}$ acting on the Fock space such that $\mathrm{UH}_{0}\mathrm{U}^{\ast}$ commutes with the particle number operator $\mathrm{N}$. It means in this case that $\mathrm{UH}_{0}\mathrm{U}^{\ast}$ is a quadratic Hamiltonian of the form (1) with $D_{0}=0$, i.e., the second quantization of a self-adjoint operator acting on the one-particle Hilbert space (see (50)).
Mathematical results on the $\mathrm{N}$–diagonalization of general quadratic operators like (1) have been obtained in [10, 11, 12], but since the sixties no new mathematical result has been performed. Note that Araki presented in 1968 [18] a very general method for the “$\mathrm{N}$–diagonalization” of bilinear Hamiltonians, but it does not solve the issue addressed here. For more details, see Section 2.4.
The second main result of this paper concerns the $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians under much more general conditions than before. Compare indeed the previously known result given below by Theorem 2.3, proven in [11, 12], with Theorem 2.4. For instance, observe that the gap condition $\Upsilon_{0}\geq\alpha\mathbf{1}$ with $\alpha\in\mathbb{R}^{+}$ and the Hilbert-Schmidt condition on the commutator $[\Upsilon_{0},D_{0}]$ in Theorem 2.3 are clearly very strong restrictions as compared to the hypotheses of Theorem 2.4, which can be applied to lower semibounded $\Upsilon_{0}$. In some cases, we are even able to fully characterize the spectrum of fermionic quadratic Hamiltonians. Note that all these results are done by mean of Bogoliubov transformations, implemented in the Fock space, as expected.
The $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians formally defined by (1) may be achieved with presumably optimal conditions in a similar way to what it is efficiently done in [8] for the bosonic case. Here, we follow another path by using the Brockett-Wegner flow [19, 20, 21, 22] like in [6], because this method of proof is interesting in its own right, while leading excellent results. Indeed, such differential equations on spaces of operators are very little developed in Mathematics, being in general very challenging. For instance, although the Brockett-Wegner flow has been often used in theoretical physics [23], there are only two rigorous results on the well-posedness of the Brockett-Wegner flow and only one study of its asymptotics for unbounded operators [6, 21]. By using fermionic quadratic operators we give another example of its well-posedness and asymptotics, by Theorem 3.8, Proposition 3.9 and Corollary 3.12. In fact, the Brockett-Wegner flow [19, 20, 21, 22] is used as a guideline in order to connect the diagonalization of quadratic Hamiltonians to a new, elegant flow on operators acting on the one-particle Hilbert space. In [24] we demonstrate that this non-linear flow presents remarkable ellipticity properties that turn out to be crucial for the study of the infinite-time limit of its solution, which is proven under relatively weak hypotheses on the initial data. It is the elliptic counterpart of an hyperbolic flow used to ($\mathrm{N}$–) diagonalize bosonic quadratic Hamiltonians in [6]. Essential results of [24] for the diagonalization of quadratic Hamiltonians are shortly explained in Section A, while the present paper gives a salient application of this elliptic flow to quantum field theory and quantum statistical mechanics.
To conclude, our main results are the following: (a) Theorem 2.15, and Corollaries 2.14 and 2.17; (b) Theorem 2.4. Observe that (a) and related results belong to the mathematical foundations of fermionic quadratic Hamiltonians, similar to Bruneau and Derezinski’s study [7, 9] on bosonic quadratic Hamiltonians. (b) focuses on their $\mathrm{N}$–diagonalization, similar to the studies [6, 8].
The paper is organized as follows: Section 2 explains fermionic quadratic Hamiltonians, by including our main results while discussing the method of proof via the Brockett-Wegner flow. Most of the technical results as well as a paradigmatic (and historical) example of fermionic quadratic Hamiltonian are gathered in Section 3. In an attempt to be as self-contained and pedagogical as possible, it also contains a review of results that are of general utility in this paper. Section A is an appendix on the elliptical flow studied in [24], which gives the pivotal results used here to diagonalize quadratic Hamiltonians.
2 Fermionic Quadratic Hamiltonians
2.1 Preliminary Definitions
To fix notation, let $\mathfrak{h}\doteq L^{2}(\mathcal{M})$ be a separable complex Hilbert space which we assume to be realized as a space of square-integrable (complex-valued) functions on a measure space $(\mathcal{M},\mathfrak{a})$. The scalar product and norm on $\mathfrak{h}$ are given by
$\left\langle f,g\right\rangle_{\mathfrak{h}}\doteq\int_{\mathcal{M}}\overline{f\left(x\right)}g\left(x\right)\mathrm{d}\mathfrak{a}\left(x\right)\qquad\text{and}\qquad\left\|f\right\|_{\mathfrak{h}}\doteq\langle f,f\rangle_{\mathfrak{h}}^{1/2}$ (2)
for any $f,g\in\mathfrak{h}$. For $f\in\mathfrak{h}$, we define its complex conjugate $\bar{f}\in\mathfrak{h}$ by $\bar{f}\left(x\right)\doteq\overline{f\left(x\right)}$, for all $x\in\mathcal{M}$.
For any bounded (linear) operator $X$ on $\mathfrak{h}$, using the Riesz lemma [25, Theorem II.4] we define its transpose $X^{\top}$ and its complex conjugate $\bar{X}$ by their sesquilinear form $\langle f,X^{\top}g\rangle_{\mathfrak{h}}\doteq\langle\bar{g},X\bar{f}\rangle_{\mathfrak{h}}$ and $\langle f,\bar{X}g\rangle_{\mathfrak{h}}\doteq\overline{\langle\bar{f},X\bar{g}\rangle_{\mathfrak{h}}}$ for $f,g\in\mathfrak{h}$. These definitions can be extended to unbounded operators $X$ with domain $\mathcal{D}(X)\subseteq\mathfrak{h}$. Note that $X\geq 0$ iff $X^{\top}\geq 0$ and the adjoint of an operator $X$ equals $X^{\ast}=\overline{X^{\top}}=\overline{X}^{\top}$, where it exists.
Remark 2.1
The assumption $\mathfrak{h}\doteq L^{2}(\mathcal{M})$ slightly simplifies the setting. More generally, one can take any separable complex Hilbert space $\mathfrak{h}$ endowed with a complex conjugation $\mathcal{C}$, which is an antiunitary involution on $\mathfrak{h}$, i.e., an antilinear mapping $\mathfrak{h}\rightarrow\mathfrak{h}$ such that $\mathcal{C}^{2}=\mathbf{1}$ and
$\left\langle\mathcal{C}f,\mathcal{C}g\right\rangle_{\mathfrak{h}}=\left\langle g,f\right\rangle_{\mathfrak{h}}\ ,\qquad f,g\in\mathfrak{h}\ .$
See, e.g., [26, Chapter 7]. Then, for any operator $X$ on $\mathfrak{h}$, $X^{\top}\doteq\mathcal{C}X^{\ast}\mathcal{C}$ and $\overline{X}\doteq\mathcal{C}X\mathcal{C}$.
The Banach space of bounded operators acting on a Banach space $\left(\mathcal{X},\left\|\cdot\right\|_{\mathcal{X}}\right)$ is denoted by $\mathcal{B}(\mathcal{X})$, with operator norm
$\left\|X\right\|_{\mathrm{op}}\doteq\sup_{f\in\mathcal{X}:\left\|f\right\|_{\mathcal{X}}=1}\left\|Xf\right\|_{\mathcal{X}}\ ,\qquad X\in\mathcal{B}\left(\mathcal{X}\right)\ .$
The unit $\mathbf{1}\in\mathcal{B}(\mathcal{X})$ is the identity operator. We also use the notation $\mathcal{L}^{2}(\mathfrak{h})$ for the Hilbert space of Hilbert-Schmidt operators acting on the separable Hilbert space $\mathfrak{h}$. The norm in $\mathcal{L}^{2}(\mathfrak{h})$ is denoted by
$\left\|X\right\|_{2}\doteq\sqrt{\mathrm{tr}\left(X^{\ast}X\right)}\ ,\qquad X\in\mathcal{L}^{2}(\mathfrak{h})\ ,$
where $\mathrm{tr}(\cdot)$ denotes the usual trace for operators.
2.2 Fermionic Quadratic Hamiltonians
Let $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ be any Hilbert-Schmidt operator satisfying $D_{0}=-D_{0}^{\top}$. Take $E_{0}\in\mathbb{R}$ and a self-adjoint operator $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ that is bounded from below on the (separable) Hilbert space $\mathfrak{h}$. Pick some real [^1] Real orthonormal bases are only taken for convenience, in order to slightly simplify our computations, but it is not a necessary condition. Note additionally that such a real basis exists, thanks to Zorn’s lemma. See [26, Lemma 4.201]. orthonormal basis $\left\{\varphi_{k}\right\}_{k=1}^{\infty}$ in the dense domain $\mathcal{D}\left(\Upsilon_{0}\right)\subseteq\mathfrak{h}$ of $\Upsilon_{0}$ and, for any $k\in\mathbb{N}$, let $a_{k}\doteq a\left(\varphi_{k}\right)$ be the corresponding annihilation operator acting on the fermionic Fock space
$\mathcal{F}\doteq\bigoplus_{n=0}^{\infty}\wedge^{n}\mathfrak{h}\ ,$ (3)
the scalar product and the norm of which are respectively denoted by $\left\langle\cdot,\cdot\right\rangle_{\mathcal{F}}$ and $\left\|\cdot\right\|_{\mathcal{F}}$. Note that this scalar product is the sum over $n\in\mathbb{N}$ of each canonical scalar product on the sector $\wedge^{n}\mathfrak{h}$. Here, $\wedge^{0}\mathfrak{h}\doteq\mathbb{C}$, while, for $n\in\mathbb{N}$, $\wedge^{n}\mathfrak{h}$ is the subspace of totally antisymmetric $n$–particle wave functions in $\mathfrak{h}^{\otimes n}$, the $n$–fold tensor product of $\mathfrak{h}$.
For any fixed $E_{0}\in\mathbb{R}$, the fermionic quadratic operator is defined through the operators $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ and $D_{0}=-D_{0}^{\top}$ by
$\mathrm{H}_{0}\doteq\sum_{k,l\in\mathbb{N}}\{\Upsilon_{0}\}_{k,l}a_{k}^{\ast}a_{l}+\{D_{0}\}_{k,l}a_{k}^{\ast}a_{l}^{\ast}+\{\bar{D}_{0}\}_{k,l}a_{l}a_{k}+E_{0}\mathbf{1}$ (4)
with
$\left\{X\right\}_{k,l}\doteq\left\langle\varphi_{k},X\varphi_{l}\right\rangle\ ,\qquad k,l\in\mathbb{N}\ ,$
for all operators $X$ acting on $\mathfrak{h}$. Since $\Upsilon_{0}$ is by assumption self-adjoint and $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ is Hilbert-Schmidt, we prove in Proposition 3.4 that the operator $\mathrm{H}_{0}$ is essentially self-adjoint on the domain
$\mathcal{D}_{0}\doteq\bigcup_{N\in\mathbb{N}}\left(\bigoplus_{n=0}^{N}\left(\wedge^{n}\mathcal{D}\left(\Upsilon_{0}\right)\right)\right)$ (5)
with $\wedge^{0}\mathcal{D}\left(\Upsilon_{0}\right)\doteq\mathbb{C}$ while for $n\in\mathbb{N}$, $\wedge^{n}\mathcal{D}\left(\Upsilon_{0}\right)$ is the subspace of totally antisymmetric $n$–particle wave functions in $\mathcal{D}\left(\Upsilon_{0}\right)^{\otimes n}$, the $n$–fold tensor product of $\mathcal{D}\left(\Upsilon_{0}\right)^{\otimes n}\subseteq\mathfrak{h}^{\otimes n}$. We use again the notation $\mathrm{H}_{0}\equiv\mathrm{H}_{0}^{\ast\ast}$ for its (uniquely defined) self-adjoint extension. Note additionally that $\mathrm{H}_{0}$ is independent of the (possibily real) orthonormal basis $\left\{\varphi_{k}\right\}_{k=1}^{\infty}\subseteq\mathcal{D}\left(\Upsilon_{0}\right)\subseteq\mathfrak{h}$ chosen, as explained after Proposition 3.4. For a general form of the quadratic Hamiltonian $\mathrm{H}_{0}$ written within an arbitrary orthonormal basis in $\mathcal{D}\left(\Upsilon_{0}\right)$, see Equation (56). Note also that a continuous integral $\int\mathrm{d}k\ \mathrm{d}\ell$ could also replace the discrete sum $\sum_{k,\ell}$ in (4), with appropriate modifications, since all results also hold in the continuous case.
Remark 2.2 (Alternative definition)
In Definition 2.10, quadratic Hamiltonians are alternatively defined as generators of strongly continuous (one-parameter) unitary groups of Bogoliubov transformations. In this approach, they are automatically self-adjoint, thanks to Stone’s theorem [44, Theorem 6.2], but an explicit form like (4) becomes an open problem. This is studied in Section 2.6.
Observe that the condition $D_{0}=-D_{0}^{\top}$ in (4) can be taken without loss of generality because the family $\{a_{k}\}_{k\in\mathbb{N}}$ satisfies the Canonical Anticommutation Relations (CAR), i.e.,
$a_{k}a_{l}+a_{l}a_{k}=0\ ,\quad a_{k}a_{l}^{\ast}+a_{l}^{\ast}a_{k}=\delta_{k,l}\mathbf{1}\ ,$ (6)
for all $k,l\in\mathbb{N}$. See [27, p. 10]. Meanwhile, the Hilbert-Schmidt condition $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ is a pivotal assumption to obtain a well-defined quadratic Hamiltonian. It is in particular a necessary condition to be able to define $\mathrm{H}_{0}$ on $\mathcal{D}_{0}$ since, using the so-called vacuum state $\Psi\doteq\left(1,0,\ldots\right)\in\mathcal{D}_{0}$, one gets
$\left\|\left(\mathrm{H}_{0}-E_{0}\mathbf{1}\right)\Psi\right\|_{\mathcal{F}}=\sqrt{2}\left\|D_{0}\right\|_{2}\ ,$ (7)
thanks to Equation (52).
Hamiltonians like (4) appear very often in theoretical physics and can for instance encode in some sense interaction effects. In fact, interparticle interactions usually play a pivotal role in physical phenomena, but they make the physical system extremely difficult to study in a mathematically rigorous way and effective models coming from different approximations and ansäzte are usually used. It is exactly what happens for the celebrated BCS theory proposed in the late 1950s to explain conventional superconductivity, which leads to the explicit diagonalization of a (very simple) quadratic operators like (4). See Section 3.7 for more details. Another example is very recently given in the context of the (generalized) Hartree-Fock theory [15], which has a strong affinity with (fermionic) mean-field models and effective quadratic Hamiltonians, as first discussed in [26, Sections 5.7 and 6.10] (with a research paper in preparation).
Fermionic quadratic Hamiltonians, as defined by (4)–(5), are generally unbounded in spite of the boundedness of fermionic creation and annihilation operators [27, Proposition 5.2.2]. For instance, if $\Upsilon_{0}=\mathbf{1}$ is the identity mapping on $\mathfrak{h}$ and $D_{0}=0$, then one obtains from (4) the so-called particle number operator
$\mathrm{N}=\sum_{k\in\mathbb{N}}a_{k}^{\ast}a_{k}\ ,$ (8)
which is unbounded for infinite-dimensional Hilbert spaces $\mathfrak{h}$. It is a very simple (well-known) quadratic Hamiltonian that can be explicitly defined by its spectral properties: its spectrum is $\sigma(\mathrm{N})=\mathbb{N}_{0}$ with
$\mathrm{N}\varphi_{n}\doteq n\varphi_{n}\ ,\qquad n\in\mathbb{N}_{0},\ \varphi_{n}\in\wedge^{n}\mathfrak{h}\subseteq\mathcal{F}\ ,$ (9)
while its domain is
$\mathcal{D}\left(\mathrm{N}\right)\doteq\left\{\left(\varphi_{n}\right)_{n\in\mathbb{N}_{0}}\in\mathcal{F}:\sum_{n\in\mathbb{N}_{0}}n^{2}\left\|\varphi_{n}\right\|^{2}<\infty\right\}\supseteq\mathcal{D}_{0}$ (10)
See, e.g., [27, p. 7]. Equality (8) means in particular that $\mathcal{D}_{0}$ is a core of $\mathrm{N}$.
More generally, quadratic Hamiltonians of the form (4) for $D_{0}=0$ are much easier Hamiltonians to study than general quadratic Hamiltonians with possibly non-zero $D_{0}\neq 0$, because their spectral features can be deduced from the spectral properties of the operator $\Upsilon_{0}$ acting on the one-particle Hilbert space $\mathfrak{h}$. As is well-known, for $D_{0}=0$, the Hamiltonian $\mathrm{H}_{0}$ is the so-called second-quantization of the one-particle Hamiltonian $\Upsilon_{0}$, see, e.g., [27, Section 5.2.1]. Hamiltonians of this form are named $\mathrm{N}$–diagonal since they are the only ones that commute with the particle number operator $\mathrm{N}$.
In the more general cases where $D_{0}\neq 0$, one can usually find a unitary transformation $\mathrm{U}$ such that $\mathrm{U\mathrm{H}}_{0}\mathrm{U}^{\ast}$ is $\mathrm{N}$–diagonal. This was first shown for fermions in the particular case of the BCS theory (Section 3.7) by implementing [^2] Bogoliubov transformations on CAR $C^{\ast}$-algebras are not necessarily implementable on the Fock space, i.e., they do not necessarily well-define a unitary transformation on the Fock space. This is discussed in detail in Section 2.4, see in particular the Shale-Stinespring theorem (Theorem 2.8). on the Fock space an algebraic transformation, named the Bogoliubov (unitary) transformation, which defines a $\ast$-isomorphism of CAR $C^{\ast}$-algebras.
2.3 Diagonalization of Quadratic Hamiltonians
General studies of fermionic quadratic Hamiltonians go back to the sixties with Berezin’s book [11] published in 1966, following Friedrichs’s work [10] written in 1953. The main result on the $\mathrm{N}$–diagonalization of quadratic operators like (4) refers to [11, Theorem 8.2], which we reproduce for the reader’s convenience:
Theorem 2.3 (Berezin)
Let $E_{0}\in\mathbb{R}$. Take $D_{0}=\bar{D}_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ and $\Upsilon_{0}=\Upsilon_{0}^{\top}=\Upsilon_{0}^{\ast}\geq\alpha\mathbf{1}$ with strictly positive $\alpha\in\mathbb{R}^{+}$. Assume additionally that the commutator $[D_{0},\Upsilon_{0}]\in\mathcal{L}^{2}(\mathfrak{h})$ and
$\Upsilon_{0}^{2}+4D_{0}D_{0}^{\ast}\pm 2[D_{0},\Upsilon_{0}]>\alpha\mathbf{1}\ .$ (11)
Then, there is a unitary transformation $\mathrm{U}$ such that $\mathrm{U\mathrm{H}}_{0}\mathrm{U}^{\ast}$ is $\mathrm{N}$–diagonal.
Note that this theorem is originally stated with the condition
$\Upsilon_{0}^{2}+4D_{0}D_{0}^{\ast}+2[D_{0},\Upsilon_{0}]>\alpha\mathbf{1}$ (12)
instead of (11). But it seems to be a mistake, see [12, Theorem 3]. The unitary transformation of Theorem 2.3 is given via a Bogoliubov transformation.
Note that Araki presented in 1968 [18] a very general method for the “$\mathrm{N}$–diagonalization” of bilinear Hamiltonians. The results are done in great generality by using an algebraic approach, but they do not solve the issue addressed here. For more details, we recommend Section 2.4, where Araki’s work is described in detail.
In fact, after 1967, to our knowledge, there is surprisingly no new general results on the $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians described here. The same situation appears for the bosonic case for which no new results on such Hamiltonians were performed since the 1960ies [10, 11, 12], until the year 2007 with the paper [7]. Then, driven by the mathematical justification of Bogoliubov’s theory or the Brockett-Wegner flow, new results on such models have been developed in recent years, see [6, 8, 9], with for instance the efficient proof on the $\mathrm{N}$–diagonalization of bosonic quadratic Hamiltonians under presumably optimal conditions [8].
By contrast, there is no recent activity on the (equally important) fermionic case and we address the issue of the $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians by means of an elliptic operator-valued flow (Section A). It leads to a continuous family of Bogoliubov transformations and gives new properties on the $\mathrm{N}$–diagonal form of $\mathrm{\mathrm{H}}_{0}$, allowing in some case to compute it explicitly. This is explained in Section 2.5 while here we only sum up the final result on the $\mathrm{N}$–diagonalization of quadratic Hamiltonians:
Theorem 2.4 (Diagonalization of quadratic Hamiltonians)
Let $E_{0}\in\mathbb{R}$. Take $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ and $\Upsilon_{0}=\Upsilon_{0}^{\ast}$, both acting on $\mathfrak{h}$. Assume that
$\Upsilon_{0}\geq-\left(\mu-\varepsilon\right)\mathbf{1}\qquad\text{and}\qquad\Upsilon_{0}+4D_{0}\left(\Upsilon_{0}^{\top}+\mu\mathbf{1}\right)^{-1}D_{0}^{\ast}\geq\mu\mathbf{1}$ (13)
for some $\0\mu\in\mathbb{R}\backslash\{0\}$ and $\varepsilon\in\mathbb{R}^{+}$. Then, there is a unitary transformation $\mathrm{U}$ such that
$\mathrm{U\mathrm{H}}_{0}\mathrm{U}^{\ast}=\sum_{k,l\in\mathbb{N}}\{\Upsilon_{\infty}\}_{k,l}a_{k}^{\ast}a_{l}+\left(E_{0}-8\int_{0}^{\infty}\left\|D_{\tau}\right\|_{2}^{2}\mathrm{d}\tau\right)\mathbf{1}$
with the operator family $D$ and
$\Upsilon_{\infty}\doteq\Upsilon_{0}+\Delta_{\infty}\geq\left|\mu\right|\mathbf{1}$
being defined via Theorems A.1 and A.2. In particular, $\mathrm{U\mathrm{H}}_{0}\mathrm{U}^{\ast}$ is $\mathrm{N}$–diagonal.
Proof. Combine Corollary 3.12 with Theorem A.2. In fact, Theorem 2.4 gives a nice application of the elliptic operator-valued flow studied in [24] to quantum field theory and quantum statistical mechanics.
Remark 2.5
Let $\alpha\in\mathbb{R}^{+}$ and assume that $\Upsilon_{0}\geq\alpha\mathbf{1}$. Then, upon choosing $\varepsilon\doteq\alpha$ and $\mu\doteq\alpha/2$, observe that (13) holds true. As a result, Theorem A.2 can be applied to all positive operators $\Upsilon_{0}$ with spectral gap. In this case, note that $\Upsilon_{\infty}\geq\alpha\mathbf{1}$.
Remark 2.6
To get a stronger exponential decay of the Hilbert-Schmidt norm $\left\|D_{t}\right\|_{2}$ and a more accurate estimate of the spectral gap of $\Upsilon_{\infty}$, one shall take the maximum value of $\mu\in\mathbb{R}$ such that (13) holds true (provided such a $\mu$ exists).
By Proposition 3.6 and Corollary 3.11, the unitary transformation $\mathrm{U}$ implements a Bogoliubov transformation. The $\mathrm{N}$–diagonal form of $\mathrm{\mathrm{H}}_{0}$ can be additionally computed in some cases. For instance, using the constant of motion (192) and Theorem A.2 with $D_{0}^{\top}=\pm D_{0}$, we have
$\mathrm{tr}\left(\Upsilon_{\infty}^{2}-\Upsilon_{0}^{2}-4D_{0}D_{0}^{\ast}\right)=0\ .$ (14)
at least if $\Upsilon_{0}$ is bounded. In the case that $\Upsilon_{0}D_{0}=D_{0}\Upsilon_{0}^{\top}$ on the domain $\mathcal{D}(\Upsilon_{0}^{\top})$ together with $\Upsilon_{0}\geq-\left(\mu-\varepsilon\right)\mathbf{1}$ for some $\mu\in\mathbb{R}$ and $\varepsilon\in\mathbb{R}^{+}$ and
$\max\left\{\|(\Upsilon_{0}+\mu\mathbf{1)}D_{0}(\Upsilon_{0}^{\top}+\mu\mathbf{1)}^{-1}\|_{\mathrm{op}},\|(\Upsilon_{0}^{\top}+\mu\mathbf{1)}D_{0}^{\ast}(\Upsilon_{0}+\mu\mathbf{1)}^{-1}\|_{\mathrm{op}}\right\}\leq\mathrm{C}$ (15)
for some strictly positive constant $\mathrm{C}\in\mathbb{R}^{+}$. Then, [24, Equation (11)] yield
$\Upsilon_{\infty}=\sqrt{\Upsilon_{0}^{2}+4D_{0}D_{0}^{\ast}}$ (16)
on the domain $\mathcal{D}(\Upsilon_{0})$ as well as
$8\int_{0}^{\infty}\left\|D_{\tau}\right\|_{2}^{2}\mathrm{d}\tau=\frac{1}{2}\mathrm{tr}\left(\sqrt{\Upsilon_{0}^{2}+4D_{0}D_{0}^{\ast}}-\Upsilon_{0}\right)\in\left[0,\infty\right)\ .$ (17)
Note that the fact that the trace in the above equation is finite is a non-trivial [^3] For a direct proof under the condition $\Upsilon_{0}=\Upsilon_{0}^{\ast}\geq\alpha\mathbf{1}$ with $\alpha\in\mathbb{R}^{+}$, see [11, Lemma 8.1]. consequence of the study. In other words, $\mathrm{U\mathrm{H}}_{0}\mathrm{U}^{\ast}$ can be fully characterized in this “commutative” case. This is for instance used in Section 3.7 to recover the exact $\mathrm{N}$–diagonal form (187) of the BCS Hamiltonian.
Note that Theorem 2.3 does not make explicit neither the $\mathrm{N}$–diagonal form of fermionic quadratic Hamiltonians nor the Bogoliubov transformation. In addition, observe that (13) looks similar to Condition (11) of Theorem 2.3, but the assumptions $\Upsilon_{0}\geq\alpha\mathbf{1}$ with $\alpha\in\mathbb{R}^{+}$ and $[\Upsilon_{0},D_{0}]\in\mathcal{L}^{2}(\mathfrak{h})$ in Theorem 2.3 are clearly strong restrictions as compared to the hypotheses of Theorem 2.4. See for instance Remark 2.5. In fact, the asssumptions of Berezin’s theorem (Theorem 2.3) are so strong that they may not even be satisfied by the elementary quadratic Hamiltonians of the BCS theory, as discussed in Section 3.7. Indeed, for these Hamiltonians, the condition $\Upsilon_{0}\geq\alpha\mathbf{1}$ may not necessarily hold for a strictly positive $\alpha\in\mathbb{R}^{+}$. Meanwhile, the assumptions of Theorem 2.4 are much more natural and general than those of Theorem 2.3 and they include important cases with negative spectra for $\Upsilon_{0}$.
To conclude, note that the conditions of Theorems 2.4 are only sufficient. As already discussed in [24] we expect the limiting case to be when
$\Upsilon_{0}>0\text{\qquad and\qquad}\Upsilon_{0}+4D_{0}\left(\Upsilon_{0}^{\top}\right)^{-1}D_{0}^{\ast}\geq 0$ (18)
with $D_{0}\left(\Upsilon_{0}^{\top}\right)^{-1}D_{0}^{\ast}$ being a well-defined operator acting on $\mathfrak{h}$. Up to singular cases, these assumptions may be necessary for the convergence of the flow because the existence of $\Upsilon_{\infty}$ seems to generally imply its positivity (up to singular cases). See the spectral properties of the operator family (195). They correspond to take $\varepsilon,\mu\rightarrow 0$ in the conditions of Theorem 2.4 with a possible limit operator $\Upsilon_{\infty}\doteq\Upsilon_{0}+\Delta_{\infty}>0$ having a priori no spectral gap. The conditions expressed by (18) look similar to the situation $\Omega_{0}>0$, without spectral gap, studied in the hyperbolic case [6]. By analogy with the hyperbolic case, one may assume that $\left(\Upsilon_{0}\right)^{-1/2}D_{0}$ is a Hilbert-Schmidt operator to ensure the existence of $\Delta_{\infty}$ as a trace-class operator. The arguments to handle the non-spectral gap case in [6] were technically involved and long (with domain issues) and the analogous study in the fermionic case under Condition (18) has probably the same mathematical complexity.
2.4 Araki’s Approach on the Level of $C^{\ast}$-Algebras
In 1968, Araki established a very general method for some “$\mathrm{N}$–diagonalization” of bilinear Hamiltonians. This refers to the well-known paper [18]. Such a diagonalization holds true in great generality. The Hilbert-Schmidt condition $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ and the semiboundedness of $\Upsilon_{0}$ are not even required in [18, Theorem 5.4]. This is because the concepts studied in Araki’s paper [18] are different than those analyzed here, albeit still related. Araki’s terminology and notation used in [18] are almost identical to the ones used in other works like in Berezin’s book [11], but they refer to differents concepts or frameworks. For instance, Araki’s study [18] uses a purely $C^{\ast}$-algebraic formulation, while Berezin and we work on the Fock space representation. See also [30] for more details on Araki’s works in relation with the Fock space representation.
In the sequel we give a brief synthesis of Araki’s results [18], highlighting the similarities and differences with the work done here. Hopefully, this section allows the non-expert reader to have an overview on quadratic/bilinear fermionic models and prevents them from possible misunderstandings. This section also allows us to introduce the celebrated Shale-Stinespring condition and the standard concept of implementable Bogoliubov transformations, which is useful afterwards and, in particular, important to keep in mind in the more general (re)definition of quadratic Hamiltonians in Section 2.6. See in particular Definition 2.10.
The paper [18] has an algebraic approach based on self-dual CAR $C^{\ast}$-algebras defined as follows: Let $\mathcal{\mathcal{H}}$ be a complex Hilbert space with either even or infinite dimension. Let $\mathfrak{A}$ be an antiunitary involution on $\mathcal{\mathcal{H}}$, see Remark 2.1. A self-dual CAR algebra
$\mathcal{A}\equiv(\mathcal{A},+,\cdot,\ast,\|\cdot\|)$ (19)
is the universal [^4] For more details on the universal $C^{\ast}$-algebras of family of polynomial relations, see [26, Section 4.8]. In the case of self-dual CAR algebras, they are uniquely defined up to Bogoliubov $\ast$–automorphisms. See [26, Section 4.8.3]. $C^{\ast}$-algebra generated by a unit $\mathfrak{1}$ and a family $\{\mathrm{B}(\psi)\}_{\psi\in\mathcal{\mathcal{H}}}$ of elements satisfying the following conditions:
  1. (a)
    The mapping $\psi\mapsto\mathrm{B}\left(\psi\right)^{\ast}$ is (complex) linear.
  2. (b)
    $\mathrm{B}(\psi)^{\ast}=\mathrm{B}(\mathfrak{A}(\psi))$ for any $\psi\in\mathcal{H}$.
  3. (c)
    The family $\{\mathrm{B}(\psi)\}_{\psi\in\mathcal{\mathcal{H}}}$ satisfies (a part of the) CAR [^5] The use of “CAR” is a slight misuse of language since $\mathrm{B}(\varphi_{1})\mathrm{B}(\varphi_{2})+\mathrm{B}(\varphi_{2})\mathrm{B}(\varphi_{1})=0$ for $\varphi_{1},\varphi_{2}\in\mathcal{\mathcal{H}}$ does not hold true, in general.: For any $\psi_{1},\psi_{2}\in\mathcal{\mathcal{H}}$,
[…]
Note that the elements $\mathrm{B}(\psi)$, $\psi\in\mathcal{\mathcal{H}}$, are the so-called fields, from which creation and annihilation elements can be defined, see for instance (22) in the Fock space representation. By the CAR (20), the antilinear mapping $\varphi\mapsto\mathrm{B}\left(\varphi\right)$ from $\mathcal{\mathcal{H}}$ to $\mathcal{A}$ is necessarily injective and contractive.
For our setting (Section 2), one takes [^6] Note that it is common to use instead $\mathcal{H}=\mathfrak{h}\oplus\mathfrak{h}^{\ast}$, but we decide to keep Araki’s choice [18, Remark 5.3] for better comparison. Here, $\oplus$ stands, as in all the paper, for the (Hilbert) direct sums of Hilbert (sub)spaces (and not the algebraic direct sum). $\mathcal{H}=\mathfrak{h}\oplus\mathfrak{h}$ and the canonical antiunitary involution of $\mathcal{H}$ is defined from the complex conjugation $\mathcal{C}$ (Remark 2.1) by
$\mathfrak{A}=\left(\begin{array}[]{cc}0&\mathcal{C}\\ \mathcal{C}&0\end{array}\right)\ .$ (21)
By contrast, we use in this paper the Fock space representation of such $C^{\ast}$-algebras, which is the representation [^7] A representation on the Hilbert space $\mathcal{X}$ of a $C^{\ast}$–algebra $\mathcal{A}$ is, by definition [32, Definition 2.3.2], a $\ast$–homomorphism $\mathbf{\pi}$ from $\mathcal{A}$ to the unital $C^{\ast}$–algebra $\mathcal{B}(\mathcal{X})$ of all bounded linear operators acting on $\mathcal{X}$. $\mathbf{\pi}_{\mathcal{F}}$ on the fermionic Fock space $\mathcal{F}$ (3) of the self-dual CAR  algebra $\mathcal{A}$, uniquely defined by
$\mathbf{\pi}_{\mathcal{F}}\left(\mathrm{B}\left(\psi\right)\right)=a(\mathrm{\psi}_{1})+a(\overline{\mathrm{\psi}}_{2})^{\ast}\ ,\qquad\psi=(\mathrm{\psi}_{1},\mathrm{\psi}_{2})\in\mathcal{\mathcal{H}}\ ,$ (22)
$a\left(\varphi\right)$ being the usual annihilation operator associated with $\varphi\in\mathfrak{h}$ and acting on the (fermionic) Fock space $\mathcal{F}$.
Bilinear Hamiltonians used in [18] are analogous to quadratic operators studied here. Nonetheless, they are no longer Hamiltonians, i.e., self-adjoint operators, acting on some Hilbert space, like the Fock space $\mathcal{F}$. In [18] a bilinear Hamiltonian is a generator of a one-parameter automorphism group of the CAR $C^{\ast}$-algebra:
  • A Bogoliubov transformation is a unitary operator $U\in\mathcal{B}(\mathcal{H})$ such that $U\mathfrak{A=A}U$. A self-dual Hamiltonian $h$ on $\mathcal{H}$ is a self-adjoint operator satisfying $h=-\mathfrak{A}h\mathfrak{A}$ with domain $\mathcal{D}(h)=\mathfrak{A}(\mathcal{D}(h))\subseteq\mathcal{\mathcal{H}}$. Basis projections are orthogonal projections $P\in\mathcal{B}(\mathcal{H})$ satisfying $\mathfrak{A}P\mathfrak{A}=P^{\bot}\doteq\mathbf{1}_{\mathcal{H}}-P$.
  • The Bogoliubov $\ast$–automorphism associated with a Bogoliubov transformation $U\in\mathcal{B}(\mathcal{H})$ is the unique $\ast$–automorphism $\mathbf{\chi}_{U}$ of $\mathcal{A}$ satisfying
    $\mathbf{\chi}_{U}\left(\mathrm{B}(\varphi)\right)=\mathrm{B}(U\varphi)\ ,\qquad\varphi\in\mathcal{\mathcal{H}}\ .$ (23)
    See for instance [26, Corollary 4.211]. Given a self-dual Hamiltonian $h$, $(\mathbf{\chi}_{\mathrm{e}^{ith}})_{t\in{\mathbb{R}}}$ is a strongly continuous group of Bogoliubov $\ast$–automorphisms of $\mathcal{A}$.
  • It is known (see, e.g. [31, Lemma 1] when $h\in\mathcal{B}(\mathcal{H})$) that the above group is generated by a symmetric derivation [^8] Symmetric derivations $\delta$ refer to (linear) operators satisfying $\delta(A^{\ast})=\delta(A)^{\ast}$ and $\delta(AB)=\delta(A)B+A\delta(B)$ on $\mathcal{A}$. $\delta_{h}$ acting on the $C^{\ast}$–algebra $\mathcal{A}$, formally denoted by
    $\delta_{h}\left(\cdot\right)\equiv i\left[\mathrm{H},\ \cdot\ \right]\ ,$ (24)
    keeping in mind that [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] usually denotes a commutator. Bilinear Hamiltonians are by definition such a derivation, or such a “$\mathrm{H}$”. The bilinear Hamiltonian “$\mathrm{H}$” is said hermitian or self-adjoint because $h=h^{\ast}$. See [18, Definition 5.1].
  • The diagonalization of bilinear Hamiltonians in Araki’s sense [18, Definition 5.1] then only means the existence of a basis projection $P$ with range $\mathfrak{h}_{P}$, or a Bogoliubov transformation (cf. [18, Lemma 3.6]), such that the self-dual Hamiltonian $h$ has the form
    $\left(\begin{array}[]{cc}h_{1,1}&0\\ 0&h_{2,2}\end{array}\right)$
    on $\mathcal{H}=\mathfrak{h}_{P}\oplus\mathfrak{h}_{P^{\bot}}$, where $h_{1,1},h_{2,2}$ are self-adjoint operators acting on $\mathfrak{h}_{P},\mathfrak{h}_{P^{\bot}}$, respectively. Applying the spectral theorem to the self-dual Hamiltonian $h$, Araki shows in [18, Theorem 5.4] that this is always possible iff the kernel of $h$ has dimension even or infinite, otherwise the same holds true but with a so-called extended Bogoliubov transformation [18, Theorem 5.6].
The motivation of these definitions comes from the fact that such a $\mathrm{H}$ in Araki’s viewpoint can be seen in some very restricted situation as a quadratic Hamiltonian in our sense. See, e.g., [18, Remark 4.3] and [33, Sections 2.2-2.3], where $\mathrm{H}$ is shown to exist as a self-adjoint element of the $C^{\ast}$-algebra $\mathcal{A}$ for trace-class Hamiltonians $h$. However, in general, “$\mathrm{H}$” is only a formal notation, without necessarily another meaning than denoting the symmetric derivation via (24). In other words, [18] analyzes another situation and, in the general case, it is an open problem to show the existence of some well-defined, self-adjoint, quadratic Hamiltonian that leads to $\delta_{h}$ in some representation. In the Fock space representation, one can answer this question provided (4) defines a self-adjoint operator on the Fock space:
Lemma 2.7 (From quadratic Hamiltonians to quasi-free dynamics)
Fix $E_{0}\in\mathbb{R}$. Let $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ and $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ be two operators acting on the Hilbert space $\mathfrak{h}$. Then, for any $t\in{\mathbb{R}}$ and $A\in\mathcal{A}$,
$\mathbf{\pi}_{\mathcal{F}}\circ\mathbf{\chi}_{\mathrm{e}^{ith}}\left(A\right)=\mathrm{e}^{it\mathrm{H}_{0}/2}\mathbf{\pi}_{\mathcal{F}}\left(A\right)\mathrm{e}^{-it\mathrm{H}_{0}/2}\ ,$
where $h$ is the self-dual Hamiltonian defined on $\mathcal{H}=\mathfrak{h}\oplus\mathfrak{h}$ by
$h\doteq\frac{1}{2}\left(\begin{array}[]{cc}\Upsilon_{0}&2D_{0}\\ -2\overline{D}_{0}&-\Upsilon_{0}^{\top}\end{array}\right)=\frac{1}{2}\left(\begin{array}[]{cc}\Upsilon_{0}&2D_{0}\\ 2D_{0}^{\ast}&-\Upsilon_{0}^{\top}\end{array}\right)\ .$ (25)
Proof. Fix all parameters of the lemma. By Proposition 3.4, Equations (4)–(5) define a self-adjoint operator $\mathrm{H}_{0}$ with the dense subset $\mathcal{D}_{0}\subseteq\mathcal{F}$ as a core. In particular, $\mathrm{e}^{it\mathrm{H}_{0}/2}$ is well defined for all times $t\in{\mathbb{R}}$ (via the functional calculus). The corresponding self-dual Hamiltonian $h$ acting on $\mathcal{H}=\mathfrak{h}\oplus\mathfrak{h}$ which would formally lead to our quadratic Hamiltonian (4) (with $E_{0}=0$, the constant being irrelevant here) is equal to (25), see (22) and [18, Equation (5.6)]. Note in this case that $h=h^{\ast}=-\mathfrak{A}h\mathfrak{A}$, thanks to $\Upsilon_{0}=\Upsilon_{0}^{\ast}$, $D_{0}=-D_{0}^{\top}$, $\mathcal{C}^{2}=\mathbf{1}$, $X^{\top}\doteq\mathcal{C}X^{\ast}\mathcal{C}$, $\overline{X}\doteq\mathcal{C}X\mathcal{C}$ (Remark 2.1) and Equation (21). To prove now the assertion, it suffices to show that
$\mathbf{\pi}_{\mathcal{F}}\circ\mathbf{\chi}_{\mathrm{e}^{ith}}\left(\mathrm{B}(\psi)\right)=\mathrm{e}^{it\mathrm{H}_{0}/2}\mathbf{\pi}_{\mathcal{F}}\left(\mathrm{B}\left(\psi\right)\right)\mathrm{e}^{-it\mathrm{H}_{0}/2}$ (26)
for any $\psi\in\mathcal{H}=\mathfrak{h}\oplus\mathfrak{h}$. The proof is relatively standard in its principle. See for instance [33, Lemma 2.8]. We outline it for completeness. As usual, $[A,B]\doteq AB-BA$ is the commutator of two elements $A,B\in\mathcal{A}$. Using the CAR (6) (or (20)), Equation (22) together with the properties of the antiunitary involution $\mathfrak{A}$, $h^{\ast}=-\mathfrak{A}h\mathfrak{A}$ and $\mathrm{B}(\psi)^{\ast}=\mathrm{B}(\mathfrak{A}(\psi))$, we compute that
$\frac{1}{2}\left[\mathrm{H}_{0},\mathbf{\pi}_{\mathcal{F}}\left(\mathrm{B}\left(\psi\right)\right)^{\ast}\right]\varphi=\mathbf{\pi}_{\mathcal{F}}\left(\mathrm{B}\left(h\psi\right)\right)^{\ast}\varphi\ ,$ (27)
for any $\psi\in\mathcal{D}\left(\Upsilon_{0}\right)\oplus\mathcal{D}\left(\Upsilon_{0}^{\top}\right)$ and $\varphi\in\mathcal{D}_{0}$, with $\mathcal{D}\left(\Upsilon_{0}\right),\mathcal{D}\left(\Upsilon_{0}^{\top}\right)\subseteq\mathfrak{h}$ being the dense domains of $\Upsilon_{0}$ and $\Upsilon_{0}^{\top}$, respectively, while $\mathcal{D}_{0}\subseteq\mathcal{F}$ is the dense subset defined by (5). This statement is analogous to [18, Equations (4.15)–(4.16)] done on the level of the algebra $\mathcal{A}$ for trace-class operators $h$ (to obtain a true bilinear Hamiltonian $\mathrm{H}\in\mathcal{A}$). A representation is automatically contractive [32, Proposition 2.3.1] and even isometric when it is injective [32, Proposition 2.3.3]. Meanwhile, the antilinear mapping $\varphi\mapsto\mathrm{B}\left(\varphi\right)$ on $\mathcal{H}$ is also contractive. Together with (23) and (27), it follows that, for any $t\in\mathbb{R}$, $\psi\in\mathcal{D}\left(\Upsilon_{0}\right)\oplus\mathcal{D}\left(\Upsilon_{0}^{\top}\right)$ and $\varphi\in\mathcal{D}_{0}$,
$\partial_{t}\left\{\mathrm{e}^{-it\mathrm{H}_{0}/2}\mathbf{\pi}_{\mathcal{F}}\circ\mathbf{\chi}_{\mathrm{e}^{ith}}\left(\mathrm{B}\left(\psi\right)\right)^{\ast}\mathrm{e}^{it\mathrm{H}_{0}/2}\varphi\right\}=0\ ,$
which in turn implies that
$\mathbf{\pi}_{\mathcal{F}}\circ\mathbf{\chi}_{\mathrm{e}^{ith}}\left(\mathrm{B}(\psi)\right)\varphi=\mathrm{e}^{it\mathrm{H}_{0}/2}\mathbf{\pi}_{\mathcal{F}}\left(\mathrm{B}\left(\psi\right)\right)\mathrm{e}^{-it\mathrm{H}_{0}/2}\varphi$ (28)
for $t\in\mathbb{R}$, $\psi\in\mathcal{D}\left(\Upsilon_{0}\right)\oplus\mathcal{D}\left(\Upsilon_{0}^{\top}\right)$ and $\varphi\in\mathcal{D}_{0}$. Since $\mathcal{D}_{0}$ is dense in $\mathcal{F}$ and the operators in (28) are all bounded, we conclude that
$\mathbf{\pi}_{\mathcal{F}}\circ\mathbf{\chi}_{\mathrm{e}^{ith}}\left(\mathrm{B}(\psi)\right)=\mathrm{e}^{it\mathrm{H}_{0}/2}\mathbf{\pi}_{\mathcal{F}}\left(\mathrm{B}\left(\psi\right)\right)\mathrm{e}^{-it\mathrm{H}_{0}/2}$ (29)
for any $\psi\in\mathcal{D}\left(\Upsilon_{0}\right)\oplus\mathcal{D}\left(\Upsilon_{0}^{\top}\right)$. Since $\mathbf{\pi}_{\mathcal{F}}$ and $\varphi\mapsto\mathrm{B}\left(\varphi\right)$ are contractive mappings, we can extend by continuity (29) to all $\psi\in\mathcal{H}=\mathfrak{h}\oplus\mathfrak{h}$ in order to get (26).
Lemma 2.7 implies that the (necessarily closed) generator of the strongly continuous group $(\mathbf{\chi}_{\mathrm{e}^{ith}})_{t\in{\mathbb{R}}}$ of Bogoliubov $\ast$–automorphisms of $\mathcal{A}$, $h$ being defined by (25), can be seen in the Fock space representation as the commutator $\left[\mathrm{H}_{0},\ \cdot\ \right]$ on some (dense) domain of $\mathcal{B}(\mathcal{F})$, making explicit the relation between Araki’s bilinear Hamiltonians and quadratic operators studied here.
[18, Theorem 5.4] would mean in this case that a quadratic Hamiltonian in the sense of (4) is diagonalizable whenever (25) is essentially self-adjoint and such that the eigenspace associated with the eigenvalue 00 has dimension even or infinite. This would be a very general result, even if $D_{0}$ is required to be Hilbert-Schmidt to ensure the assertions of Lemma 2.7. For instance, the operator $\Upsilon_{0}$ is not even semibounded in this case. However, [18, Definition 5.1], defining the diagonalization of bilinear Hamiltonians, only involves the existence of a Bogoliubov transformation $U\in\mathcal{B}(\mathcal{H})$ such that $U\mathfrak{A=A}U$, which may not define a unitary transformation in the Fock space representation. This is well known via the celebrated Shale-Stinespring condition which refers to the Hilbert-Schmidt assumption.
Theorem 2.8 (Implementation of Bogoliubov transformations)
Let $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ and $\mathfrak{A}$ be defined by (21). Take any Bogoliubov transformation $U$. In particular, $U$ has the form [^9] Use $\overline{X}=\mathcal{C}X\mathcal{C}$ (Remark 2.1) and $U\mathfrak{=A}U\mathfrak{A}$ together with (21). The fact that $UU^{\ast}=U^{\ast}U=1$ also implies other constraints on $\mathbf{u},\mathbf{v}\in\mathcal{B}\left(\mathfrak{h}\right)$, but this is not important here.
$U=\left(\begin{array}[]{cc}\mathbf{u}&\mathbf{v}\\ \overline{\mathbf{v}}&\overline{\mathbf{u}}\end{array}\right)$
[…]
$\mathrm{U}a\left(\varphi\right)\mathrm{U}^{\ast}=a(\mathbf{u}\varphi)+a^{\ast}(\mathbf{v}\overline{\varphi})\ ,\qquad\varphi\in\mathfrak{h}\ .$
See [16, Theorems 3.1 and 6.2] or [17, Theorem 7] for a complete proof. This theorem originally goes back to Shale and Stinespring [34]. For more recent or pedagogical references, see [35, Theorem 9.5] or [15, Equation (2a.9) and Theorem 2.2]. Again the Hilbert-Schmidt condition is referred to as the Shale-Stinespring condition. When it is satisfied, one says that $U$ is (unitarily) implementable. Note that one can extend this result to the case where $\mathrm{U}$ is not necessarily unitary. This refers to [36, Proposition 1.2].
Since Araki’s construction of the Bogoliubov transformation, or the corresponding basis projection, is only based on general considerations applying the spectral theorem to the self-dual Hamiltonian $h$ (see [18, Lemma 3.6 and 4.8]), the main result [18, Theorem 5.4] provides no information on the Shale-Stinespring condition. In other words, we do not know whether the corresponding Bogoliubov transformation given by [18, Theorem 5.4] is implementable or not. This can certainly be done in some general cases, but further studies are in any case required to get a result like Theorem 2.4.
Remark 2.9 (Implementation of non-autonomous flows of Bogoliubov transformations)
In Theorems 3.5 and 3.10 as well as in Proposition 3.6 and Corollary 3.11, we provide a non-autonomous version [^10] By non-autonomous, we refer to the fact that, instead of having a one-parameter group of Bogoliubov transformations, we have a two-parameter family that satisfies the cocycle property. of implemented Bogoliubov transformations, by using the theory of non-autonomous (Kato-hyperbolic) evolution equations. This is used to find a solution to the Brockett-Wegner flow on quadratic Hamiltonians as well as its asymptotics.
2.5 Brockett-Wegner Flow and Elliptic Operator Flows
The $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians may be shown by using an approach that could be similar to [8], done in the bosonic case. However, we follow here another path by using the Brockett-Wegner flow [19, 20, 21, 22] as a guideline. This method of proof is interesting in its own right. Indeed, the Brockett-Wegner flow is very little developed in Mathematics, although it has been often used in theoretical physics [23]. In our setting, it allows one to connect the $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians to an elegant non-linear elliptic flow on operators acting on the one-particle Hilbert space, which is studied in detail in [24] (see Section A).
In this section, we thus formally explain this connection. This allows the interested reader to understand the general strategy to prove Theorem 2.4, before going through the rigorous arguments performed in Section 3. In addition, this demonstrates how the formal Brockett-Wegner flow can produce novel and elegant ordinary differential equations (ODEs) on operator spaces. Remark in this context that differential equations on non-commutative spaces (e.g., operator spaces) are clearly much less known than its commutative counterpart (i.e., ODEs or PDEs on function spaces), which forms a well-established, albeit still very active, research field of Mathematics. This section can thus be viewed as an invitation to develop this research direction on operator spaces.
The Brockett-Wegner flow [19, 20, 21, 22] is a (quadratically) non-linear first-order differential equation for operators. It is used to diagonalize self-adjoint operators. On a formal level, it is easy to describe, but the mathematically rigorous treatment can be rather involved, in particular for unbounded operators.
Aiming at the $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians, one can apply the Brockett-Wegner flow to a fermionic quadratic Hamiltonian $\mathrm{H}_{0}$ via the particle number operator $\mathrm{N}$ defined by (8), both acting on the Fock space $\mathcal{F}$ (3). Formally, it corresponds to the evolution equation
$\forall t\in\mathbb{R}_{0}^{+}:\qquad\partial_{t}\mathrm{H}_{t}=\left[\mathrm{H}_{t},\left[\mathrm{H}_{t},\mathrm{N}\right]\right]\ ,\qquad\mathrm{H}_{t=0}\doteq\mathrm{H}_{0}\ ,$ (30)
with $[A,B]\doteq AB-BA$ being the commutator between two operators $A$ and $B$ (provided it exists on some domain). This operator flow is very useful because it is awaited that the solution to (30) is a family of unitarily equivalent quadratic operators: For any $s\in\mathbb{R}_{0}^{+}$ and $t\in\left[s,\infty\right)$, $\mathrm{H}_{t}=\mathrm{U}_{t,s}\mathrm{\mathrm{H}}_{s}\mathrm{U}_{t,s}^{\ast}$ with $\left(\mathrm{U}_{t,s}\right)_{t\geq s}$ being a (evolution) family of unitary operators solution to the non-autonomous (hyperbolic in Kato’s terminology [28, 29], named Kato-hyperbolic here to avoid confusions) evolution equation
$\forall s\in\mathbb{R}_{0}^{+},\ t\in\left[s,\infty\right):\qquad\partial_{t}\mathrm{U}_{t,s}=-i\mathrm{G}_{t}\mathrm{U}_{t,s}\ ,\qquad\mathrm{U}_{s,s}\doteq\mathbf{1}\ ,$ (31)
with generator $\mathrm{G}_{t}\doteq i\left[\mathrm{N},\mathrm{H}_{t}\right]$. What is more, we expect the convergence in some sense of $\mathrm{H}_{t}$ and $\mathrm{U}_{t,0}$ in the limit $t\rightarrow\infty$, leading to a unitarily equivalent Hamiltonian $\mathrm{H}_{\infty}=\mathrm{U}_{\infty,0}\mathrm{\mathrm{H}}_{0}\mathrm{U}_{\infty,0}^{\ast}$ satisfying $\left[\mathrm{H}_{\infty},\mathrm{N}\right]=0$.
Nonetheless, establishing the well-posedness of the Brockett-Wegner flow (30) is highly non-trivial, because quadratic Hamiltonians are generally unbounded operators. The only known general result in the unbounded case concerns the local existence of its solution under some restricted conditions [21], which cannot be applied here. Moreover, even if the solution was proven to exist for all times, its asymptotics would even be harder to study for general Brockett-Wegner flows.
In order to avoid technical issues related to unboundedness and cumbersome environment inherent to the Fock space, it is of course natural to transfer the problem in the one-particle space. To this end, as in Section 2.4, on can use the space $\mathcal{H}=\mathfrak{h}\oplus\mathfrak{h}$ endowed with the antiunitary involution $\mathfrak{A}$ (21), instead of the Fock space. In this context, the particle number operator refers to the bounded (self-dual) operator
$\mathcal{N}\doteq\frac{1}{2}\left(\begin{array}[]{cc}\mathbf{1}&0\\ 0&-\mathbf{1}\end{array}\right)\in\mathcal{B}\left(\mathfrak{h}\oplus\mathfrak{h}\right)\ .$
Compare with Lemma 2.7 for $\Upsilon_{0}=\mathbf{1}$ and $D_{0}=0$. Formally, the self-dual version of the Brockett-Wegner flow (30) is the following differential equation in the strong sense on (some domain of) $\mathcal{H}=\mathfrak{h}\oplus\mathfrak{h}$:
$\forall t\in\mathbb{R}_{0}^{+}:\qquad\partial_{t}h_{t}=4\left[h_{t},\left[h_{t},\mathcal{N}\right]\right]\ ,\qquad h_{t=0}\doteq\frac{1}{2}\left(\begin{array}[]{cc}\Upsilon_{0}&2D_{0}\\ 2D_{0}^{\ast}&-\Upsilon_{0}^{\top}\end{array}\right)\ ,$ (32)
where $D_{0}=-D_{0}^{\top}\in\mathcal{B}\left(\mathfrak{h}\right)$ and $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ is a self-adjoint operator that is bounded from below. Having Lemma 2.7 in mind, one guesses that the solution to the last differential equation should be of the form
$h_{t}=\frac{1}{2}\left(\begin{array}[]{cc}\Upsilon_{t}&2D_{t}\\ -2\overline{D}_{t}&-\Upsilon_{t}^{\top}\end{array}\right)=\frac{1}{2}\left(\begin{array}[]{cc}\Upsilon_{t}&2D_{t}\\ 2D_{t}^{\ast}&-\Upsilon_{t}^{\top}\end{array}\right)\ ,\qquad t\in\mathbb{R}_{0}^{+}\ ,$ (33)
for some strongly continuous families of operators $\left(\Upsilon_{t}\right)_{t\geq 0}$ and $\left(D_{t}\right)_{t\geq 0}$ such that $D_{t}=-D_{t}^{\top}\in\mathcal{B}\left(\mathfrak{h}\right)$ and $\Upsilon_{t}=\Upsilon_{t}^{\ast}$ is a self-adjoint operator that is bounded from below, with always the time-independent domain $\mathcal{D}(\Upsilon_{0})$. In this case, one formally computes the commutators
$g_{t}\doteq 2i\left[\mathcal{N},h_{t}\right]=\frac{i}{2}\left(\begin{array}[]{cc}0&4D_{t}\\ -4D_{t}^{\ast}&0\end{array}\right)\ .$
Compare this generator with Lemma 2.7 and Equation (57). As a consequence,
$\partial_{t}h_{t}=4\left[h_{t},\left[h_{t},\mathcal{N}\right]\right]=2i\left[h_{t},g_{t}\right]=\frac{1}{2}\left(\begin{array}[]{cc}16D_{t}D_{t}^{\ast}&-4\left(\Upsilon_{t}D_{t}+D_{t}\Upsilon_{t}^{\top}\right)\\ -4\left(\Upsilon_{t}^{\top}D_{t}^{\ast}+D_{t}^{\ast}\Upsilon_{t}\right)&-16D_{t}^{\ast}D_{t}\end{array}\right)$
for all times $t\in\mathbb{R}_{0}^{+}$. By identification this leads precisely to the flow (191) (with $\Upsilon_{t}=\Upsilon_{0}+\Delta_{t}$) studied in [24].
Analogously to Equation (31) in the Fock space, one defines an evolution family of unitary operators solution to the (Kato-hyperbolic) non-autonomous evolution equation
$\forall s\in\mathbb{R}_{0}^{+},\ t\in\left[s,\infty\right):\qquad\partial_{t}U_{t,s}=-ig_{t}U_{t,s}\ ,\qquad U_{s,s}\doteq\mathbf{1}\ .$ (34)
By (21) and $D_{t}=-D_{t}^{\top}$ together with Remark 2.1, $g_{t}\mathfrak{A}=\mathfrak{A}g_{t}$ for all $t\in\mathbb{R}_{0}^{+}$ and, via the uniqueness of the solution to the above evolution equation, one can check that $\left(U_{t,s}\right)_{t\geq s}$ is a family of Bogoliubov transformations. In particular, $U_{t,s}\mathfrak{A=A}U_{t,s}$. Moreover, assuming a unique solution to the flow (32), we should have $h_{t}=U_{t,0}h_{0}U_{t,0}^{\ast}$ for all $t\in\mathbb{R}_{0}^{+}$. For instance, the Hilbert-Schmidt norm of $h_{(\cdot)}$ (if it exists) is constant, which can be rewritten as
$\frac{1}{4}\mathrm{tr}\left(\Upsilon_{t}^{2}+4D_{t}D_{t}^{\ast}\right)=\frac{1}{4}\mathrm{tr}\left(\Upsilon_{0}^{2}+4D_{0}D_{0}^{\ast}\right),\qquad t\in\mathbb{R}_{0}^{+}\ ,$
see (33). This gives an intuition of the elliptic nature of the flow on the (self-dual) one-particle space. (In the bosonic case, one does not use an inner product, but a symplectic form, leading heuristically to an hyperbolic flow through the same argument.) Rigorous treatment of the flow with generalized elliptic properties is developed in [24].
The Brockett-Wegner flow (32) takes place in the one-particle space, but we have to lift it again to the Fock space. In particular, the above family $\left(U_{t,s}\right)_{t\geq s}$ of Bogoliubov transformations has to be implemented, as explained in Remark 2.9. Mutatis mutandis for $\left(h_{t}\right)_{t\geq 0}$ and the relation $h_{t}=U_{t,0}h_{0}U_{t,0}^{\ast}$. This is done in Sections 3.2 and 3.3 whereas we present results on general implementation criteria of quadratic Hamiltonians in Section 2.6.
In fact, using the information we obtained in the one-particle space, the solution to the Brockett-Wegner flow (30) should be the following fermionic quadratic operator:
$\mathrm{H}_{t}\doteq\sum_{k,l\in\mathbb{N}}\left\{\Upsilon_{t}\right\}_{k,l}a_{k}^{\ast}a_{l}+\left\{D_{t}\right\}_{k,l}a_{k}^{\ast}a_{l}^{\ast}+\left\{\bar{D}_{t}\right\}_{k,l}a_{l}a_{k}+\left(E_{0}-8\int_{0}^{t}\left\|D_{\tau}\right\|_{2}^{2}\mathrm{d}\tau\right)\mathbf{1}$ (35)
with $\Upsilon_{t}\doteq\Upsilon_{0}+\Delta_{t}$ for each $t\in\mathbb{R}_{0}^{+}$ and where $\Delta$ and $D$ are solution to the elliptic operator-valued flow (191) studied in detail in [24]. Observe in this case that $D_{t}$ needs to be a Hilbert-Schmidt operator for all $t\in\mathbb{R}_{0}^{+}$ in order to invoke Proposition 3.4 for the self-adjointness of $\mathrm{H}_{t}$. In fact, using additionally the CAR (6), we should have
$D_{t}=-D_{t}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})\ ,\qquad t\in\mathbb{R}_{0}^{+}\ ,$ (36)
which is satisfied provided $D_{0}=D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$, see Theorem A.1. Therefore, if $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ and $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ then $\Upsilon_{t}$ is always self-adjoint (by definition of the elliptic flow) and the operator $\mathrm{H}_{t}$ is essentially self-adjoint on the domain $\mathcal{D}_{0}$ (5), thanks to Proposition 3.4. We use again the notation $\mathrm{H}_{t}\equiv\mathrm{H}_{t}^{\ast\ast}$ for its self-adjoint extension.
In other words, the Brockett-Wegner flow (30) on the many-fermion system can be indirectly analyzed via the Brockett-Wegner flow (32), i.e., the elliptic flow (191) on operators acting on the one-particle Hilbert space $\mathfrak{h}$. In particular, the existence of $\mathrm{H}_{\infty}$ and $\mathrm{U}_{\infty,0}$ can be studied rigorously via a detailed analysis of asymptotics of the operator families $\left(\Upsilon_{t}\right)_{t\geq 0}$ and $\left(D_{t}\right)_{t\geq 0}$. For instance, the convergence of the commutator $\left[\mathrm{H}_{t},\mathrm{N}\right]$ towards zero as $t\rightarrow\infty$ is related to the elimination of the operator $D_{t}$ in (35). The same strategy was used for the bosonic case in [6] by using an hyperbolic version of the elliptic flow (191) studied in [24].
In Section 3, we give the rigorous arguments based on this strategy, starting with the existence and uniqueness of a solution to the non-autonomous (Kato-hyperbolic) evolution equation (31) associated with the Brockett-Wegner flow (30). This approach eventually leads to Theorem 2.4.
2.6 Bach, Lieb and Solovej’s Approach to Quadratic Hamiltonians
Bach, Lieb and Solovej [15] introduced a definition of quadratic Hamiltonians that a priori differs from the one used here. This definition is inspired by Araki’s definition [18, Definition 5.1] and Theorem 2.8. For the reader’s convenience, it is reproduced below within the terminology and notation used in Section 2.4. We take here $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ and a Bogoliubov transformation is a unitary operator $U\in\mathcal{B}(\mathcal{H})$ such that $U\mathfrak{A=A}U$.
Definition 2.10 (Quadratic Hamilonians [15])
A self-adjoint operator $\mathrm{H}$ on the fermionic Fock space $\mathcal{F}$ is said to be a quadratic Hamiltonian if the unitary operator $\mathrm{U}_{t}=\mathrm{e}^{it\mathrm{H}/2}$ implements a Bogoliubov transformation $U_{t}$ for any $t\in\mathbb{R}$, i.e.,
$\mathrm{U}_{t}\left(\mathbf{\pi}_{\mathcal{F}}\left(A\right)\right)\mathrm{U}_{t}^{\ast}=\mathbf{\pi}_{\mathcal{F}}\circ\mathbf{\chi}_{U_{t}}\left(A\right)\ ,\qquad A\in\mathcal{A},\ t\in\mathbb{R}\ ,$ (37)
where $\mathbf{\chi}_{U_{t}}$ is the unique $\ast$–automorphism $\mathbf{\chi}_{U}$ of $\mathcal{A}$ satisfying (23) for $U=U_{t}$ while $\mathbf{\pi}_{\mathcal{F}}$ is the Fock space representation of the universal $C^{\ast}$-algebra $\mathcal{A}$ (19).
This point of view is also studied for bosonic systems, for instance by Bruneau and Derezinski via their definition of Bogoliubov Hamiltonians [7]. Observe also that, if $\mathrm{H}$ is a quadratic Hamiltonian as in Definition 2.10 then, for any real constant $C\in\mathbb{R}$$\mathrm{H}+C\mathbf{1}$ is of course a quadratic Hamiltonian leading to the same Bogoliubov transformation.
Quadratic operators in Bach, Lieb and Solovej’s sense are therefore automatically self-adjoint. In Section 2.2, following Berezin’s approach, we define quadratic operators on some restricted domain by using explicit sums over annihilation and creation operators acting on the fermionic Fock space. In this case, their essentially self-adjointness is an open issue solved by Proposition 3.4. It leads to self-adjoint operators that are always quadratic Hamiltonians in Bach, Lieb and Solovej’s sense:
Corollary 2.11 (From Berezin’s view point to BLS’s approach)
Fix $E_{0}\in\mathbb{R}$. Let $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ and $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ be two operators acting on the Hilbert space $\mathfrak{h}$. Then, the self-adjoint extension $\mathrm{H}_{0}\equiv\mathrm{H}_{0}^{\ast\ast}$ of the quadratic operator defined by (4)–(5) on the domain (5) is a quadratic Hamiltonian with respect to Definition 2.10.
Proof. Combine Lemma 2.7 with Definition 2.10.
We now discuss whether any quadratic Hamiltonian with respect to Definition 2.10 is the self-adjoint extension of a quadratic operators defined by (4) on the domain (5). To this end, we recall that $\mathfrak{A}$ is the antiunitary involution of $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ defined by (21), while $\{\mathrm{B}(\psi)\}_{\psi\in\mathcal{\mathcal{H}}}$ is the family of generators of the universal $C^{\ast}$-algebra $\mathcal{A}$ (19). We start with a first elementary, but important, lemma:
Lemma 2.12 (Generators of Bogoliubov transformations)
Let $\mathrm{H}$ be a quadratic Hamiltonian with respect to Definition 2.10. Then, (37) defines a strongly continuous group $(U_{t}=\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$, where $h$ is a self-adjoint operator acting on $\mathcal{H}$ satisfying $h=h^{\ast}=-\mathfrak{A}h\mathfrak{A}$ with domain $\mathcal{D}(h)=\mathfrak{A}(\mathcal{D}(h))\subseteq\mathcal{\mathcal{H}}$.
Proof. As done in [15, Equation (2b.36)], for any quadratic Hamiltonian $\mathrm{H}$ with respect to Definition 2.10, one computes from the CAR (20) and $U_{t}\mathfrak{A=A}U_{t}$ that, for any $\varphi,\psi\in\mathcal{H}$ and $t\in\mathbb{R}$,
[…]
where $\left\{A,B\right\}\doteq AB+BA$ is the usual anticommutator. So, by elementary computations using the strong continuity of the family $\left(\mathrm{e}^{it\mathrm{H}/2}\right)_{t\in\mathbb{R}}$ of unitary operators, we deduce that $(U_{t})_{t\in\mathbb{R}}$ is a strongly continuous one-parameter group of Bogoliubov transformations, see [31, Chapter I, Theorem 5.8]. In particular, we infer from Stone’s theorem [44, Theorem 6.2] the existence of some self-adjoint operator $h=h^{\ast}$ acting on $\mathcal{H}$ such that $U_{t}=\mathrm{e}^{ith}$ for any $t\in\mathbb{R}$. The assertion then follows from Lemma 3.13.
A priori, one may represent on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ the Hamiltonian $h$ of Lemma 2.12 by some $2\times 2$ matrix
$h=\left(\begin{array}[]{cc}P_{\mathfrak{h}}hP_{\mathfrak{h}}&P_{\mathfrak{h}}hP_{\mathfrak{h}}^{\perp}\\ P_{\mathfrak{h}}^{\perp}hP_{\mathfrak{h}}&P_{\mathfrak{h}}^{\perp}hP_{\mathfrak{h}}^{\perp}\end{array}\right)$ (38)
with operator-valued coefficients [^11] By a slight abuse of notation, we use in an appropriate way the identification $\mathfrak{h}\oplus\{0\}\equiv\mathfrak{h}$ and $\{0\}\oplus\mathfrak{h}\equiv\mathfrak{h}$ to see $P_{\mathfrak{h}}hP_{\mathfrak{h}}$, $P_{\mathfrak{h}}hP_{\mathfrak{h}}^{\perp}$, $P_{\mathfrak{h}}^{\perp}hP_{\mathfrak{h}}$ and $P_{\mathfrak{h}}^{\perp}hP_{\mathfrak{h}}^{\perp}$ as four operators acting on $\mathfrak{h}$., $P_{\mathfrak{h}}$ being the orthogonal projection on $\mathfrak{h}\oplus\{0\}$ and $P_{\mathfrak{h}}^{\perp}\doteq\mathbf{1}-P_{\mathfrak{h}}$. If $h\in\mathcal{B}\left(\mathcal{H}\right)$ then one easily computes from $h=h^{\ast}=-\mathfrak{A}h\mathfrak{A}$ (Lemma 2.12) together with (21) and $X^{\top}\doteq\mathcal{C}X^{\ast}\mathcal{C}$, $\overline{X}\doteq\mathcal{C}X\mathcal{C}$ and $\overline{X}^{\top}=\overline{X^{\top}}=X^{\ast}$ (Remark 2.1) that the block operator matrix (38) is well-defined and of the form
$h=\frac{1}{2}\left(\begin{array}[]{cc}\Upsilon_{0}&2D_{0}\\ 2D_{0}^{\ast}&-\Upsilon_{0}^{\top}\end{array}\right)=\frac{1}{2}\left(\begin{array}[]{cc}\Upsilon_{0}&2D_{0}\\ -2\overline{D}_{0}&-\Upsilon_{0}^{\top}\end{array}\right)\ ,$ (39)
where $\Upsilon_{0}\doteq P_{\mathfrak{h}}hP_{\mathfrak{h}}=\Upsilon_{0}^{\ast}$ and $D_{0}\doteq P_{\mathfrak{h}}hP_{\mathfrak{h}}^{\perp}=-D_{0}^{\top}$ are two bounded operators acting on $\mathfrak{h}$. By comparing Equation (39) with Lemma 2.7, one sees that quadratic Hamiltonians in the sense of Definition 2.10 or (4)–(5) are similarly encoded on the (dubbed) Hilbert space $\mathcal{H}$, at least when the Hamiltonian $h$ is bounded.
Nevertheless, general block operator matrices like (38) with unbounded operators are generally nontrivial to define. See for instance [37, 38, 39, 40, 41, 42, 43] which analyze properties of operators (e.g., self-adjointness, closeness) given via a $2\times 2$ matrix of unbounded operator-valued coefficients. In the present case, if the Hamiltonian $h$ of Lemma 2.12 is unbounded, with domain $\mathcal{D}(h)\subseteq\mathcal{\mathcal{H}}$, then it is for instance necessary to ensure at least that $\mathfrak{h}\cap\mathcal{D}(h)\neq 0$ in order to give a mathematically rigorous meaning of the coefficients $P_{\mathfrak{h}}hP_{\mathfrak{h}}$, $P_{\mathfrak{h}}hP_{\mathfrak{h}}^{\perp}$, $P_{\mathfrak{h}}^{\perp}hP_{\mathfrak{h}}$ and $P_{\mathfrak{h}}^{\perp}hP_{\mathfrak{h}}^{\perp}$ of (38) as well-defined operators acting on $\mathfrak{h\equiv h}\oplus\{0\}\equiv\{0\}\oplus\mathfrak{h}$. Additionally, even if $\mathfrak{h}\cap\mathcal{D}(h)\neq 0$, it is not clear that $\Upsilon_{0}$ is a (essentially) self-adjoint operator while one has to make sense of the sum $\Upsilon_{0}+2D_{0}$ with $\Upsilon_{0}\doteq P_{\mathfrak{h}}hP_{\mathfrak{h}}$ and $D_{0}\doteq P_{\mathfrak{h}}hP_{\mathfrak{h}}^{\perp}$ being possibly unbounded operators. There is in particular important domain issues associated with these operators $\Upsilon_{0}$ and $D_{0}$.
For instance, in Lemma 3.14 (ii), we show that, as soon as the kernel of $h$ has even or infinite dimension, $h$ can be represented on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ by (39), where $\Upsilon_{0}$ and $D_{0}$ are two (possibly unbounded) operators acting on $\mathfrak{h}$ with some dense domain
$\mathcal{Y}=\mathcal{D}\left(\Upsilon_{0}\right)\subseteq\mathfrak{h\qquad}\text{and}\mathfrak{\qquad}\mathcal{D}\left(D_{0}\right)=\mathcal{CY}\ ,$ (40)
where $\mathcal{C}$ is the complex conjugation of Remark 2.1. Note in this case that
$\mathcal{D}\left(h\right)=\mathcal{Y}\oplus\mathcal{CY}\ .$ (41)
However, even in this situation, $\Upsilon_{0}$ is only a symmetric operator while $D_{0}=-D_{0}^{\top}$ on its domain of definition. To get a self-adjoint operator $\Upsilon_{0}=\Upsilon_{0}^{\ast}$, it is necessary to construct a self-adjoint extension of the previous symmetric operator. This can be performed via the use of the Cayley and Krein transforms or the notion of boundary triplets [44, Part VI]. For example, as a consequence of the Cayley transform, von Neumann’s extension theorem [44, Theorem 13.10] demonstrates the existence of self-adjoint extensions of a densely defined symmetric operator $S$ iff the deficiency indices of $S$ are identical [^12] If this is not the case, one may use larger Hilbert space than $\mathcal{H}$, by [44, Corollary 13.11].. This property holds true whenever $S$ is semibounded, thanks to [44, Proposition 3.3]. Applied to the densely defined symmetric operator $\Upsilon_{0}$ of Lemma 3.14 (ii), these mathematical observations strongly highlight our condition $\Upsilon_{0}\geq-\tilde{\mu}\mathbf{1}$ for some $\tilde{\mu}\in\mathbb{R}$ given in Theorem 2.4, as a sufficient condition to obtain a self-adjoint operator $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ from Lemma 3.14 (ii).
A general and exhaustive discussion of fully unbounded cases like in Lemma 3.14 (ii) is therefore nontrivial with rather technical arguments and proofs. This simply goes beyond the main objective of the paper. In fact, even Araki in [18, Theorems 5.4 and 5.6] or Bruneau and Derezinski in the bosonic case [7] use (essentially) self-adjoint operators $h$ which can be written as a block operator matrix like (39) (in the fermionic case) and we thus restrict the definition of quadratic Hamiltonians given in Definition 2.10 by using the notion of compatible quadratic Hamiltonians:
Definition 2.13 (Compatible quadratic Hamilonians)
A quadratic Hamiltonian of Definition 2.10 is said to be compatible with the (one-particle) Hilbert space $\mathfrak{h}$ whenever its associated Hamiltonian $h$, as given by Lemma 2.12, can be written as (39) with domain $\mathcal{D}\left(h\right)=\mathcal{Y}\oplus\mathcal{CY}$, where $\mathcal{Y}\subseteq\mathfrak{h}$ is a dense subspace while $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ and $D_{0}=-D_{0}^{\top}$ are two operators with domains $\mathcal{D}\left(\Upsilon_{0}\right)\supseteq\mathcal{Y}$ and $\mathcal{D}\left(D_{0}\right)\supseteq\mathcal{C}\mathcal{Y}$.
From now, we only study compatible quadratic Hamiltonians. This case is one standard situation in Physics for which $\Upsilon_{0}$ is interpreted as the Hamiltonian acting on the one-particle Hilbert space describing free fermions while $D_{0}$ could result from effective theories, as already explained in Section 2.2. We emphasize again that a generator of a group of Bogoliubov transformations is not necessarily compatible, in spite of Lemma 3.14 (ii), which can only gives the existence of densely defined symmetric operators $\Upsilon_{0}$, as explained above.
It turns out that the set of quadratic Hamiltonians defined via (4)–(5) includes all compatible quadratic Hamiltonians of Definition 2.10 having a Hilbert-Schmidt operator $D_{0}$:
Corollary 2.14 (From BLS’s approach to Berezin’s view point)
Take a compatible quadratic Hamiltonian $\mathrm{H}$ with respect to Definitions 2.10 and 2.13. If $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ in (39) then, for some constant $E_{0}\in\mathbb{R}$,
$\mathrm{H}=\mathrm{H}_{0}^{\ast\ast}\equiv\mathrm{H}_{0}\doteq\sum_{k,l\in\mathbb{N}}\{\Upsilon_{0}\}_{k,l}a_{k}^{\ast}a_{l}+\{D_{0}\}_{k,l}a_{k}^{\ast}a_{l}^{\ast}+\{\bar{D}_{0}\}_{k,l}a_{l}a_{k}+E_{0}\mathbf{1}\ .$
Proof. Fix all parameters of the corollary. Using the operators $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ and $D_{0}=-D_{0}^{\top}$ given by Definition 2.13 (see (39)) together with $E_{0}=0$, we can define via (4)–(5) a self-adjoint operator $\mathrm{H}_{0}^{\ast\ast}\equiv\mathrm{H}_{0}$ under the condition that $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$. This is a consequence of Proposition 3.4. By Lemma 2.7 and Definition 2.10, the (unitary) groups $\left(\mathrm{e}^{it\mathrm{H}/2}\right)_{t\in\mathbb{R}}$ and $\left(\mathrm{e}^{it\mathrm{H}_{0}/2}\right)_{t\in\mathbb{R}}$ implement the same Bogoliubov transformation and so, we deduce that
$\mathrm{e}^{-it\mathrm{H}_{0}/2}\mathrm{e}^{it\mathrm{H}/2}\mathbf{\pi}_{\mathcal{F}}\left(A\right)=\mathbf{\pi}_{\mathcal{F}}\left(A\right)\mathrm{e}^{-it\mathrm{H}_{0}/2}\mathrm{e}^{it\mathrm{H}/2}\ ,\qquad t\in{\mathbb{R}},\text{ }A\in\mathcal{A}\ ,$
i.e., for any $t\in{\mathbb{R}}$, $\mathrm{e}^{-it\mathrm{H}_{0}/2}\mathrm{e}^{it\mathrm{H}/2}$ is in the commutant of $\pi_{\mathcal{F}}(A)$. Hence, by irreducibility of the Fock representation,
$\mathrm{e}^{-it\mathrm{H}_{0}/2}\mathrm{e}^{it\mathrm{H}/2}=c_{t}\in\mathbb{C}\ ,$ (42)
see Equations (169)–(170) for more details. Using the group property of $(\mathrm{e}^{it\mathrm{H}/2})_{t\in\mathbb{R}}$ and $(\mathrm{e}^{it\mathrm{H}_{0}/2})_{t\in\mathbb{R}}$ note that
$c_{t+s}\mathrm{e}^{i\left(t+s\right)\mathrm{H}_{0}/2}=\mathrm{e}^{i\left(t+s\right)\mathrm{H}/2}=c_{t}c_{s}\mathrm{e}^{i\left(t+s\right)\mathrm{H}_{0}/2}\ ,\qquad s,t\in{\mathbb{R}}\ .$ (43)
Additionally, for any $s,t\in{\mathbb{R}}$ and every normalized vector $\varphi\in\mathcal{F}$, i.e., vectors $\varphi$ in the Fock space $\mathcal{F}$ satisfying $\left\|\varphi\right\|_{\mathcal{F}}=1$,
$\left|c_{t}\right|=\left\|c_{t}\varphi\right\|_{\mathcal{F}}=\left\|\mathrm{e}^{-it\mathrm{H}_{0}/2}\mathrm{e}^{it\mathrm{H}/2}\varphi\right\|_{\mathcal{F}}=\left\|\varphi\right\|_{\mathcal{F}}=1$ (44)
as well as
$\left|c_{t}-c_{s}\right|$ $=$ $\left\|\left(c_{t}-c_{s}\right)\varphi\right\|_{\mathcal{F}}\leq\left\|\left(\mathrm{e}^{-it\mathrm{H}_{0}/2}\mathrm{e}^{it\mathrm{H}/2}-\mathrm{e}^{-is\mathrm{H}_{0}/2}\mathrm{e}^{is\mathrm{H}/2}\right)\varphi\right\|_{\mathcal{F}}$ (45)
$\leq$ $\left\|\mathrm{e}^{-it\mathrm{H}_{0}/2}\right\|_{\mathrm{op}}\left\|\left(\mathrm{e}^{it\mathrm{H}/2}-\mathrm{e}^{is\mathrm{H}/2}\right)\varphi\right\|_{\mathcal{F}}+\left\|\left(\mathrm{e}^{-it\mathrm{H}_{0}/2}-\mathrm{e}^{-is\mathrm{H}_{0}/2}\right)\mathrm{e}^{is\mathrm{H}/2}\varphi\right\|_{\mathcal{F}}\ ,$
thanks to the triangle inequality. Since $(\mathrm{e}^{it\mathrm{H}/2})_{t\in\mathbb{R}}$ and $(\mathrm{e}^{it\mathrm{H}_{0}/2})_{t\in\mathbb{R}}$ are strongly continuous groups of unitary operators, it follows from (43)–(45) that $(c_{t})_{t\in\mathbb{R}}\in C\left(\mathbb{R};\mathbb{C}\right)$ is a continuous complex-valued function satisfying
$\left|c_{t}\right|=1\qquad\text{and}\qquad c_{t+s}=c_{t}c_{s}$
for any $s,t\in{\mathbb{R}}$. As is well-known, there is therefore a real constant $\theta\in\mathbb{R}$ such that $c_{t}=\mathrm{e}^{it\theta/2}$ for any $t\in{\mathbb{R}}$. Combining this last property with (42) and the uniqueness of the generator of the strongly continuous group $\left(\mathrm{e}^{it\mathrm{H}/2}=\mathrm{e}^{it\left(\mathrm{H}_{0}+\theta\right)/2}\right)_{t\in\mathbb{R}}$ of unitary operators (see [44, Theorem 6.2]), we arrive at the assertion.
Corollary 2.14 is of course the converse of Corollary 2.11 in the Hilbert-Schmidt case. Note, however, that it is not clear whether the off-diagonal operator $D_{0}$ in Definition 2.13 (see (39)) is a bounded operator while it must be a Hilbert-Schmidt operator for the quadratic Hamiltonians defined in Section 2.2, as seen in Corollaries 2.11 and 2.14.
On the one hand, the Hilbert-Schmidt condition for $D_{0}$ results from the (natural) use of the domain $\mathcal{D}_{0}$ (5), which is a core for all our quadratic Hamiltonians. See Proposition 3.4. In particular, the vacuum state $\Psi\doteq\left(1,0,\ldots\right)$ of the fermionic Fock space $\mathcal{F}$ always belongs to the domain of our quadratic Hamiltonians, see Equation (7). By contrast, $\Psi$ may not be in the domain of a quadratic Hamiltonian with respect to Definition 2.10. But, a quadratic Hamiltonian $\mathrm{H}$ of Definition 2.10 that does not include $\Psi$ in its domain would be intrinsically different from our quadratic Hamiltonians and would refer to an abnormal model in Physics. We strongly suspect that the vacuum state would not belong to the domain of H when the operator $D_{0}$ of Definition 2.13 (see (39)) is not a Hilbert-Schmidt operator. This conjecture is based on our next theorems.
Theorem 2.15 (Shale-Stinespring-like condition for generators – I)
Take any compatible quadratic Hamiltonian $\mathrm{H}$ with respect to Definitions 2.10 and 2.13, the domain of which is $\mathcal{D}\left(\mathrm{H}\right)\subseteq\mathcal{F}$. Assume that, for some $c\in\mathbb{R}^{+}\cup\{\infty\}$,
$\mathrm{e}^{it\Upsilon_{0}}\mathcal{Y}\subseteq\mathcal{Y}\ ,\qquad t\in\left(-c,c\right)\ ,$ (46)
with the operator family $(\overline{D}_{0}\mathrm{e}^{it\Upsilon_{0}})_{t\in(-c,c)}$ being strongly continuous on $\mathcal{Y}$. Then, $\Psi\doteq\left(1,0,\ldots\right)\in\mathcal{D}\left(\mathrm{H}\right)$ iff $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ [^13] By a slight abuse of notation, we mean here the original $D_{0}$ or its (well-defined) continuous extension to the whole Hilbert space $\mathfrak{h}$..
Proof. Combine Corollaries 2.14 and 3.27 (ii), keeping in mind Definitions 2.10 and 2.13.
Remark 2.16
Condition (46) implies that $\mathcal{Y}$ must be a core for the self-adjoint operator $\Upsilon_{0}$, by [44, Proposition 6.3]. In other words, $\Upsilon_{0}$ restricted to the subspace $\mathcal{Y}\subseteq\mathcal{D}\left(\Upsilon_{0}\right)$ is essentially self-adjoint.
In applications, it is very natural to have an off-diagonal operator-valued coefficient $\overline{D}_{0}$ that is relatively bounded with respect of the diagonal one $\Upsilon_{0}$. For the block operator matrix (39) with densely defined linear operators $\Upsilon_{0}$ and $\overline{D}_{0}$ acting on $\mathfrak{h}$, it means that $\mathcal{D}(\overline{D}_{0})\supseteq\mathcal{D}(\Upsilon_{0})$ and that there exist constants $\alpha_{1},\alpha_{2}\in\mathbb{R}^{+}$ such that, for any $\varphi\in\mathcal{D}\left(\Upsilon_{0}\right)$,
$\left\|\overline{D}_{0}\varphi\right\|_{\mathfrak{h}}\leq\alpha_{1}\left\|\varphi\right\|_{\mathfrak{h}}+\alpha_{2}\left\|\Upsilon_{0}\varphi\right\|_{\mathfrak{h}}\ .$ (47)
See [45, Equation (X.19a)]. It is an usual assumption which enables well-defined block operator matrices like (39). See for instance [37, 43, 46]. In Physics, it means that the off-diagonal terms $D_{0}$,$\overline{D}_{0}$ can be viewed as a bounded perturbation of the free system, characterized  by the free (one particle) Hamiltonian $\Upsilon_{0}$. On the level of quadratic Hamiltonians, it means that the non-$\mathrm{N}$–diagonal part (see (95)) should be relatively bounded with respect to the $\mathrm{N}$–diagonal part of the given quadratic Hamiltonian. The latter is in fact true whenever $D_{0}$ is not only relatively bounded with respect $\Upsilon_{0}$ in the operator norm as in (47) but in the Hilbert-Schmidt norm, in the sense of Lemma 3.1 with $\theta=\Upsilon_{0}+\mu\mathbf{1}\geq 0$ for some appropriate $\mu\in\mathbb{R}$. The Shale-Stinespring-like condition for such generators can be studied by applying directly Theorem 2.15:
Corollary 2.17 (Shale-Stinespring-like condition for generators – II)
Take any compatible quadratic Hamiltonian $\mathrm{H}$ with respect to Definitions 2.10 and 2.13, the domain of which is $\mathcal{D}\left(\mathrm{H}\right)\subseteq\mathcal{F}$. Assume that $D_{0}$ is relatively bounded with respect of $\Upsilon_{0}$. Then, $\Psi\doteq\left(1,0,\ldots\right)\in\mathcal{D}\left(\mathrm{H}\right)$ iff $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$.
Proof. Combine Theorem 2.15 with Lemma 3.16.
Similar to the usual Shale-Stinespring condition given by Theorem 2.8, the property $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ in Theorem 2.15 and Corollary 2.17 can be seen as a Shale-Stinespring condition for generators of implementable groups of Bogoliubov transformations. This condition is directly related with the question of whether the vacuum state belongs to the domain of the quadratic Hamiltonian or not. When it is the case, we obtain the quadratic Hamiltonians diagonalized in the present paper.
We now conclude this section by studying the specific case for which the off-diagonal operator-valued coefficient $D_{0}$ of Definition 2.13 (see (39)) is not only relatively bounded with respect to $\Upsilon_{0}$, but already bounded. In this situation, we lurch from Corollary 2.17 by not assuming that the vacuum state $\Psi\doteq\left(1,0,\ldots\right)\in\mathcal{F}$ belongs to the domain of the compatible quadratic Hamiltonians. Even in this case, the operator $D_{0}$ of Definition 2.13 turns out to be very close to a Hilbert-Schmidt operator. To see this, we define the continuous operator family $(\varkappa_{t})_{t\in\mathbb{R}}\subseteq\mathcal{B}(\mathcal{B}(\mathfrak{h}))$ by
$\varkappa_{t}\left(A\right)\doteq\int_{0}^{t}\mathrm{e}^{-i\tau\Upsilon_{0}/2}A\mathrm{e}^{-i\tau\Upsilon_{0}^{\top}/2}\mathrm{d}\tau\ ,\qquad A\in\mathcal{B}(\mathfrak{h})\ ,$ (48)
for each time $t\in\mathbb{R}$. The integral is well-defined for any vector $\varphi\in\mathcal{H}$ since $(\mathrm{e}^{it\Upsilon_{0}})_{t\in\mathbb{R}}$ and $(\mathrm{e}^{it\Upsilon_{0}^{\top}})_{t\in\mathbb{R}}$ are strongly continuous groups of unitary operators.
Theorem 2.18 (Shale-Stinespring-like condition for generators – III)
Take any compatible quadratic Hamiltonian $\mathrm{H}$ with respect to Definitions 2.10 and 2.13, the domain of which is $\mathcal{D}\left(\mathrm{H}\right)\subseteq\mathcal{F}$. Assume $D_{0}\in\mathcal{B}(\mathfrak{h})$. Then, the following assertions hold true:
  1. (i)
    $(\varkappa_{t}(D_{0}))_{t\in\mathbb{R}}\subseteq\mathcal{B}(\mathfrak{h})$ belongs to $\mathcal{L}^{2}(\mathfrak{h})$ for times $t$ in a neighborhood of 00 and is even continuous in $\mathcal{L}^{2}(\mathfrak{h})$ at $t=0$.
  2. (ii)
    If $D_{0}\Upsilon_{0}=-\Upsilon_{0}^{\top}D_{0}$ then $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ and $\Psi\in\mathcal{D}\left(\mathrm{H}\right)$.
  3. (iii)
    If $\Upsilon_{0}\in\mathcal{B}(\mathfrak{h})$ and $D_{0}\Upsilon_{0}=\Upsilon_{0}^{\top}D_{0}$ then $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ and $\Psi\in\mathcal{D}\left(\mathrm{H}\right)$.
Proof. Assertion (i) results from Corollary 3.27 combined with Definitions 2.10 and 2.13. We deduce Assertions (ii)–(iii) by means of Theorem 2.8, Corollary 2.17 and Lemma 3.24, keeping in mind again Definitions 2.10 and 2.13. See also Remark 3.25 for Assertion (iii).
Note that, for simplicity, we restrict the last theorem to bounded off-diagonal elements $D_{0}\in\mathcal{B}(\mathfrak{h})$ since a general and exhaustive discussion of the fully unbounded cases would a priori deserve much longer discussions with rather technical arguments and new proofs. Here, the most general case is given by Theorem 2.18 (i), which shows that implementable strongly continuous groups of Bogoliubov transformations lead to operators $D_{0}$ that are very close to a Hilbert-Schmidt operator (if it is not always a Hilbert-Schmidt) since, for any $\varphi\in\mathfrak{h}$ and $A\in\mathcal{B}(\mathfrak{h})$,
$\varkappa_{t}\left(A\right)\varphi=tA\varphi+\mathcal{O}\left(1\right)\ ,\qquad\text{as\quad}t\rightarrow 0\ .$
This concludes our study of Bach, Lieb and Solovej’s approach to quadratic Hamiltonians in relation with Berezin’s approach used here.
3 Technical Results
Before starting the proofs, we add a few general definitions. First, $s\wedge t$ and $s\vee t$ stand respectively for the minimum and the maximum of real numbers $s,t\in\mathbb{R}$. This notation has not to be mixed with the exterior product defining the subspace $\wedge^{n}\mathfrak{h}$ of totally antisymmetric $n$–particle wave functions in $\mathfrak{h}^{\otimes n}$.
The Banach space of bounded operators from a Banach space $\left(\mathcal{X},\left\|\cdot\right\|_{\mathcal{X}}\right)$ to another Banach space $\left(\mathcal{Y},\left\|\cdot\right\|_{\mathcal{Y}}\right)$ is denoted by $\mathcal{B}(\mathcal{X},\mathcal{Y})$, with operator norm
$\left\|X\right\|_{\mathrm{op}}\doteq\sup_{f\in\mathcal{X}:\left\|f\right\|_{\mathfrak{h}}=1}\left\|Xf\right\|_{\mathcal{Y}}\ ,\qquad X\in\mathcal{B}(\mathcal{X},\mathcal{Y})\ .$
If $\mathcal{X=Y}$ then observe that $\mathcal{B}(\mathcal{X})\doteq\mathcal{B}(\mathcal{X},\mathcal{X})$ and recall that $\mathbf{1}\in\mathcal{B}(\mathcal{X})$ is the identity operator. Finally, $C\left(I;\mathcal{X}\right)$ denotes the Banach space of continuous functions from a closed set $I\subseteq\mathbb{R}$ to a Banach space $\left(\mathcal{X},\left\|\cdot\right\|_{\mathcal{X}}\right)$, along with the norm
$\left\|X\right\|_{\infty}\doteq\sup_{t\in I}\left\|X_{t}\right\|_{\mathcal{X}}\ ,\qquad X\in C\left(I;\mathcal{X}\right)\ .$
3.1 Self-adjointness of Quadratic Operators
In Equation (4) we define operators that are quadratic in the creation and annihilation operators on the fermionic Fock space $\mathcal{F}$ (3): Given $E_{0}\in\mathbb{R}$ and two operators $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ and $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ acting on the separable Hilbert space $\mathfrak{h}$, we formally define the operator
$\mathrm{H}_{0}\doteq d\Gamma(\Upsilon_{0})+\mathbf{D}+\mathbf{D}^{\ast}+E_{0}\mathbf{1}\ ,$ (49)
where
$d\Gamma\left(\theta\right)\doteq\sum_{k,l\in\mathbb{N}}\{\theta\}_{k,l}a_{k}^{\ast}a_{l}$ (50)
is the (well-known) second-quantization of any self-adjoint operator $\theta=\theta^{\ast}$ (here with domain $\mathcal{D}\left(\theta\right)\supseteq\left\{\varphi_{k}\right\}_{k=1}^{\infty}$), while
$\mathbf{D}\doteq\sum_{k,l\in\mathbb{N}}\{\bar{D}_{0}\}_{k,l}a_{l}a_{k}\ .$ (51)
Recall that $\left\{X\right\}_{k,l}\doteq\left\langle\varphi_{k},X\varphi_{l}\right\rangle_{\mathfrak{h}}$ for all operators $X$ acting on $\mathfrak{h}$, where $\left\{\varphi_{k}\right\}_{k=1}^{\infty}$ is a real orthonormal basis in the dense domain $\mathcal{D}\left(\Upsilon_{0}\right)\subseteq\mathfrak{h}$ of $\Upsilon_{0}$. Since $\mathfrak{h}$ can have infinite dimension, it is not clear that such quadratic operators are well-defined as (possibly unbounded) self-adjoint operators on $\mathcal{F}$. This issue is already studied in the literature, see for instance [11, Theorem 6.1]. However, the proof of [11, Theorem 6.1] is not completely rigorous. Another proof is given by [13, Proposition 2.1], but under the assumption of bounded operators $\Upsilon_{0}\in\mathcal{B}(\mathfrak{h})$. As a consequence, we explicitly establish the essential self-adjointness of $\mathrm{H}_{0}$ by translating the results of [7, Theorem 5.3] (itself inspired by [11, Theorem 6.1], following [14]) on bosonic quadratic operators to the fermionic case.
To this end, we first need a technical lemma, which is not only useful to the self-adjointness of fermionic quadratic operators but also pivotal to study the Brockett-Wegner flow (30) in Sections 3.23.5. In fact, we first show that the operators $\mathbf{D}$ and $d\Gamma(\Delta)$ with $\Delta=\Delta^{\ast}\in\mathcal{L}^{2}(\mathfrak{h})$ are well-defined on the domain $\mathcal{D}\left(\mathrm{N}\right)$ of the particle number operator $\mathrm{N}$, which is a self-adjoint operator that is explicitly defined by (9)–(10), see, e.g., [27, p. 7]. This is a simple corollary of the next assertion, which is similar to [6, Lemma 65] done in the bosonic case:
Lemma 3.1 (Relative boundedness of quadratic operators – I)
Let $\Delta=\Delta^{\ast}\in\mathcal{L}^{2}(\mathfrak{h})$ and $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ be two Hilbert-Schmidt operators. Given a positive, invertible, operator $\theta=\theta^{\ast}$ acting on $\mathfrak{h}$ with domain $\mathcal{D}\left(\theta\right)\supseteq\left\{\varphi_{k}\right\}_{k=1}^{\infty}$, the following inequalities hold true:
$\|\mathbf{D}\left(d\Gamma\left(\theta\right)+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}$$\leq$ $\left\|\theta^{-1/2}D_{0}\theta^{-1/2}\right\|_{2}\ ,$
$\|\mathbf{D}^{\ast}\left(d\Gamma\left(\theta\right)+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}$$\leq$ $\sqrt{\left\|\theta^{-1/2}D_{0}\theta^{-1/2}\right\|_{2}^{2}+2\left\|D_{0}\right\|_{2}^{2}}\ ,$
$\|d\Gamma(\Delta)\left(d\Gamma\left(\theta\right)+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}$$\leq$ $\sqrt{\left\|\theta^{-1/2}\Delta\theta^{-1/2}\right\|_{2}^{2}+\left\|\theta^{-1/2}\Delta^{2}\theta^{-1/2}\right\|_{2}}\ .$
Proof. The second quantization (50) of one-particle Hamiltonians does not depend of the choice of the orthonormal basis and we thus do the proof only when $\theta\geq\alpha\mathbf{1}$ for some $\alpha\in\mathbb{R}^{+}$ and
$d\Gamma\left(\theta\right)=\sum_{k\in\mathbb{N}}\left\{\theta\right\}_{k,k}a_{k}^{\ast}a_{k}\equiv\sum_{k\in\mathbb{N}}\theta_{k}a_{k}^{\ast}a_{k}$
where we use the shorter notation $\theta_{k}\equiv\left\{\theta\right\}_{k,k}$. The general case is completly identical, thanks to the spectral theorem. For more details, we recommend [6, Lemma 65]. In fact, in this paper we only need the case $\theta=\mathbf{1}$, i.e., $\theta_{k}=1$ for $k\in\mathbb{N}$. Take once and for all $\varphi\in\mathcal{F}$, $\Delta=\Delta^{\ast}\in\mathcal{L}^{2}(\mathfrak{h})$ and $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$. Using the Cauchy-Schwarz inequality we estimate that
$\left\|\mathbf{D}\varphi\right\|_{\mathcal{F}}$ $\leq\sum_{k,l\in\mathbb{N}}\frac{\left|\{\bar{D}_{0}\}_{k,l}\right|}{\theta_{k}^{1/2}\theta_{l}^{1/2}}\left\|\theta_{k}^{1/2}\theta_{l}^{1/2}a_{k}a_{l}\varphi\right\|_{\mathcal{F}}$
$\leq\left(\sum_{k,l\in\mathbb{N}}\theta_{k}^{-1}\left|\{D_{0}\}_{k,l}\right|^{2}\theta_{l}^{-1}\right)^{1/2}\left(\sum_{k,l\in\mathbb{N}}\theta_{k}\theta_{l}\left\|a_{k}a_{l}\varphi\right\|_{\mathcal{F}}^{2}\right)^{1/2}\ .$
Then, by using the CAR (6), it follows that
$\left\|\mathbf{D}\varphi\right\|_{\mathcal{F}}^{2}\leq\left\|\theta^{-1/2}D_{0}\theta^{-1/2}\right\|_{2}^{2}\sum_{k,l\in\mathbb{N}}\theta_{k}\theta_{l}\left\langle a_{l}\varphi,a_{k}^{\ast}a_{k}a_{l}\varphi\right\rangle_{\mathcal{F}}\\ =\left\|\theta^{-1/2}D_{0}\theta^{-1/2}\right\|_{2}^{2}\sum_{k,l\in\mathbb{N}}\left(1-\delta_{k,l}\right)\left\langle\theta_{l}a_{l}^{\ast}a_{l}\varphi,\theta_{k}a_{k}^{\ast}a_{k}\varphi\right\rangle_{\mathcal{F}}\leq\left\|\theta^{-1/2}D_{0}\theta^{-1/2}\right\|_{2}^{2}\left\|d\Gamma\left(\theta\right)\varphi\right\|_{\mathcal{F}}^{2}\ ,$
which proves the first inequality. Recall now that $\mathrm{tr}(\cdot)$ denotes the usual trace for operators. For the inequality with $\mathbf{D}^{\ast}$ we compute again from the CAR (6) that
$\left\|\mathbf{D}^{\ast}\varphi\right\|_{\mathcal{F}}^{2}$ $=$ $\sum_{k,l,p,q\in\mathbb{N}}\{\bar{D}_{0}\}_{k,l}\left\{D_{0}\right\}_{p,q}\left\langle\varphi,a_{l}a_{k}a_{p}^{\ast}a_{q}^{\ast}\varphi\right\rangle_{\mathcal{F}}$ (52)
$=$ $\sum_{k,l,p,q\in\mathbb{N}}\{\bar{D}_{0}\}_{k,l}\left\{D_{0}\right\}_{p,q}\left\langle a_{q}a_{p}\varphi,a_{l}a_{k}\varphi\right\rangle_{\mathcal{F}}$
$-\sum_{k,p\in\mathbb{N}}\left\{D_{0}D_{0}^{\ast}\right\}_{p,k}\left\langle a_{p}\varphi,a_{k}\varphi\right\rangle_{\mathcal{F}}$
$+\sum_{l,p\in\mathbb{N}}\left\{D_{0}\bar{D}_{0}\right\}_{p,l}\left\langle a_{p}\varphi,a_{l}\varphi\right\rangle_{\mathcal{F}}$
$-\sum_{k,q\in\mathbb{N}}\{\bar{D}_{0}D_{0}\}_{k,q}\left\langle a_{q}\varphi,a_{k}\varphi\right\rangle_{\mathcal{F}}$
$+\sum_{l,q\in\mathbb{N}}\{D_{0}^{\ast}D_{0}\}_{l,q}\left\langle a_{q}\varphi,a_{l}\varphi\right\rangle_{\mathcal{F}}+2\mathrm{tr}\left(D_{0}D_{0}^{\ast}\right)\left\langle\varphi,\varphi\right\rangle_{\mathcal{F}}$
$=$ $\sum_{k,l,p,q\in\mathbb{N}}\{\bar{D}_{0}\}_{k,l}\left\{D_{0}\right\}_{p,q}\left\langle a_{q}a_{p}\varphi,a_{l}a_{k}\varphi\right\rangle_{\mathcal{F}}+2\mathrm{tr}\left(D_{0}D_{0}^{\ast}\right)\left\langle\varphi,\varphi\right\rangle_{\mathcal{F}}$
using that $D_{0}=-D_{0}^{\top}$. It follows from the Cauchy-Schwarz inequality and explicit computations that
$\left\|\mathbf{D}^{\ast}\varphi\right\|_{\mathcal{F}}^{2}\leq\left\|\theta^{-1/2}D_{0}\theta^{-1/2}\right\|_{2}^{2}\left\|d\Gamma\left(\theta\right)\varphi\right\|_{\mathcal{F}}^{2}+2\left\|D_{0}\right\|_{2}^{2}\left\|\varphi\right\|_{\mathcal{F}}^{2}$ (53)
from which we deduce the upper norm estimate related to the adjoint $\mathbf{D}^{\ast}$. It remains to study the norms of $d\Gamma(\Delta)$ relative to $d\Gamma\left(\theta\right)$: By the CAR (6) and the Cauchy-Schwarz inequality, using the same trick to insert the coefficients of $\theta$, we find that
$\|d\Gamma(\Delta)\varphi\|_{\mathcal{F}}^{2}=-\sum_{k,l,p,q\in\mathbb{N}}\{\Delta\}_{k,l}\{\bar{\Delta}\}_{p,q}\left\langle a_{k}a_{q}\varphi,a_{p}a_{l}\varphi\right\rangle_{\mathcal{F}}+\sum_{l,q\in\mathbb{N}}\{\Delta^{2}\}_{q,l}\left\langle a_{q}\varphi,a_{l}\varphi\right\rangle_{\mathcal{F}}\\ \leq\left\|\theta^{-1/2}\Delta\theta^{-1/2}\right\|_{2}^{2}\left\|d\Gamma\left(\theta\right)\varphi\right\|_{\mathcal{F}}^{2}+\left\|\theta^{-1/2}\Delta^{2}\theta^{-1/2}\right\|_{2}\left\|d\Gamma\left(\theta\right)^{1/2}\varphi\right\|_{\mathcal{F}}^{2}\ ,$
which directly yields the last inequality.
Corollary 3.2 (Relative boundedness of quadratic operators – II)
Let $\Delta=\Delta^{\ast}\in\mathcal{L}^{2}(\mathfrak{h})$ and $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ be two Hilbert-Schmidt operators. Then, the following inequalities hold true:
$\|\mathbf{D}\left(\mathrm{N}+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}$$\leq$ $\left\|D_{0}\right\|_{2}\ ,$
$\|\mathbf{D}^{\ast}\left(\mathrm{N}+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}$$\leq$ $\sqrt{3}\left\|D_{0}\right\|_{2}\ ,$
$\|d\Gamma(\Delta)\left(\mathrm{N}+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}$$\leq$ $\sqrt{2}\left\|\Delta\right\|_{2}\ .$
Proof. This is an immediate application of Lemma 3.1, upon choosing $\theta\doteq\mathbf{1}$ because $d\Gamma\left(\mathbf{1}\right)=\mathrm{N}$.
Remark 3.3
The bounds of Lemma 3.1 and Corollary 3.2 are not supposed to be optimal or innovative, but are only sufficient for our uses. They are in fact standard estimates, the proofs being given only for completeness and to be as pedagogical as possible.
Having now Corollary 3.2 at our disposal, we are in a position to ensure the self-adjointness of $\mathrm{H}_{0}$ on the dense domain (5), that is,
$\mathcal{D}_{0}\doteq\bigcup_{N\in\mathbb{N}}\left(\bigoplus_{n=0}^{N}\left(\wedge^{n}\mathcal{D}\left(\Upsilon_{0}\right)\right)\right)\subseteq\mathcal{F}\ ,$
where $\mathcal{D}\left(\Upsilon_{0}\right)\subseteq\mathfrak{h}$ is the (dense) domain of $\Upsilon_{0}=\Upsilon_{0}^{\ast}$, $\wedge^{0}\mathcal{D}\left(\Upsilon_{0}\right)\doteq\mathbb{C}$ while for $n\in\mathbb{N}$, $\wedge^{n}\mathcal{D}\left(\Upsilon_{0}\right)$ is the (dense) subspace of totally antisymmetric $n$–particle wave functions in $\mathcal{D}\left(\Upsilon_{0}\right)^{\otimes n}$, the $n$–fold tensor product of $\mathcal{D}\left(\Upsilon_{0}\right)^{\otimes n}\subseteq\mathfrak{h}^{\otimes n}$.
Proposition 3.4 (Self–adjointness of quadratic operators)
Fix $E_{0}\in\mathbb{R}$. Let $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ and $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ be two operators acting on the Hilbert space $\mathfrak{h}$. Then, the operator $\mathrm{H}_{0}$ defined by (49)–(51) on the domain (5) is essentially self-adjoint.
[…]
$\varphi_{\pm\epsilon}\doteq\left(1\pm i\epsilon d\Gamma(\Upsilon_{0})\right)^{-1}\varphi\ ,$
[…]
$z\left\|P_{n}\varphi\right\|_{\mathcal{F}}^{2}=z\langle P_{n}\varphi,\varphi\rangle_{\mathcal{F}}=\lim_{\epsilon\rightarrow 0}z\langle P_{n}\varphi_{\epsilon},\varphi\rangle_{\mathcal{F}}=\lim_{\epsilon\rightarrow 0}\langle P_{n}\varphi_{\epsilon},\mathrm{H}_{0}^{\ast}\varphi\rangle_{\mathcal{F}}=\lim_{\epsilon\rightarrow 0}\langle\mathrm{H}_{0}P_{n}\varphi_{\epsilon},\varphi\rangle_{\mathcal{F}}$
and, in the same way,
$\overline{z}\left\|P_{n}\varphi\right\|_{\mathcal{F}}^{2}=\lim_{\epsilon\rightarrow 0}\langle\varphi,\mathrm{H}_{0}P_{n}\varphi_{-\epsilon}\rangle_{\mathcal{F}}\ .$
So, by summing and using (49)–(51) we obtain that
$2i\mathrm{Im}\left(z\left\|P_{n}\varphi\right\|_{\mathcal{F}}^{2}\right)$ $=$ $\lim_{\epsilon\rightarrow 0}\left\{\langle d\Gamma(\Upsilon_{0})P_{n}\varphi_{\epsilon},\varphi\rangle_{\mathcal{F}}-\langle\varphi,d\Gamma(\Upsilon_{0})P_{n}\varphi_{-\epsilon}\rangle_{\mathcal{F}}\right.$
$\left.+\langle(\mathbf{D}+\mathbf{D}^{\ast})P_{n}\varphi_{\epsilon},\varphi\rangle_{\mathcal{F}}-\langle\varphi,(\mathbf{D}+\mathbf{D}^{\ast})P_{n}\varphi_{-\epsilon}\rangle_{\mathcal{F}}\right\}$
[…]
$2i\mathrm{Im}\left(z\left\|P_{n}\varphi\right\|_{\mathcal{F}}^{2}\right)$ $=$ $\langle(\mathbf{D}+\mathbf{D}^{\ast})P_{n}\varphi,\varphi\rangle_{\mathcal{F}}-\langle\varphi,(\mathbf{D}+\mathbf{D}^{\ast})P_{n}\varphi\rangle_{\mathcal{F}}$
$=$ $\langle\mathbf{D}P_{n}\varphi,P_{n-2}\varphi\rangle_{\mathcal{F}}+\langle\mathbf{D}^{\ast}P_{n}\varphi,P_{n+2}\varphi\rangle_{\mathcal{F}}$
$-\langle\mathbf{D}P_{n+2}\varphi,P_{n}\varphi\rangle_{\mathcal{F}}-\langle\mathbf{D}^{\ast}P_{n-2}\varphi,P_{n}\varphi\rangle_{\mathcal{F}}$
[…]
$2i\mathrm{Im}\left(z\sum_{n=0}^{N}\left\|P_{n}\varphi\right\|_{\mathcal{F}}^{2}\right)$ $=$ $-\langle\mathbf{D}P_{N+2}\varphi,P_{N}\varphi\rangle_{\mathcal{F}}-\langle\mathbf{D}P_{N+1}\varphi,P_{N-1}\varphi\rangle_{\mathcal{F}}$
$+\langle\mathbf{D}^{\ast}P_{N}\varphi,P_{N+2}\varphi\rangle_{\mathcal{F}}+\langle\mathbf{D}^{\ast}P_{N-1}\varphi,P_{N+1}\varphi\rangle_{\mathcal{F}}\ .$
Then, we use the Cauchy-Schwarz inequality and apply Corollary 3.2 to arrive at the inequalities
$2\mathrm{Im}\left(z\sum_{n=0}^{N}\left\|P_{n}\varphi\right\|_{\mathcal{F}}^{2}\right)$ $\leq\left\|D_{0}\right\|_{2}\left(2(N+2)\left\|P_{N}\varphi\right\|_{\mathcal{F}}\left\|P_{N+2}\varphi\right\|_{\mathcal{F}}+2(N+1)\left\|P_{N-1}\varphi\right\|_{\mathcal{F}}\left\|P_{N+1}\varphi\right\|_{\mathcal{F}}\right)$
$\leq(N+2)\left\|D_{0}\right\|_{2}\left(\left\|P_{N-1}\varphi\right\|_{\mathcal{F}}^{2}+\left\|P_{N}\varphi\right\|_{\mathcal{F}}^{2}+\left\|P_{N+1}\varphi\right\|_{\mathcal{F}}^{2}+\left\|P_{N+2}\varphi\right\|_{\mathcal{F}}^{2}\right)\ .$ (54)
(Note that the last upper bound is obtained from the trivial inequality $2xy\leq x^{2}+y^{2}$ for $x,y\in\mathbb{R}_{0}^{+}$.) By contradiction, suppose that $\varphi\neq 0$. Necessarily there exist $N_{0}\in\mathbb{N}$ and a constant $C\in\mathbb{R}^{+}$ such that for all natural numbers $N\geq N_{0}$,
$\sum_{n=0}^{N}\left\|P_{n}\varphi\right\|_{\mathcal{F}}^{2}\geq\sum_{n=0}^{N_{0}}\left\|P_{n}\varphi\right\|_{\mathcal{F}}^{2}=C>0\ .$
As a result, for all natural numbers $N\geq N_{0}$, we deduce from (54) that
$\frac{2C\mathrm{Im}z}{N+2}\leq\left\|D_{0}\right\|_{2}\sum_{j=N-1}^{N+2}\left\|P_{j}\varphi\right\|_{\mathcal{F}}^{2}\ .$
If we sum up to $M\in\mathbb{N}$ with $M\geq N_{0}$ we obtain
$\sum_{N=N_{0}}^{M}\frac{2C\mathrm{Im}z}{N+2}\leq\left\|D_{0}\right\|_{2}\sum_{N=N_{0}}^{M}\sum_{j=N-1}^{N+2}\left\|P_{j}\varphi\right\|_{\mathcal{F}}^{2}\ .$
[…]
We conclude this section by showing that the essentially self-adjoint operator $\mathrm{H}_{0}$ defined by (49)–(51) on the domain (5) is independent of the (real) orthonormal basis $\left\{\varphi_{k}\right\}_{k=1}^{\infty}\subseteq\mathcal{D}\left(\Upsilon_{0}\right)\subseteq\mathfrak{h}$ chosen. So, take another orthonormal basis $\left\{\psi_{k}\right\}_{k=1}^{\infty}\subseteq\mathcal{D}\left(\Upsilon_{0}\right)\subseteq\mathfrak{h}$ and observe that, for any $k\in\mathbb{N}$,
$a_{k}\doteq a(\varphi_{k})=\sum_{n\in\mathbb{N}}\left\langle\varphi_{k},\psi_{n}\right\rangle_{\mathfrak{h}}a\left(\psi_{n}\right)$ (55)
by anti-linearity of the mapping $\varphi\mapsto a(\varphi)$ from $\mathfrak{h}$ to $\mathcal{B}\left(\mathcal{F}\right)$. The series are not necessarily absolutely convergent but they are always well-defined because of the CAR:
$a(\psi)a(\varphi)+a(\varphi)a(\psi)=0\ ,\qquad a(\psi)a(\varphi)^{\ast}+a(\varphi)^{\ast}a(\psi)=\langle\psi,\varphi\rangle_{\mathfrak{h}}\mathbf{1}$
for any $\varphi,\psi\in\mathfrak{h}$. These CAR are a generalization of (6). Therefore, the essentially self-adjoint operator $\mathrm{H}_{0}$ defined by (49)–(51) for $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ and $D_{0}=-D_{0}^{\top}$ is also equal to
$\mathrm{H}_{0}=\sum_{k,l\in\mathbb{N}}\left\{\Upsilon_{0}\right\}_{k,l}a\left(\psi_{k}\right)^{\ast}a\left(\psi_{l}\right)+\left\{D_{0}\mathcal{C}\right\}_{k,l}a\left(\psi_{k}\right)^{\ast}a\left(\psi_{l}\right)^{\ast}+\overline{\left\{D_{0}\mathcal{C}\right\}}_{k,l}a\left(\psi_{l}\right)a\left(\psi_{k}\right)+E_{0}\mathbf{1}$ (56)
on the domain (5), where $\mathcal{C}$ is the complex conjugation (Remark 2.1) and $\left\{X\right\}_{k,l}\doteq\left\langle\psi_{k},X\psi_{l}\right\rangle_{\mathfrak{h}}$ for all operators $X$ acting on $\mathfrak{h}$. As a consequence, the quadratic Hamiltonian is written exactly like in (49)–(51) when $\left\{\psi_{k}\right\}_{k=1}^{\infty}$ is a real orthonormal basis. In fact, $\mathrm{H}_{0}$ is basis independent and Equation (56) is the general form of the quadratic Hamiltonian $\mathrm{H}_{0}$ written with an arbitrary orthonormal basis in $\mathcal{D}\left(\Upsilon_{0}\right)$.
3.2 Continuous Flow of Bogoliubov Transformations
We need to implement on the Fock space a continuous family of Bogoliubov transformations generated by the solution to a flow on the one-particle Hilbert space. At first glance, this could be done directly via Theorem 2.8 and the Shale-Stinespring condition. For our applications, however, we need a non-autonomous flow of Bogoliubov transformations. To implement them, we invoke the theory of non-autonomous (Kato-hyperbolic) evolution equations. So, we use it on the Fock space because of the time-dependency of the generator in the Fock space.
Indeed, the continuous flow that implements Bogoliubov transformations is constructed from the unique solution to the non-autonomous evolution equation (31) with infinitesimal generator
$\mathrm{G}_{t}\doteq 2i\sum_{k,l\in\mathbb{N}}\left\{D_{t}\right\}_{k,l}a_{k}^{\ast}a_{l}^{\ast}+\left\{\bar{D}_{t}\right\}_{k,l}a_{k}a_{l}\ ,\qquad t\in\mathbb{R}_{0}^{+}\ ,$ (57)
for some continuous Hilbert-Schmidt operator family $D\in C(\mathbb{R}_{0}^{+};\mathcal{L}^{2}(\mathfrak{h}))$ satisfying $D_{t}=-D_{t}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}_{0}^{+}$. In our applications, $D$ is solution to the elliptic operator-valued flow (191) studied in detail in [24]. Note that the initial condition $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ leads to $D_{t}=-D_{t}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}_{0}^{+}$, thanks to Theorem A.1. Additionally, in this case, observe that the generator is formally equal to the commutator $\mathrm{G}_{t}=i\left[\mathrm{N},\mathrm{H}_{t}\right]$ with $\mathrm{N}$ and $\mathrm{H}_{t}$ being the particle number operator and the general quadratic Hamiltonian respectively defined by (8) and (35).
We need to prove that the corresponding non-autonomous evolution equation is well-established with time-dependent generators (57). Given $D\in C(\mathbb{R}_{0}^{+};\mathcal{L}^{2}(\mathfrak{h}))$ with $D_{t}=-D_{t}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}_{0}^{+}$, Proposition 3.4 proves that quadratic operators like the infinitesimal generators $\mathrm{G}_{t}$, $t\in\mathbb{R}_{0}^{+}$, all defined on the same (dense) domain
$\mathcal{D}_{1}\doteq\bigcup_{N\in\mathbb{N}}\left(\bigoplus_{n=0}^{N}\wedge^{n}\mathfrak{h}\right)\supseteq\mathcal{D}_{0}$
(see (5)), are essentially self-adjoint and, as already mentioned, we use the notation $\mathrm{G}_{t}\equiv\mathrm{G}_{t}^{\ast\ast}$, $t\in\mathbb{R}_{0}^{+}$, for their self-adjoint extension. In particular, the domain $\mathcal{D}_{1}$ is a core for all these self-adjoint operators, but also for the particle number operator $\mathrm{N}$ (see (8)) and we infer from Corollary 3.2 that the (now self-adjoint) generators $\mathrm{G}_{t}$, $t\in\mathbb{R}_{0}^{+}$, are relatively bounded with respect to the particle number operator $\mathrm{N}$: $\mathcal{D}(\mathrm{G}_{t})\supseteq\mathcal{D}(\mathrm{N})$ (see (10)) and
$\left\|\mathrm{G}_{t}\left(\mathrm{N}+\mathbf{1}\right)^{-1}\right\|_{\mathrm{op}}\leq 2(1+\sqrt{3})\left\|D_{t}\right\|_{2}<+\infty\ ,\qquad t\in\mathbb{R}_{0}^{+}\ ,$ (58)
thanks to $D_{t}=-D_{t}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ and Corollary 3.2.
The precise definition of infinitesimal generators of the non-autonomous evolution equation (31) being established, we show below the existence of a unique solution to this Kato-hyperbolic evolution equation, analogously to the line of arguments used in [6, Section VI.1]. Recall that $s\wedge t$ and $s\vee t$ stand respectively for the minimum and the maximum of real numbers $s,t\in\mathbb{R}$.
Theorem 3.5 (Continuous Bogoliubov transformations)
Assume $D\in C(\mathbb{R}_{0}^{+};\mathcal{L}^{2}(\mathfrak{h}))$ with $D_{t}=-D_{t}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}_{0}^{+}$. There is a family $(\mathrm{U}_{t,s})_{s,t\in\mathbb{R}_{0}^{+}}\subseteq\mathcal{B}\left(\mathcal{F}\right)$ of bounded operators satisfying the following properties:
  1. (i)
    For any $s,t\in\mathbb{R}_{0}^{+}$, $\mathrm{U}_{t,s}=\mathrm{U}_{s,t}^{\ast}$ is a unitary operator.
  2. (ii)
    It satisfies the cocycle property $\mathrm{U}_{t,x}\mathrm{U}_{x,s}=\mathrm{U}_{t,s}$ for any $s,x,t\in\mathbb{R}_{0}^{+}$.
  3. (iii)
    It is jointly strongly continuous in $s$ and $t$ on $(\mathbb{R}_{0}^{+})^{2}$.
  4. (iv)
    There is $C\in\mathbb{R}^{+}$ such that, for any $s,t\in\mathbb{R}_{0}^{+}$,
    $\|(\mathrm{N}+\mathbf{1})\mathrm{U}_{t,s}(\mathrm{N}+\mathbf{1})^{-1}\|_{\mathrm{op}}\leq\mathrm{\exp}\left\{C\int_{s\wedge t}^{s\vee t}\|D_{\tau}\|_{2}\mathrm{d}\tau\right\}$
    and the mapping $\left(s,t\right)\mapsto(\mathrm{N}+\mathbf{1})\mathrm{U}_{t,s}(\mathrm{N}+\mathbf{1})^{-1}$ is strongly continuous on $(\mathbb{R}_{0}^{+})^{2}$.
  5. (v)
[…]
    $\forall s,t\in\mathbb{R}_{0}^{+}:\qquad\left\{\begin{array}[]{llll}\partial_{t}\mathrm{U}_{t,s}=-i\mathrm{G}_{t}\mathrm{U}_{t,s}&,&\mathrm{U}_{s,s}\doteq\mathbf{1}&.\\ \partial_{s}\mathrm{U}_{t,s}=i\mathrm{U}_{t,s}\mathrm{G}_{s}&,&\mathrm{U}_{t,t}\doteq\mathbf{1}&.\end{array}\right.$
Proof. The proof is done in several steps:
Step 1: We first ensure the existence of a unique, bounded evolution operator $\left(\mathrm{U}_{t,s}\right)_{t\geq s\geq 0}$ satisfying on the domain $\mathcal{D}\left(\mathrm{N}\right)$ the non-autonomous evolution equations [^15] The derivatives $\partial_{t}$ and $\partial_{s}$ on the borderline $t=s$ or $s=0$ have to be understood as either right or left derivatives.
$\forall t\geq s\geq 0:\qquad\left\{\begin{array}[]{llll}\partial_{t}\mathrm{U}_{t,s}=-i\mathrm{G}_{t}\mathrm{U}_{t,s}&,&\mathrm{U}_{s,s}\doteq\mathbf{1}&.\\ \partial_{s}\mathrm{U}_{t,s}=i\mathrm{U}_{t,s}\mathrm{G}_{s}&,&\mathrm{U}_{t,t}\doteq\mathbf{1}&.\end{array}\right.$ (59)
This is done by using the general results of [6, Section VII.1] for the infinitesimal generator $G_{t}=\mathrm{G}_{t}$, the closed auxiliary operator $\Theta\doteq\mathrm{N}+\mathbf{1}$ and the space
$\mathcal{Y}\doteq\mathcal{D}\left(\mathrm{N}\right)=\left\{\varphi\in\mathcal{F}:\left\|\varphi\right\|_{\mathcal{Y}}\doteq\left\|\left(\mathrm{N}+\mathbf{1}\right)\varphi\right\|<\infty\right\}\ .$ (60)
Since $\mathrm{G}_{t}$ is self-adjoint for all times $t\in\mathbb{R}_{0}^{+}$, the assumption B1 of [6, Section VII.1] (with $m=1$ and $\beta_{0}(t)=0$) is immediate and we only need to verify a priori two conditions: [6, B2–B3 in Section VII.1]. It amounts to prove that the relative norms $\|\mathrm{G}_{t}(\mathrm{N}+\mathbf{1})^{-1}\|_{\mathrm{op}}$ and $\|[\mathrm{N},\mathrm{G}_{t}](\mathrm{N}+\mathbf{1})^{-1}\|_{\mathrm{op}}$ are bounded for any time $t\in\mathbb{R}_{0}^{+}$, with the function $t\mapsto\|\mathrm{G}_{t}(\mathrm{N}+\mathbf{1})^{-1}\|_{\mathrm{op}}$ being continuous on $\mathbb{R}_{0}^{+}$. Since a formal computation using the CAR shows that
$\left[\mathrm{N},\mathrm{G}_{t}\right]=4i\sum_{k,l\in\mathbb{N}}\left\{D_{t}\right\}_{k,l}a_{k}^{\ast}a_{l}^{\ast}+\left\{\bar{D}_{t}\right\}_{k,l}a_{l}a_{k}\ ,$ (61)
[6, B2–B3 in Section VII.1] are direct consequences of $D\in C(\mathbb{R}_{0}^{+};\mathcal{L}^{2}(\mathfrak{h}))$ together with Corollary 3.2 and Equation (57), which yield
$\max\left\{\|\mathrm{G}_{t}\left(\mathrm{N}+\mathbf{1}\right)^{-1}\|_{\mathrm{op}},\|\left[\mathrm{N},\mathrm{G}_{t}\right]\left(\mathrm{N}+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}\right\}\leq C\|D_{t}\|_{2}\ ,\qquad t\in\mathbb{R}_{0}^{+}\ ,$ (62)
for some (sufficiently large) constant $C\in\mathbb{R}^{+}$. Note that $\|D_{t}\|_{2}$ is uniformly bounded on compacta. As a consequence, we can apply [6, Theorem 88 and Lemma 92] and infer the existence of a unique, bounded evolution operator $(\mathrm{U}_{t,s})_{t\geq s\geq 0}$ satisfying the non-autonomous evolution equations (59) in the strong sense on the domain $\mathcal{D}\left(\mathrm{N}\right)$. Moreover, one obtains from (62) that, for any $t\geq s\geq 0$,
$\left\|(\mathrm{N}+\mathbf{1})\mathrm{U}_{t,s}(\mathrm{N}+\mathbf{1})^{-1}\right\|_{\mathrm{op}}\leq\mathrm{\exp}\left\{C\int_{s}^{t}\left\|D_{\tau}\right\|_{2}\mathrm{d}\tau\right\}\ .$
In particular, $\mathrm{U}_{t,s}$ conserves the domain $\mathcal{D}\left(\mathrm{N}\right)$ for all $t\geq s\geq 0$. The strong continuity of the bounded operator family
$\left\{(\mathrm{N}+\mathbf{1})\mathrm{U}_{t,s}(\mathrm{N}+\mathbf{1})^{-1}\right\}_{t\geq s\geq 0}$ (63)
also results from [6, Lemma 92].
Step 2: Observe that Step 1 neither implies that the adjoint $\mathrm{U}_{t,s}^{\ast}$ of the evolution operator $\mathrm{U}_{t,s}$ conserves the dense domain $\mathcal{D}\left(\mathrm{N}\right)$ of the particle number operator $\mathrm{N}$, nor that it is jointly strongly continuous in $s$ and $t$ for all $t\geq s\geq 0$. Therefore, in the same way we prove Step 1, one checks the existence of a unique, bounded evolution operator $(\mathrm{V}_{t,s})_{t\geq s\geq 0}$ satisfying on the domain $\mathcal{D}\left(\mathrm{N}\right)$ the non-autonomous evolution equations [^16] The derivatives $\partial_{t}$ and $\partial_{s}$ on the borderline $t=s$ or $s=0$ have to be understood as either right or left derivatives.
$\forall t\geq s\geq 0:\qquad\left\{\begin{array}[]{llll}\partial_{s}\mathrm{V}_{t,s}=-i\mathrm{G}_{s}\mathrm{V}_{t,s}&,&\mathrm{V}_{t,t}\doteq\mathbf{1}&.\\ \partial_{t}\mathrm{V}_{t,s}=i\mathrm{V}_{t,s}\mathrm{G}_{t}&,&\mathrm{V}_{s,s}\doteq\mathbf{1}&.\end{array}\right.$ (64)
Moreover, $\mathrm{V}_{t,s}$ conserves the domain $\mathcal{D}\left(\mathrm{N}\right)$ for all $t\geq s\geq 0$ as
$\|(\mathrm{N}+\mathbf{1})\mathrm{V}_{t,s}(\mathrm{N}+\mathbf{1})^{-1}\|_{\mathrm{op}}\leq\mathrm{\exp}\left\{C\int_{s}^{t}\|D_{\tau}\|_{2}\mathrm{d}\tau\right\}$ (65)
for some sufficiently large constant $C\in\mathbb{R}^{+}$, thanks to (62).
Step 3: We prove here the unitarity of the operators $\mathrm{U}_{t,s}$ for $t\geq s\geq 0$. For $t\geq s\geq 0$, note that [6, Theorem 88] gives both the operators $\mathrm{V}_{t,s}$ and $\mathrm{U}_{t,s}$ as limit $\lambda\rightarrow\infty$ of (evolution) operators $\mathrm{V}_{t,s,\lambda}=\mathrm{U}_{t,s,\lambda}^{\ast}$ and $\mathrm{U}_{t,s,\lambda}$ for $\lambda\gg 1$ defined like in [6, Equation VII.6], both having a norm convergent representation as a Dyson series. From this, one deduces that $\mathrm{V}_{t,s}=\mathrm{U}_{t,s}^{\ast}$ for all $t\geq s\geq 0$. For more details, see the arguments proving [6, Lemma 68]. Using now Steps 1 and 2, we deduce that, for any $t\geq s\geq 0$,
$(\mathbf{1}-\mathrm{U}_{t,s}\mathrm{U}_{t,s}^{\ast})(\mathrm{N}+\mathbf{1})^{-1}=\int_{s}^{t}\partial_{\tau}\left\{\mathrm{U}_{t,\tau}\mathrm{U}_{t,\tau}^{\ast}\right\}(\mathrm{N}+\mathbf{1})^{-1}\mathrm{d}\tau=0$
and
$(\mathbf{1}-\mathrm{U}_{t,s}^{\ast}\mathrm{U}_{t,s})(\mathrm{N}+\mathbf{1})^{-1}=-\int_{s}^{t}\partial_{\tau}\left\{\mathrm{U}_{\tau,s}^{\ast}\mathrm{U}_{\tau,s}\right\}(\mathrm{N}+\mathbf{1})^{-1}\mathrm{d}\tau=0\ .$
Thus, for all $t\geq s\geq 0$, $\mathrm{U}_{t,s}^{\ast}\mathrm{U}_{t,s}=\mathrm{U}_{t,s}\mathrm{U}_{t,s}^{\ast}=\mathbf{1}$ on the dense domain $\mathcal{D}\left(\mathrm{N}\right)\subseteq\mathcal{F}$, which immediately implies the unitarity of $\mathrm{U}_{t,s}\in\mathcal{B}\left(\mathcal{F}\right)$.
Step 4: We finally use the definition $\mathrm{U}_{t,s}\doteq\mathrm{U}_{s,t}^{\ast}$ for $s\geq t\geq 0$. Then, by combining Steps 1-3, one deduces all the assertions of the theorem.
We analyze now the effect of the unitary operators $\mathrm{U}_{t,s}$, $t,s\in\mathbb{R}_{0}^{+}$, on annihilation/creation operators $\left\{a_{k}\right\}_{k\in\mathbb{N}}$ and $\left\{a_{k}^{\ast}\right\}_{k\in\mathbb{N}}$. The aim is to relate the method used here to the so-called Bogoliubov transformation. To this end, we define the bounded operators
$\mathbf{u}_{t,s}$ $\doteq$ $\mathbf{1}+\sum_{n\in\mathbb{N}}\left(-1\right)^{n}4^{2n}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{s}^{\tau_{2n-1}}\mathrm{d}\tau_{2n}D_{\tau_{1}}D_{\tau_{2}}^{\ast}\cdots D_{\tau_{2n-1}}D_{\tau_{2n}}^{\ast}\ ,$ (66)
$\mathbf{v}_{t,s}$ $\doteq$ $\sum_{n\in\mathbb{N}_{0}}\left(-1\right)^{n}4^{2n+1}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{s}^{\tau_{2n}}\mathrm{d}\tau_{2n+1}D_{{\tau_{1}}}D_{\tau_{2}}^{\ast}\cdots D_{\tau_{2n}}^{\ast}D_{\tau_{2n+1}}\ ,$ (67)
with $\tau_{0}\doteq t$, while $D\in C(\mathbb{R}_{0}^{+};\mathcal{L}^{2}(\mathfrak{h}))$ with $D_{t}=-D_{t}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$. In our applications, it is solution to the elliptic operator flow (191). Note that the mappings $(s,t)\mapsto(\mathbf{u}_{t,s}-\mathbf{1})$ and $(s,t)\mapsto\mathbf{v}_{t,s}$ from $(\mathbb{R}_{0}^{+})^{2}$ to $\mathcal{L}^{2}(\mathfrak{h})$ are well-defined and continuous in the Hilbert-Schmidt topology, thanks to the bounds
$\left\|\mathbf{u}_{t_{1},s_{1}}-\mathbf{u}_{t_{2},s_{2}}\right\|_{2}$$\leq$ $\cosh\left(4\int_{t_{1}\wedge t_{2}}^{t_{1}\vee t_{2}}\left\|D_{\tau}\right\|_{2}\mathrm{d}\tau\right)+\cosh\left(4\int_{s_{1}\wedge s_{2}}^{s_{1}\vee s_{2}}\left\|D_{\tau}\right\|_{2}\mathrm{d}\tau\right)-2$
$\left\|\mathbf{v}_{t_{1},s_{1}}-\mathbf{v}_{t_{2},s_{2}}\right\|_{2}$$\leq$ $\sinh\left(4\int_{t_{1}\wedge t_{2}}^{t_{1}\vee t_{2}}\left\|D_{\tau}\right\|_{2}\mathrm{d}\tau\right)+\sinh\left(4\int_{s_{1}\wedge s_{2}}^{s_{1}\vee s_{2}}\left\|D_{\tau}\right\|_{2}\mathrm{d}\tau\right)$
for all $s_{1},s_{2},t_{1},t_{2}\in\mathbb{R}_{0}^{+}$. (Note that the above inequalities are straightforwardly deduced from the triangle inequality.) In particular, the series in (66) and (67) are absolutely convergent in $\mathcal{L}^{2}(\mathfrak{h})$. These coefficients allow one to define a Bogoliubov transformation which is implemented (in the sense of Theorem 2.8) by the unitary evolution system $(\mathrm{U}_{t,s})_{s,t\in\mathbb{R}_{0}^{+}}$ of Theorem 3.5:
Proposition 3.6 (Implementation of non-autonomous Bogoliubov transformations)
Assume $D\in C(\mathbb{R}_{0}^{+};\mathcal{L}^{2}(\mathfrak{h}))$ with $D_{t}=-D_{t}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}_{0}^{+}$. For all $s,t\in\mathbb{R}_{0}^{+}$,
$\mathrm{U}_{t,s}a\left(\varphi\right)\mathrm{U}_{t,s}^{\ast}=a(\mathbf{u}_{t,s}\varphi)+a^{\ast}(\mathbf{v}_{t,s}\overline{\varphi})\ ,\qquad\varphi\in\mathfrak{h}\ ,$ (68)
with the operators $\mathbf{u}_{t,s}$ and $\mathbf{v}_{t,s}$ satisfying the following equalities:
[…]
$\mathbf{u}_{t,s}\mathbf{u}_{t,s}^{\ast}-\mathbf{v}_{t,s}\mathbf{v}_{t,s}^{\ast}$ $=$ $\mathbf{1}\ ,\qquad\mathbf{u}_{t,s}\mathbf{v}_{t,s}^{\top}+\mathbf{v}_{t,s}\mathbf{u}_{t,s}^{\top}=0\ .$ (70)
Proof. Since $(D_{t})_{t\in\mathbb{R}_{0}^{+}}\in C(\mathbb{R}_{0}^{+};\mathcal{L}^{2}(\mathfrak{h}))$ , the operator families $(\mathbf{u}_{t,s})_{s,t\in\mathbb{R}_{0}^{+}}$ and $(\mathbf{v}_{t,s})_{s,t\in\mathbb{R}_{0}^{+}}$ satisfy [^17] The derivatives $\partial_{t}$ and $\partial_{s}$ on the borderline $t=0$ or $s=0$ have to be understood as right derivatives.
$\partial_{t}\mathbf{u}_{t,s}=4D_{t}\overline{\mathbf{v}}_{t,s}\qquad\text{and}\qquad\partial_{t}\mathbf{v}_{t,s}=4D_{t}\overline{\mathbf{u}}_{t,s}$ (71)
for all $s,t\in\mathbb{R}_{0}^{+}$, in the Hilbert-Schmidt topology. Therefore, using the continuity [^18] This follows from the inequalities $\|A\|_{2}=\|A^{\ast}\|_{2}=\|A^{\top}\|_{2}$ for any $A\in\mathcal{L}^{2}(\mathfrak{h})$. of the mappings $A\mapsto A^{\ast}$ and $A\mapsto A^{\top}$ in $\mathcal{L}^{2}(\mathfrak{h})$, one computes that
$\partial_{t}\left\{\mathbf{u}_{t,s}^{\ast}\mathbf{u}_{t,s}+\mathbf{v}_{t,s}^{\top}\overline{\mathbf{v}}_{t,s}\right\}=0\qquad\text{and}\qquad\partial_{t}\left\{\mathbf{u}_{t,s}^{\ast}\mathbf{v}_{t,s}+\mathbf{v}_{t,s}^{\top}\overline{\mathbf{u}}_{t,s}\right\}=0\ ,$
which in turn imply that
$\mathbf{u}_{t,s}^{\ast}\mathbf{u}_{t,s}+\mathbf{v}_{t,s}^{\top}\overline{\mathbf{v}}_{t,s}=\mathbf{u}_{s,s}^{\ast}\mathbf{u}_{s,s}+\mathbf{v}_{s,s}^{\top}\overline{\mathbf{v}}_{s,s}=\mathbf{1}$
as well as
$\mathbf{u}_{t,s}^{\ast}\mathbf{v}_{t,s}+\mathbf{v}_{t,s}^{\top}\overline{\mathbf{u}}_{t,s}=\mathbf{u}_{s,s}^{\ast}\mathbf{v}_{s,s}+\mathbf{v}_{s,s}^{\top}\overline{\mathbf{u}}_{s,s}=0\ .$
I.e., one gets Equation (69). To prove Equation (70), one performs similar computations together with the observation that
$\partial_{s}\mathbf{u}_{t,s}=4\mathbf{v}_{t,s}D_{s}^{\ast}\qquad\text{and}\qquad\partial_{s}\mathbf{v}_{t,s}=4\mathbf{u}_{t,s}D_{s}\ ,$
for all $s,t\in\mathbb{R}_{0}^{+}$. Note that these derivatives [^19] The derivatives $\partial_{t}$ and $\partial_{s}$ on the borderline $t=0$ or $s=0$ have to be understood as right derivatives. follow from $D_{t}=-D_{t}^{\top}$ and the equalities
$\mathbf{u}_{t,s}$ $=$ $\mathbf{1}+\sum_{n\in\mathbb{N}}\left(-1\right)^{n}4^{2n}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{\tau_{2n-1}}^{t}\mathrm{d}\tau_{2n}D_{\tau_{2n}}D_{\tau_{2n-1}}^{\ast}\cdots D_{\tau_{2}}D_{\tau_{1}}^{\ast}\ ,$
$\mathbf{v}_{t,s}$ $=$ $\sum_{n\in\mathbb{N}_{0}}\left(-1\right)^{n}4^{2n+1}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{\tau_{2n}}^{t}\mathrm{d}\tau_{2n+1}D_{{\tau_{2n+1}}}D_{\tau_{2n}}^{\ast}\cdots D_{\tau_{2}}^{\ast}D_{\tau_{1}}\ ,$
with $\tau_{0}\doteq s$.
Now, recall that $\left\{\varphi_{k}\right\}_{k=1}^{\infty}$ is some real orthonormal basis in $\mathcal{D}\left(\Upsilon_{0}\right)\subseteq\mathfrak{h}$. So, to prove Equation (68) for all $\varphi\in\mathfrak{h}$, it suffices to prove it for any basis element $\varphi_{k}$, thanks to the antilinearity of the mapping $\varphi\mapsto a(\varphi)$ from $\mathfrak{h}$ to $\mathcal{B}\left(\mathcal{F}\right)$. So, for each $k\in\mathbb{N}$ and $s,t\in\mathbb{R}_{0}^{+}$, we define the bounded operator $\tilde{a}_{t,s,k}\in\mathcal{B}\left(\mathcal{F}\right)$ by
$\tilde{a}_{t,s,k}\doteq a(\mathbf{u}_{t,s}\varphi_{k})+a^{\ast}(\mathbf{v}_{t,s}\varphi_{k})=\sum_{\ell\in\mathbb{N}}\left\{\mathbf{u}_{t,s}^{\ast}\right\}_{k,\ell}a_{\ell}+\left\{\mathbf{v}_{t,s}^{\top}\right\}_{k,\ell}a_{\ell}^{\ast}\ ,$ (72)
where $a_{k}\doteq a\left(\varphi_{k}\right)$ is the annihilation operator acting on the fermionic Fock space $\mathcal{F}$. Because of (69)–(70), for each $t\in\mathbb{R}_{0}^{+}$ the operator family $\{\tilde{a}_{t,s,k},\tilde{a}_{t,s,k}^{\ast}\}_{k=1}^{\infty}$ satisfies the CAR and since
$\left\|a(\varphi)\right\|_{\mathrm{op}}=\left\|\varphi\right\|_{\mathfrak{h}}\ ,\qquad\varphi\in\mathfrak{h}\ ,$ (73)
and $D_{t}=-D_{t}^{\top}$, we have the equalities
$\partial_{t}\tilde{a}_{t,s,k}=a(4D_{t}\overline{\mathbf{v}}_{t,s}\varphi_{k})+a^{\ast}(4D_{t}\overline{\mathbf{u}}_{t,s}\varphi_{k})$ (74)
in the Banach space $\mathcal{B}(\mathcal{F})$ for each $k\in\mathbb{N}$ and $t\in\mathbb{R}_{0}^{+}$, thanks to Equation (71). Meanwhile, observe from a formal computation using the CAR that, for all $k\in\mathbb{N}$ and $t\in\mathbb{R}_{0}^{+}$,
$\left[\mathrm{N}\mathbf{,}\tilde{a}_{t,s,k}\right]=\sum_{\ell\in\mathbb{N}}\left\{-\mathbf{u}_{t,s}^{\ast}\right\}_{k,\ell}a_{\ell}+\left\{\mathbf{v}_{t,s}^{\top}\right\}_{k,\ell}a_{\ell}^{\ast}=a(-\mathbf{u}_{t,s}\varphi_{k})+a^{\ast}(\mathbf{v}_{t,s}\varphi_{k})\ ,$
and from this observation and (72)–(73) one in fact checks that
$\left\|(\mathrm{N}+\mathbf{1)}\tilde{a}_{t,s,k}(\mathrm{N}+\mathbf{1)}^{-1}\right\|_{\mathrm{op}}\leq 2\|\mathbf{u}_{t,s}\|_{\mathrm{op}}+2\|\mathbf{v}_{t,s}\|_{\mathrm{op}}<\infty\ .$ (75)
Note that another formal computation using the CAR shows that
$i\left[\tilde{a}_{t,s,k},\mathrm{G}_{t}\right]$ $=$ $4\sum_{\ell\in\mathbb{N}}\left\{\mathbf{v}_{t,s}^{\top}D_{t}^{\ast}\right\}_{k,\ell}a_{\ell}-4\sum_{\ell\in\mathbb{N}}\left\{\mathbf{u}_{t,s}^{\ast}D_{t}\right\}_{k,\ell}a_{\ell}^{\ast}$
$=$ $a(4D_{t}\overline{\mathbf{v}}_{t,s}\varphi_{k})+a^{\ast}(4D_{t}\overline{\mathbf{u}}_{t,s}\varphi_{k})\ ,$
which is in fact well-defined on $\mathcal{D}(\mathrm{N})$.Then, by Equations (74)–(3.2) and Theorem 3.5, we arrive at the equality
$\partial_{t}\left\{\mathrm{U}_{t,s}^{\ast}\tilde{a}_{t,s,k}\mathrm{U}_{t,s}\psi\right\}=\mathrm{U}_{t,s}^{\ast}\left(\partial_{t}\tilde{a}_{t,s,k}-i\left[\tilde{a}_{t,s,k},\mathrm{G}_{t}\right]\right)\mathrm{U}_{t,s}\psi=0$ (77)
for all vectors $\psi\in\mathcal{D}(\mathrm{N})$ and $s,t\in\mathbb{R}_{0}^{+}$. Using again Theorem 3.5, we deduce that
$\tilde{a}_{t,s,k}\psi=\mathrm{U}_{t,s}\tilde{a}_{k}\mathrm{U}_{t,s}^{\ast}\psi$ (78)
for all $k\in\mathbb{N}$, $\psi\in\mathcal{D}(\mathrm{N})$ and $s,t\in\mathbb{R}_{0}^{+}$. The domain $\mathcal{D}(\mathrm{N})\subseteq\mathcal{F}$ is dense and both $\tilde{a}_{t,s,k}$ and $\mathrm{U}_{t,s}\tilde{a}_{k}\mathrm{U}_{t,s}^{\ast}$ are bounded operators. As a consequence, using a standard continuity argument one deduces that (78) holds true for all $\psi\in\mathcal{F}$.
Proposition 3.6 expresses the fact that the isospectral flow defined via the unitary evolution system $\left(\mathrm{U}_{t,s}\right)_{s,t\in\mathbb{R}_{0}^{+}}$ is a time-dependent family of Bogoliubov $\ast$–automorphisms. In other words, this evolution system implements on the Fock space the family of Bogoliubov transformations
$U_{t,s}=\left(\begin{array}[]{cc}\mathbf{u}_{t,s}&\mathbf{v}_{t,s}\\ \overline{\mathbf{v}}_{t,s}&\overline{\mathbf{u}}_{t,s}\end{array}\right)\ ,\qquad s,t\in\mathbb{R}_{0}^{+}\ ,$
defined on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$. Compare Proposition 3.6 with Theorem 2.8.
Remark 3.7
There is a small inconsistency or typo in the first assertion of [6, Lemma 71] as one can see from its proof (which is correct). [6, Lemma 71] done in the bosonic case has be analogous to Proposition 3.6.
3.3 Solution to the Brockett-Wegner Flow on Quadratic Hamiltonians
The continuous flow of Bogoliubov transformations used here is based on an elliptic flow (191) on operators acting on the one-particle Hilbert space. This flow is not studied here, but in [24], some important results of which are shortly explained in Section A. Therefore, the kernel of our proofs on the diagonalization of quadratic Hamiltonians lies on the one-particle Hilbert space.
This being said, we had to implement on the Fock space the corresponding Bogoliubov transformations generating by the solution to the elliptic flow. This was done in Section 3.2 and it remains to shows that the Brockett-Wegner flow (30), applied to a fermionic quadratic Hamiltonian $\mathrm{H}_{0}$ and the particle number operator $\mathrm{N}$, has a solution as a unitary equivalent quadratic Hamiltonian. As a consequence, we have to deal here with domain issues on the Fock space, which can be sometimes cumbersome. However, as explained in Section 2.5, the heuristic idea is very easy to understand and thus mathematically elegant from our viewpoint.
In fact, our aim here is not only to obtain a diagonalization of quadratic Hamiltonians, but also to highlight the Brockett-Wegner flow [19, 20, 21, 22] by given another rigorous solution (and even its asymptotics) of this (usually) formal method for an unbounded case. In particular, trying to diagonalize quadratic Hamiltonians by adapting the method developed in [8] for the bosonic case may be efficient, but it would definitively suppress our second aim here, namely the study of the Brockett-Wegner flow and ODEs on operator spaces like the elliptic flow (191). Indeed, differential equations on non-commutative algebras of operators (like the Brockett-Wegner flow [19, 20, 21, 22]) are very little developed in Mathematics (in particular for infinite-dimensional spaces). Hence, any mathematical result on relevant differential equations for operators is useful and possibly liminal. This is one important objective of [24] and the present paper, the diagonalization of quadratic Hamiltonians under much more general conditions being one salient application achieved via asymptotics results of Section 3.5. Moreover, note that comparatively to [8] (done in the bosonic case), we obtain here some expressions for the diagonal form as integrals of elements of the elliptic flow, and in some cases a very explicit expressions (see (16) and (17)).
To this end, in this section we aim at proving that the Brockett-Wegner flow (30), applied to a fermionic quadratic Hamiltonian $\mathrm{H}_{0}$ and the particle number operator $\mathrm{N}$, has a solution as a unitary equivalent quadratic Hamiltonian $\mathrm{H}_{t}=\mathrm{U}_{t,s}\mathrm{\mathrm{H}}_{s}\mathrm{U}_{t,s}^{\ast}$ for any $s,t\in\mathbb{R}_{0}^{+}$, with the evolution system $\left(\mathrm{U}_{t,s}\right)_{s,t\in\mathbb{R}_{0}^{+}}$ being the family of unitary operators described in Theorem 3.5. This refers to the following assertion:
Theorem 3.8 (Solution to the Brockett-Wegner flow on quadratic Hamiltonians)
The quadratic Hamiltonians defined by (35) satisfy the equality
$\mathrm{H}_{t}=\mathrm{U}_{t,s}\mathrm{\mathrm{H}}_{s}\mathrm{U}_{t,s}^{\ast}\ ,\qquad s,t\in\mathbb{R}_{0}^{+},$
with $\left(\mathrm{U}_{t,s}\right)_{s,t\in\mathbb{R}_{0}^{+}}$ being the (evolution) family of unitary operators of Theorem 3.5 for the solution $D$ to the elliptic operator-valued flow (191) with initial condition $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$.
Proof. Note first that the solution $D$ to the elliptic operator-valued flow (191) with initial condition $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ satisfies $D\in C(\mathbb{R}_{0}^{+};\mathcal{L}^{2}(\mathfrak{h}))$ and $D_{t}=-D_{t}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}_{0}^{+}$, thanks to Theorem A.1. We can in particular apply Theorem 3.5 for this continuous family of Hilbert-Schmidt operators. Observe next that the theorem is equivalent to the assertion that
$\left(\mathrm{H}_{t}+i\lambda\mathbf{1}\right)^{-1}=\mathrm{U}_{t,s}\left(\mathrm{H}_{s}+i\lambda\mathbf{1}\right)^{-1}\mathrm{U}_{t,s}^{\ast}\ ,\qquad s,t\in\mathbb{R}_{0}^{+}\ ,$ (79)
for some $\0\lambda\in\mathbb{R}\backslash\{0\}$. The proof is the same as the one done in various lemmata of [6, Section VI.2], except one has to replace [6, Lemma 65] with Corollary 3.2. For the reader’s convenience, we give below the important steps, but without much details.
Step 1: For any $t\in\mathbb{R}_{0}^{+}$, one writes the operator
$(\mathrm{N}+\mathbf{1})(\mathrm{H}_{t}+i\lambda\mathbf{1})^{-1}(\mathrm{N}+\mathbf{1})^{-1}=\Big{(}\mathrm{H}_{t}-i\mathrm{G}_{t}(\mathrm{N}+\mathbf{1})^{-1}+i\lambda\mathbf{1}\Big{)}^{-1}$ (80)
as a norm convergent Neumann series (see [21, Equation (VI.100)]) for any $\lambda\in\mathbb{R}$ such that
$\left|\lambda\right|>C\|D_{t}\|_{2}\geq\|\mathrm{G}_{t}\left(\mathrm{N}+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}$ (81)
$C\in\mathbb{R}^{+}$ being some fixed, sufficiently large, time-independent constant (see (62)). It yields
$\|(\mathrm{N}+\mathbf{1})(\mathrm{H}_{t}+i\lambda\mathbf{1})^{-1}(\mathrm{N}+\mathbf{1})^{-1}\|_{\mathrm{op}}$$\leq$ $\frac{1}{|\lambda|-C\|D_{t}\|_{2}}$ (82)
$\|(\mathrm{N}+\mathbf{1})\mathrm{H}_{t}(\mathrm{H}_{t}+i\lambda\mathbf{1})^{-1}(\mathrm{N}+\mathbf{1})^{-1}\|_{\mathrm{op}}$$\leq$ $\frac{|\lambda|\left(1+C\|D_{t}\|_{2}\right)}{|\lambda|-C\|D_{t}\|_{2}}$ (83)
for any $t\in\mathbb{R}_{0}^{+}$ and $\lambda\in\mathbb{R}$ satisfying (81).
Step 2: By applying Corollary 3.2, one directly arrives at
$\|(\mathrm{H}_{t}-\mathrm{H}_{s})(\mathrm{N}+\mathbf{1)}^{-1}\|_{\mathrm{op}}\leq C\left(\int_{s\wedge t}^{s\vee t}\|D_{\tau}\|_{2}^{2}\mathrm{d}\tau+\|D_{t}-D_{s}\|_{2}\right)$ (84)
for some fixed (time-independent) constant $C\in\mathbb{R}^{+}$ and any $s,t\in\mathbb{R}_{0}^{+}$. Using carefully the resolvent identity together with Step 1 and Equation (84), one proves that
$t\mapsto\left(\mathrm{H}_{t}+i\lambda\mathbf{1}\right)^{-1}\in C\left(\left[s,T\right];\mathcal{B}(\mathcal{Y},\mathcal{F}_{b})\right)$ (85)
for any $s\in\mathbb{R}_{0}^{+}$, $T\in\left(s,\infty\right)$ and $\lambda\in\mathbb{R}$ satisfying
$\left|\lambda\right|>C\sup_{t\in\left[s,T\right]}\|D_{t}\|_{2}\ .$ (86)
In Equation (85), note that $\mathcal{Y}=\mathcal{D}\left(\mathrm{N}\right)$ is the dense domain of the particle number operator $\mathrm{N}$, see (60). In particular, the resolvent $t\mapsto\left(\mathrm{H}_{t}+i\lambda\mathbf{1}\right)^{-1}$ is strongly continuous on $\left[0,T\right]$ for any fixed $T\in\mathbb{R}^{+}$ and $\lambda\in\mathbb{R}$ satisfying (86) at $s=0$.
Step 3: Using Steps 1 and 2 with a cautious attention to domain issues, one proves that, for any non-zero initial time $s\in\mathbb{R}^{+}$, $T\in\left(s,\infty\right)$ and $\lambda\in\mathbb{R}$ satisfying (86), the resolvent $\left(\mathrm{H}_{t}+i\lambda\mathbf{1}\right)^{-1}$ satisfies, for any $\varphi,\psi\in\mathcal{D}\left(\mathrm{N}\right)$ and $t\in\left[s,T\right]$, the differential equation
$\langle\psi,(\partial_{t}Y_{t})\varphi\rangle=\langle\psi,(i\left[Y_{t},\mathrm{G}_{t}\right])\varphi\rangle\ ,\quad Y_{s}\doteq\left(\mathrm{H}_{s}+i\lambda\mathbf{1}\right)^{-1}\ .$
Note that Theorem A.1 and Corollary 3.2 are pivotal to arrive at this assertion.
Step 4: Now, using Steps 1–3, one shows that, for any non-zero initial time $s\in\mathbb{R}^{+}$, $T\in\left(s,\infty\right)$ and $\lambda\in\mathbb{R}$ satisfying (86),
$\langle\left(\mathrm{N}+\mathbf{1}\right)^{-1}\psi,\partial_{t}\{\mathrm{U}_{t,s}^{\ast}\left(\mathrm{H}_{t}+i\lambda\mathbf{1}\right)^{-1}\mathrm{U}_{t,s}\}\left(\mathrm{N}+\mathbf{1}\right)^{-1}\varphi\rangle=0$ (87)
for any $\varphi,\psi\in\mathcal{F}$ and all $t\in\left[s,T\right]$, i.e.,
$\langle\left(\mathrm{N}+\mathbf{1}\right)^{-1}\psi,(\mathrm{U}_{t,s}^{\ast}\left(\mathrm{H}_{t}+i\lambda\mathbf{1}\right)^{-1}\mathrm{U}_{t,s}-\left(\mathrm{H}_{s}+i\lambda\mathbf{1}\right)^{-1})\left(\mathrm{N}+\mathbf{1}\right)^{-1}\varphi\rangle=0\ .$ (88)
Since the domain $\mathcal{D}\left(\mathrm{N}\right)$ is dense and the resolvent $\left(\mathrm{H}_{t}+i\lambda\mathbf{1}\right)^{-1}$ as well as the unitary operator $\mathrm{U}_{t,s}$ are bounded, we infer from (88) that
$\mathrm{U}_{t,s}^{\ast}\left(\mathrm{H}_{t}+i\lambda\mathbf{1}\right)^{-1}\mathrm{U}_{t,s}=\left(\mathrm{H}_{s}+i\lambda\mathbf{1}\right)^{-1}$ (89)
for all $s\in\mathbb{R}^{+}$, $T\in\left(s,\infty\right)$, $t\in\left[s,T\right]$ and $\lambda\in\mathbb{R}$ satisfying (86). It remains to perform in the above equality the limit $s\rightarrow 0^{+}$, by using Theorems A.1 and 3.5 as well as the fact that $t\mapsto\left(\mathrm{H}_{t}+i\lambda\mathbf{1}\right)^{-1}$ is strongly continuous on $\left[0,T\right]$ for all $T\in\mathbb{R}^{+}$ and $\lambda\in\mathbb{R}$ satisfying (86) with $s=0$. This leads to the assertion.
3.4 Asymptotics of the Solution to the Brockett-Wegner Flow
Keeping in mind Brockett-Wegner’s strategy applied to the diagonalization of quadratic Hamiltonians, we study the limit $t\rightarrow\infty$ of the solution $\left(\mathrm{H}_{t}\right)_{t\in\mathbb{R}_{0}^{+}}$ to the Brockett-Wegner flow (30). See Equation (35) and Theorem 3.8. It is directly related to the square-integrability of the mapping $t\mapsto\|D_{t}\|_{2}$ on $\mathbb{R}_{0}^{+}$, where $D$ is solution to the elliptic operator flow (191). Sufficient conditions to ensure this feature are given by Theorem A.2.
The square-integrability of $\|D_{t}\|_{2}$ allows us to define the self-adjoint operator
$\Upsilon_{\infty}\doteq\Upsilon_{0}+16\int_{0}^{\infty}D_{\tau}D_{\tau}^{\ast}\mathrm{d}\tau$
with domain $\mathcal{D}\left(\Upsilon_{\infty}\right)=\mathcal{D}\left(\Upsilon_{0}\right)$ (the integral being absolutely convergent in the space $\mathcal{L}^{1}(\mathfrak{h})$ of trace-class operators) and therefore,
$\mathrm{H}_{\infty}\doteq\sum_{k,l\in\mathbb{N}}\left\{\Upsilon_{\infty}\right\}_{k,l}a_{k}^{\ast}a_{l}+\left(E_{0}-8\int_{0}^{\infty}\left\|D_{\tau}\right\|_{2}^{2}\mathrm{d}\tau\right)\mathbf{1}=d\Gamma\left(\Upsilon_{\infty}\right)+\left(E_{0}-8\int_{0}^{\infty}\left\|D_{\tau}\right\|_{2}^{2}\mathrm{d}\tau\right)\mathbf{1}\ ,$ (90)
on the domain $\mathcal{D}_{0}$ (5), keeping always in mind that $D$ is solution to the elliptic operator flow (191). See (35) extended to $t=\infty$. Note $\mathrm{H}_{\infty}\equiv\mathrm{H}_{\infty}^{\ast\ast}=\mathrm{H}_{\infty}^{\ast}$ is well-defined, for instance via Proposition 3.4. Up to a constant, it is the well-known second-quantization *$d\Gamma(\Upsilon_{\infty})$ (50) of $\Upsilon_{\infty}$ and its spectral properties can be directly obtained from the ones of the one-particle self-adjoint operator $\Upsilon_{\infty}$. In fact, $\mathrm{H}_{\infty}$ is $\mathrm{N}$–diagonal since it commutes with the particle number operator $\mathrm{N}$. Last but not least, it is the limit $t\rightarrow\infty$, in the strong resolvent sense, of the family $\left(\mathrm{H}_{t}\right)_{t\geq 0}$ of unitary equivalent quadratic Hamiltonians.
Proposition 3.9 (Quasi $\mathrm{N}$–diagonalization of quadratic Hamiltonians)
Assume the square-integrability [^20] This condition is clearly satisfied under the assumptions of Theorem A.2. of the mapping $t\mapsto\|D_{t}\|_{2}$ on $\mathbb{R}_{0}^{+}$, $D$ being solution to (191) with $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$. Then,
$\underset{t\rightarrow\infty}{\lim}\left\|(\mathrm{H}_{\infty}-\mathrm{H}_{t})(\mathrm{N}+\mathbf{1)}^{-1}\right\|_{\mathrm{op}}=0$ (91)
and for any non-zero real $\0\lambda\in\mathbb{R}\backslash\{0\}$ and vector $\varphi\in\mathcal{F}$,
$\underset{t\rightarrow\infty}{\lim}\left\|\{(\mathrm{H}_{\infty}+i\lambda\mathbf{1})^{-1}-(\mathrm{H}_{t}+i\lambda\mathbf{1})^{-1}\}\varphi\right\|_{\mathcal{F}}=0\ .$ (92)
Proof. Because of the square-integrability of the mapping $t\mapsto\|D_{t}\|_{2}$ on $\mathbb{R}_{0}^{+}$, we can extend (84) to all $t,s\in\mathbb{R}_{0}^{+}\cup\{\infty\}$ with $D_{\infty}\doteq 0$ and obviously deduce (91). To prove next that $\left(\mathrm{H}_{t}\right)_{t\geq 0}$ converges to $\mathrm{H}_{\infty}$ in the strong resolvent sense, observe first that, for any non-zero $\0\lambda\in\mathbb{R}\backslash\{0\}$ and $C\in\mathbb{R}^{+}$, there is $s_{\lambda}\in\mathbb{R}_{0}^{+}$ such that
$\left|\lambda\right|>C\sup_{t\in\left[s_{\lambda},\infty\right)}\|D_{t}\|_{2}\ ,$
because, by assumption, the Hilbert-Schmidt norm $\|D_{t}\|_{2}$ vanishes when $t\rightarrow\infty$. If the mapping $t\mapsto\|D_{t}\|_{2}$ is square-integrable on $\mathbb{R}_{0}^{+}$ then Equations (82)–(83) are satisfied for $t=\infty$ and sufficiently large $s\in\mathbb{R}_{0}^{+}$. As a consequence, using the resolvent identity, for any $\0\lambda\in\mathbb{R}\backslash\{0\}$, sufficiently large initial time $s\in\mathbb{R}_{0}^{+}$ and $t\in\left[s,\infty\right)$,
$\left\|\{(\mathrm{H}_{\infty}+i\lambda\mathbf{1})^{-1}-(\mathrm{H}_{t}+i\lambda\mathbf{1})^{-1}\}(\mathrm{N}+\mathbf{1})^{-1}\right\|_{\mathrm{op}}$ (93)
$\leq$ $\|\left(\mathrm{H}_{\infty}+i\lambda\mathbf{1}\right)^{-1}\|_{\mathrm{op}}\|\left(\mathrm{H}_{\infty}-\mathrm{H}_{t}\right)\left(\mathrm{N}+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}$
$\|\left(\mathrm{N}+\mathbf{1}\right)\left(\mathrm{H}_{t}+i\lambda\mathbf{1}\right)^{-1}\left(\mathrm{N}+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}\ ,$
which in turn yields the limit
$\underset{t\rightarrow\infty}{\lim}\left\|\{(\mathrm{H}_{\infty}+i\lambda\mathbf{1})^{-1}-(\mathrm{H}_{t}+i\lambda\mathbf{1})^{-1}\}(\mathrm{N}+\mathbf{1})^{-1}\right\|_{\mathrm{op}}=0$ (94)
for any $\0\lambda\in\mathbb{R}\backslash\{0\}$, thanks to (82)–(83) for $t=\infty$ and (91). The domain $\mathcal{D}(\mathrm{N})$ is dense in $\mathcal{F}$ and both resolvents $(\mathrm{H}_{\infty}+i\lambda\mathbf{1})^{-1}$ and $(\mathrm{H}_{t}+i\lambda\mathbf{1})^{-1}$ in (94) are bounded operators for any $t\in\mathbb{R}_{0}^{+}$ and for any fixed $\0\lambda\in\mathbb{R}\backslash\{0\}$. As a consequence, the strong resolvent convergence follows from (94) and a standard continuity argument.
By Equations (57)–(58), for any $t\in\mathbb{R}_{0}^{+}$, the non-$\mathrm{N}$–diagonal part [^21] Recall that Proposition 3.4 shows that $\mathrm{Z}_{t}$, $t\in\mathbb{R}_{0}^{+}$, as operators defined on the domain (5), are essentially self-adjoint and we use again the notation $\mathrm{Z}_{t}\equiv\mathrm{Z}_{t}^{\ast\ast}$, $t\in\mathbb{R}_{0}^{+}$, for their self-adjoint extension.
$\mathrm{Z}_{t}\doteq\sum_{k,l\in\mathbb{N}}\{D_{t}\}_{k,l}a_{k}^{\ast}a_{l}^{\ast}+\{\bar{D}_{t}\}_{k,l}a_{l}a_{k}$ (95)
of the quadratic operator $\mathrm{H}_{t}=\mathrm{U}_{t,0}\mathrm{H}_{0}\mathrm{U}_{t,0}^{\ast}$ originally defined by (35) satisfies $\mathcal{D}(\mathrm{Z}_{t})\supseteq\mathcal{D}(\mathrm{N})$ with
$\|\mathrm{Z}_{t}\left(\mathrm{N}+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}\leq(1+\sqrt{3})\|D_{t}\|_{2}\ ,\qquad t\in\mathbb{R}_{0}^{+}\ .$
Hence, if the Hilbert-Schmidt norm $\left\|D_{t}\right\|_{2}$ vanishes in the limit $t\rightarrow\infty$ then the non-$\mathrm{N}$–diagonal part $\mathrm{Z}_{t}$ of the quadratic operator $\mathrm{H}_{t}$ vanishes on the domain $\mathcal{D}(\mathrm{N})$, as $t\rightarrow\infty$. In this sense, Proposition 3.9 asserts a quasi $\mathrm{N}$–diagonalization of $\mathrm{H}_{0}$. In other words, performing the limit $t\rightarrow+\infty$ removes the non-$\mathrm{N}$–diagonal part of the Hamiltonian, but at this stage we have no information on the link between $\mathrm{H}_{\infty}$ and $\mathrm{H}_{0}$. This study is the aim of the next section.
3.5 Diagonalization of Quadratic Hamiltonians
Following Brockett-Wegner’s strategy – or Theorem 3.8 and Proposition 3.9 in our example – the diagonalization of quadratic Hamiltonians $\mathrm{H}_{0}$ (4) via a Bogoliubov $\mathbf{u}$$\mathbf{v}$ transformation should be achieved by performing the limits $t\rightarrow\infty$ of bounded operator families $\left(\mathrm{U}_{t,s}\right)_{t,s\in\mathbb{R}_{0}^{+}}$ (Theorems 3.5) and $\left(\mathbf{u}_{t,s}\right)_{t,s\in\mathbb{R}^{+}}$, $\left(\mathbf{v}_{t,s}\right)_{t,s\in\mathbb{R}_{0}^{+}}$ (see (66)–(67) and Proposition 3.6). Recall indeed that the limit $t\rightarrow\infty$ of the solution $\left(\mathrm{H}_{t}\right)_{t\in\mathbb{R}_{0}^{+}}$ to the Brockett-Wegner flow (30) leads to the $\mathrm{N}$–diagonal quadratic Hamiltonian $\mathrm{H}_{\infty}$ (90), thanks to Proposition 3.9. This refers to a quasi $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians. To get a true $\mathrm{N}$–diagonalization we prove that $\mathrm{H}_{\infty}$ is unitary equivalent to any $\mathrm{H}_{t}$, $t\in\mathbb{R}_{0}^{+}$.
To this end, we first show that $\left(\mathrm{U}_{t,s}\right)_{t,s\in\mathbb{R}_{0}^{+}}$ can be extended to infinite times $s,t=\{\infty\}$ by using now the integrability of the mapping $t\mapsto\left\|D_{t}\right\|_{2}$ on $\mathbb{R}_{0}^{+}$, instead of its square-integrability only, which is used in Proposition 3.9.
Theorem 3.10 (Extension of the unitary flow to infinite times)
Take $D\in C(\mathbb{R}_{0}^{+};\mathcal{L}^{2}(\mathfrak{h}))$ with $D_{t}=-D_{t}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}_{0}^{+}$ and assume the integrability of the mapping $t\mapsto\|D_{t}\|_{2}$ on $\mathbb{R}_{0}^{+}$. Then the family of Theorem 3.5 can be extended to a family $(\mathrm{U}_{t,s})_{s,t\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}}$ of unitary operators satisfying, on $\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}$, all properties (i)–(v) of Theorem 3.5.
Proof. From straightforward, albeit tedious, estimates using the triangle inequality, Equations (57), (62), Theorem 3.5 and Corollary 3.2, we obtain that, for any $s_{1},t_{1},s_{2},t_{2}\in\mathbb{R}_{0}^{+}$,
$\|(\mathrm{U}_{t_{2},s_{2}}-\mathrm{U}_{t_{1},s_{1}})(\mathrm{N}+\mathbf{1})^{-1}\|_{\mathrm{op}}$ (97)
$\leq$ $C\int_{s_{1}\wedge s_{2}}^{s_{1}\vee s_{2}}\|D_{\tau}\|_{2}\ \mathrm{d}\tau+C\int_{t_{1}\wedge t_{2}}^{t_{1}\vee t_{2}}\|D_{\tau}\|_{2}\mathrm{e}^{C\int_{s_{2}\wedge\tau}^{s_{2}\vee\tau}\|D_{\alpha}\|_{2}\mathrm{d}\alpha}\ \mathrm{d}\tau$
$=$ $C\int_{s_{1}\wedge s_{2}}^{s_{1}\vee s_{2}}\|D_{\tau}\|_{2}\ \mathrm{d}\tau+\mathrm{e}^{C\int_{s_{2}}^{\left(s_{2}\vee\left(t_{1}\wedge t_{2}\right)\right)\wedge\left(t_{1}\vee t_{2}\right)}\|D_{\tau}\|_{2}\mathrm{d}\tau}\left(\mathrm{e}^{C\int_{\left(s_{2}\vee\left(t_{1}\wedge t_{2}\right)\right)\wedge\left(t_{1}\vee t_{2}\right)}^{t_{1}\vee t_{2}}\|D_{\tau}\|_{2}\mathrm{d}\tau}-1\right)$
$+\mathrm{e}^{-C\int_{s_{2}}^{\left(s_{2}\vee\left(t_{1}\wedge t_{2}\right)\right)\wedge\left(t_{1}\vee t_{2}\right)}\|D_{\tau}\|_{2}\mathrm{d}\tau}\left(\mathrm{e}^{-C\int_{\left(s_{2}\vee\left(t_{1}\wedge t_{2}\right)\right)\wedge\left(t_{1}\vee t_{2}\right)}^{t_{1}\wedge t_{2}}\|D_{\tau}\|_{2}\mathrm{d}\tau}-1\right)$
for some time-independent constant $C\in\mathbb{R}^{+}$. In particular, if the mapping $t\mapsto\|D_{t}\|_{2}$ is integrable on $\mathbb{R}_{0}^{+}$ then the operator
$\{\mathrm{U}_{t_{2},s_{2}}-\mathrm{U}_{t_{1},s_{1}}\}(\mathrm{N}+\mathbf{1})^{-1}$
converges to zero in the norm topology either when $t_{1}\wedge t_{2}\rightarrow\infty$ and $s_{1}=s_{2}$ or when $s_{1}\wedge s_{2}\rightarrow\infty$ and $t_{1}=t_{2}$. Hence, since the domain $\mathcal{D}\left(\mathrm{N}\right)$ is dense with
$(\mathrm{U}_{t,s})_{s,t\in\mathbb{R}_{0}^{+}}\subseteq\mathcal{B}\left(\mathcal{F}\right)\ ,$ (98)
one infers from a continuity argument the existence of families $\left(\mathrm{U}_{\infty,s}\right)_{s\in\mathbb{R}_{0}^{+}}$ and $\left(\mathrm{U}_{t,\infty}\right)_{t\in\mathbb{R}_{0}^{+}}$ of bounded operators respectively defined, for any $\varphi\in\mathcal{F}$ and $s,t\in\mathbb{R}_{0}^{+}$, by the strong limits
$\mathrm{U}_{\infty,s}\varphi\doteq\underset{t\rightarrow\infty}{\lim}\left\{\mathrm{U}_{t,s}\varphi\right\}\ ,\quad\mathrm{U}_{t,\infty}\varphi\doteq\underset{s\rightarrow\infty}{\lim}\left\{\mathrm{U}_{t,s}\varphi\right\}\ .$ (99)
In particular, by the unitarity of the operator family (98),
$\max\left\{\left\|\mathrm{U}_{\infty,s}\right\|_{\mathrm{op}},\left\|\mathrm{U}_{t,\infty}\right\|_{\mathrm{op}}\right\}\leq 1$ (100)
for any $s,t\in\mathbb{R}_{0}^{+}$. From (97) we also obtain that, for any $\varphi\in\mathcal{F}$,
$\mathrm{U}_{\infty,\infty}\varphi\doteq\underset{t\rightarrow\infty}{\lim}\underset{s\rightarrow\infty}{\lim}\left\{\mathrm{U}_{t,s}\varphi\right\}=\underset{s\rightarrow\infty}{\lim}\underset{t\rightarrow\infty}{\lim}\left\{\mathrm{U}_{t,s}\varphi\right\}=\varphi\ .$ (101)
In particular, $\mathrm{U}_{\infty,\infty}=\mathbf{1}$. It is also straightforward to verify from (97) and (99) that $\left(\mathrm{U}_{\infty,s}\right)_{s\in\mathbb{R}_{0}^{+}}$ and $\left(\mathrm{U}_{t,\infty}\right)_{t\in\mathbb{R}_{0}^{+}}$ are strongly continuous in $s$ and $t$, respectively, while $\mathrm{U}_{t,s}=\mathrm{U}_{t,x}\mathrm{U}_{x,s}$ for any $x,t,s\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}$. In the same way, one easily checks that they are family of unitary operators with $\mathrm{U}_{\infty,s}^{\ast}=\mathrm{U}_{s,\infty}$ for any $s\in\mathbb{R}_{0}^{+}$. Moreover, using the triangle inequality, one gets that, for any $\0\epsilon\in\mathbb{R}\backslash\{0\}$ with sufficiently small $\left|\epsilon\right|>0$, all times $s,t\in\mathbb{R}^{+}$ and any vector $\varphi\in\mathcal{D}\left(\mathrm{N}^{2}\right)\subseteq\mathcal{D}\left(\mathrm{N}\right)\subseteq\mathcal{D}\left(\mathrm{G}_{s}\right)$,
$\|\{\epsilon^{-1}(\mathrm{U}_{\infty,s+\epsilon}-\mathrm{U}_{\infty,s})-i\mathrm{U}_{\infty,s}\mathrm{G}_{s}\}\varphi\|_{\mathcal{F}}$ (102)
$\leq$ $\|\{\mathrm{U}_{\infty,s}-\mathrm{U}_{t,s}\}\mathrm{G}_{s}\varphi\|_{\mathcal{F}}+\epsilon^{-1}\int_{s}^{s+\epsilon}\|\{\mathrm{G}_{\tau}-\mathrm{G}_{s}\}\varphi\|_{\mathcal{F}}\mathrm{d}\tau$
$+\epsilon^{-1}\int_{s}^{s+\epsilon}\|\{\mathrm{U}_{t,\tau}-\mathrm{U}_{t,s}\}\left(\mathrm{N}+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}\|\mathrm{G}_{s}\left(\mathrm{N}+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}\left\|\left(\mathrm{N}+\mathbf{1}\right)^{2}\varphi\right\|_{\mathcal{F}}\mathrm{d}\tau$
$+\epsilon^{-1}\int_{s}^{s+\epsilon}\|\{\mathrm{U}_{t,\tau}-\mathrm{U}_{t,s}\}\left(\mathrm{N}+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}\|\left[\mathrm{N},\mathrm{G}_{s}\right]\left(\mathrm{N}+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}\left\|\left(\mathrm{N}+\mathbf{1}\right)^{2}\varphi\right\|_{\mathcal{F}}\mathrm{d}\tau$
$+\epsilon^{-1}\|\{\mathrm{U}_{\infty,s}-\mathrm{U}_{t,s}\}\varphi\|_{\mathcal{F}}+\epsilon^{-1}\|\{\mathrm{U}_{\infty,s+\epsilon}-\mathrm{U}_{t,s+\epsilon}\}\varphi\|_{\mathcal{F}}\ .$
Since $D\in C(\mathbb{R}_{0}^{+};\mathcal{L}^{2}(\mathfrak{h}))$, by Corollary 3.2 and Equation (57),
$\mathrm{G}_{t}\in C(\mathbb{R}_{0}^{+};\mathcal{B}(\mathcal{Y},\mathcal{F}))\qquad\text{and}\qquad\left[\mathrm{N},\mathrm{G}_{t}\right]\left(\mathrm{N}+\mathbf{1}\right)^{-1}\in\mathcal{B}\left(\mathcal{F}\right)$ (103)
with $\mathcal{Y\doteq D}\left(\mathrm{N}\right)$, see, e.g., Equation (62). Therefore, we take $t=t(\epsilon)\rightarrow\infty$ as $\epsilon\rightarrow 0$ in (102) and use (97) together with standard arguments to get that $\partial_{s}\mathrm{U}_{\infty,s}\varphi=i\mathrm{U}_{\infty,s}\mathrm{G}_{s}\varphi$ for any $\varphi\in\mathcal{D}\left(\mathrm{N}^{2}\right)\subseteq\mathcal{D}\left(\mathrm{G}_{s}\right)$ and $s\in\mathbb{R}^{+}$. Since $\mathcal{D}\left(\mathrm{N}^{2}\right)$ is a core for the particle number operator $\mathrm{N}$, using a continuity argument together with (62), (97) and (100), we deduce that $\partial_{s}\mathrm{U}_{\infty,s}=i\mathrm{U}_{\infty,s}\mathrm{G}_{s}$ holds true on the domain $\mathcal{D}\left(\mathrm{N}\right)$, for any $s\in\mathbb{R}^{+}$. The case $s=0$ is done exactly in the same way by using only $\epsilon>0$.
We need now to prove that
$s\mapsto(\mathrm{N}+\mathbf{1})\mathrm{U}_{\infty,s}(\mathrm{N}+\mathbf{1})^{-1}\qquad\text{and}\qquad t\mapsto(\mathrm{N}+\mathbf{1})\mathrm{U}_{t,\infty}(\mathrm{N}+\mathbf{1})^{-1}$ (104)
are strongly continuous mappings from $\mathbb{R}_{0}^{+}$ to $\mathcal{B}\left(\mathcal{F}\right)$. First, in the same way one proves (62), we obtain from the CAR (6) and Corollary 3.2 that
$\|[(\mathrm{N}+\mathbf{1)}^{2},\mathrm{G}_{t}](\mathrm{N}+\mathbf{1)}^{-2}\|_{\mathrm{op}}\leq C\|D_{t}\|_{2}$
for some (time-independent) constant $C\in\mathbb{R}^{+}$. Similar to the proof of Theorem 3.5 (iv), one then finds that, for any $s,t\in\mathbb{R}_{0}^{+}$,
$\|(\mathrm{N}+\mathbf{1})^{2}\mathrm{U}_{t,s}(\mathrm{N}+\mathbf{1})^{-2}\|_{\mathrm{op}}\leq\mathrm{\exp}\left\{C\int_{s\wedge t}^{s\vee t}\|D_{\tau}\|_{2}\mathrm{d}\tau\right\}\ ,$ (105)
keeping in mind that $C\in\mathbb{R}^{+}$ is some fixed (time-independent) constant. Next, using again the triangle inequality, Equations (57), (62), (105), Theorem 3.5 and Corollary 3.2, like in Inequality (97), we get that, for any $s_{1},t_{1},s_{2},t_{2}\in\mathbb{R}_{0}^{+}$,
$\left\|(\mathrm{N}+\mathbf{1})(\mathrm{U}_{t_{2},s_{2}}-\mathrm{U}_{t_{1},s_{1}})(\mathrm{N}+\mathbf{1})^{-2}\right\|_{\mathrm{op}}$ (106)
$\leq$ $C\int_{s_{1}\wedge s_{2}}^{s_{1}\vee s_{2}}\left\|D_{\tau}\right\|_{2}\mathrm{e}^{C\int_{t_{1}\wedge\tau}^{t_{1}\vee\tau}\left\|D_{\alpha}\right\|_{2}\mathrm{d}\alpha}\mathrm{d}\tau+C\int_{t_{1}\wedge t_{2}}^{t_{1}\vee t_{2}}\left\|D_{\tau}\right\|_{2}\mathrm{e}^{C\int_{s_{2}\wedge\tau}^{s_{2}\vee\tau}\left\|D_{\alpha}\right\|_{2}\mathrm{d}\alpha}\mathrm{d}\tau$
for some time-independent constant $C\in\mathbb{R}^{+}$. Compare (106) with (97). In particular, an explicit expression for the upper bound of (106) can be obtained, similar to (97). The particle number operator $\mathrm{N}$ is self-adjoint and in particular a closed operator. Using this property and the last inequality together with a continuity argument and (99)–(101), it follows that, for any $\varphi\in\mathcal{F}$ and $s,t\in\mathbb{R}_{0}^{+}$,
$(\mathrm{N}+\mathbf{1})\mathrm{U}_{\infty,s}(\mathrm{N}+\mathbf{1})^{-1}\varphi$ $=$ $\underset{t\rightarrow\infty}{\lim}\left\{(\mathrm{N}+\mathbf{1})\mathrm{U}_{t,s}(\mathrm{N}+\mathbf{1})^{-1}\varphi\right\}\ ,$
$(\mathrm{N}+\mathbf{1})\mathrm{U}_{t,\infty}(\mathrm{N}+\mathbf{1})^{-1}\varphi$ $=$ $\underset{s\rightarrow\infty}{\lim}\left\{(\mathrm{N}+\mathbf{1})\mathrm{U}_{t,s}(\mathrm{N}+\mathbf{1})^{-1}\varphi\right\}\ ,$
as well as
$\underset{t\rightarrow\infty}{\lim}\underset{s\rightarrow\infty}{\lim}\left\{(\mathrm{N}+\mathbf{1})\mathrm{U}_{t,s}(\mathrm{N}+\mathbf{1})^{-1}\varphi\right\}=\underset{s\rightarrow\infty}{\lim}\underset{t\rightarrow\infty}{\lim}\left\{(\mathrm{N}+\mathbf{1})\mathrm{U}_{t,s}(\mathrm{N}+\mathbf{1})^{-1}\varphi\right\}\\ =\varphi=(\mathrm{N}+\mathbf{1})\mathrm{U}_{\infty,\infty}(\mathrm{N}+\mathbf{1})^{-1}\ .$
It is then straightforward to verify that the two mappings of Equations (104) are strongly continuous in $s$ and $t$, respectively. Moreover, using again the triangle inequality and Theorem 3.5, one finds that, for any $\epsilon\in\mathbb{R}$ with sufficiently small $\left|\epsilon\right|>0$, all times $t,s\in\mathbb{R}^{+}$ and any vector $\varphi\in\mathcal{D}\left(\mathrm{N}^{2}\right)$,
$\|\{\epsilon^{-1}(\mathrm{U}_{t+\epsilon,\infty}-\mathrm{U}_{t,\infty})+i\mathrm{G}_{t}\mathrm{U}_{t,\infty}\}\varphi\|_{\mathcal{F}}$ (107)
$\leq$ $\|\mathrm{G}_{t}\left(\mathrm{N}+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}\|\left(\mathrm{N}+\mathbf{1}\right)\{\mathrm{U}_{t,\infty}-\mathrm{U}_{t,s}\}\varphi\|_{\mathcal{F}}$
$+\epsilon^{-1}\int_{t}^{t+\epsilon}\|\{\mathrm{G}_{\tau}-\mathrm{G}_{t}\}\left(\mathrm{N}+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}\|\left(\mathrm{N}+\mathbf{1}\right)\mathrm{U}_{\tau,s}\left(\mathrm{N}+\mathbf{1}\right)^{-1}\varphi\|_{\mathcal{F}}\mathrm{d}\tau$
$+\epsilon^{-1}\int_{t}^{t+\epsilon}\|\mathrm{G}_{t}\left(\mathrm{N}+\mathbf{1}\right)^{-1}\|_{\mathrm{op}}\|\left(\mathrm{N}+\mathbf{1}\right)\left(\mathrm{U}_{\tau,s}-\mathrm{U}_{t,s}\right)\left(\mathrm{N}+\mathbf{1}\right)^{-1}\varphi\|_{\mathrm{op}}\mathrm{d}\tau$
$+\epsilon^{-1}\|\{\mathrm{U}_{t+\epsilon,\infty}-\mathrm{U}_{t+\epsilon,s}\}\varphi\|_{\mathcal{F}}+\epsilon^{-1}\|\{\mathrm{U}_{t,\infty}-\mathrm{U}_{t,s}\}\varphi\|_{\mathcal{F}}\ .$
Therefore, we take $s=s(\epsilon)\rightarrow\infty$ as $\epsilon\rightarrow 0$ in (107) and use (106) together with (62), Theorem A.1 and standard arguments to get that $\partial_{t}\mathrm{U}_{t,\infty}\varphi=-i\mathrm{G}_{t}\mathrm{U}_{t,\infty}\varphi$ for any $\varphi\in\mathcal{D}\left(\mathrm{N}^{2}\right)$. Using again a continuity argument together with (62), (97) and (100), we deduce that $\partial_{t}\mathrm{U}_{t,\infty}=-i\mathrm{G}_{t}\mathrm{U}_{t,\infty}$ holds true on the domain $\mathcal{D}\left(\mathrm{N}\right)$, for any $t\in\mathbb{R}^{+}$. The case $t=0$ is done exactly in the same way by using $\epsilon>0$.
Note that Theorem 3.10 for the bosonic case is only partially proven in [6], even though it can also be shown in this situation. Remark additionally that $n\in\mathbb{N}$ iterations of the formal computation (61) to compute the $n$-folds muticommutators $[\mathrm{N},[\mathrm{N},\cdots,[\mathrm{N},\mathrm{G}_{t}]\cdots]$ together with Corollary 3.2 yield
$\|[(\mathrm{N}+\mathbf{1)}^{n},\mathrm{G}_{t}](\mathrm{N}+\mathbf{1)}^{-n}\|_{\mathrm{op}}\leq C_{n}\|D_{t}\|_{2}\ ,\qquad t\in\mathbb{R}_{0}^{+}\ ,$
for some (time-independent) constant $C_{n}\in\mathbb{R}^{+}$. By using the same arguments of the proof of Theorem 3.10 one thus checks that, for any $n\in\mathbb{N}$ and $s,t\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}$,
$\|(\mathrm{N}+\mathbf{1})^{n}\mathrm{U}_{t,s}(\mathrm{N}+\mathbf{1})^{-n}\|_{\mathrm{op}}\leq\mathrm{\exp}\left\{C_{n}\int_{s\wedge t}^{s\vee t}\|D_{\tau}\|_{2}\mathrm{d}\tau\right\}\ ,$
and both
$s\mapsto(\mathrm{N}+\mathbf{1})^{n}\mathrm{U}_{t,s}(\mathrm{N}+\mathbf{1})^{-n},\quad t\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\},\quad\text{and}\quad t\mapsto(\mathrm{N}+\mathbf{1})^{n}\mathrm{U}_{t,s}(\mathrm{N}+\mathbf{1})^{-n},\quad s\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\},$
are strongly continuous mapping from $\mathbb{R}_{0}^{+}$ to $\mathcal{B}\left(\mathcal{F}\right)$. These properties can be used to compute higher order derivatives of the evolution system $(\mathrm{U}_{t,s})_{s,t\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}}$. Mutatis mutandis for the bosonic case studied in [6].
Theorem 3.10 means in particular that, as $t\rightarrow\infty$ (resp. $s\rightarrow\infty$), the unitary operator family $(\mathrm{U}_{t,s})_{s,t\in\mathbb{R}_{0}^{+}}$ strongly (pointwise) converges to a unitary operator family $\left(\mathrm{U}_{\infty,s}\right)_{s\in\mathbb{R}_{0}^{+}}$ (resp. $\left(\mathrm{U}_{t,\infty}\right)_{t\in\mathbb{R}_{0}^{+}}=\left(\mathrm{U}_{\infty,t}^{\ast}\right)_{t\in\mathbb{R}_{0}^{+}}$). The limit families are strongly continuous in $s\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}$ or $t\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}$ and they satisfy the non-autonomous evolution equations
$\forall s,t\in\mathbb{R}_{0}^{+}:\qquad\left\{\begin{array}[]{llll}\partial_{t}\mathrm{U}_{t,\infty}=-i\mathrm{G}_{t}\mathrm{U}_{t,\infty}&,&\mathrm{U}_{\infty,\infty}\doteq\mathbf{1}&,\\ \partial_{s}\mathrm{U}_{\infty,s}=i\mathrm{U}_{\infty,s}\mathrm{G}_{s}&,&\mathrm{U}_{\infty,\infty}\doteq\mathbf{1}&,\end{array}\right.$
on the domain $\mathcal{D}\left(\mathrm{N}\right)$ and the cocycle property $\mathrm{U}_{t,s}=\mathrm{U}_{t,x}\mathrm{U}_{x,s}$ for $x,t,s\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}$. What is more, the extended family of Theorem 3.10 still realizes a Bogoliubov $\mathbf{u}$$\mathbf{v}$ transformation and the solution $\left(\mathrm{H}_{t}\right)_{t\in\mathbb{R}_{0}^{+}}$, defined by (35), to the Brockett-Wegner flow (30) and its asymptotics $\mathrm{H}_{\infty}$, defined by (90) (see Proposition 3.9), are unitarily equivalent:
Corollary 3.11 (Implementation of non-autonomous Bogoliubov transformations)
Take $D\in C(\mathbb{R}_{0}^{+};\mathcal{L}^{2}(\mathfrak{h}))$ with $D_{t}=-D_{t}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}_{0}^{+}$ and assume the integrability of the mapping $t\mapsto\|D_{t}\|_{2}$ on $\mathbb{R}_{0}^{+}$. Then, Proposition 3.6 can be extended to all $s,t\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}$, where $\mathbf{u}_{\infty,\infty}\doteq\mathbf{1}$ and $\mathbf{v}_{\infty,\infty}\doteq 0$.
Proof. If the mapping $t\mapsto\|D_{t}\|_{2}$ is integrable on $\mathbb{R}_{0}^{+}$ then the operator families $(\mathbf{u}_{t,s}-\mathbf{1})_{s,t\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}}$ and $(\mathbf{v}_{t,s})_{s,t\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}}$, with the definition $\mathbf{u}_{\infty,\infty}\doteq\mathbf{1}$ and $\mathbf{v}_{\infty,\infty}\doteq 0$, are well-defined in $\mathcal{L}^{2}(\mathfrak{h})$ by (66)-(67) and continuous in the Hilbert-Schmidt topology. Using the same arguments of Proposition 3.6 one verifies (69)–(70) for all $s,t\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}$. Additionally, the bounded operators (72) can also be extended to $s,t\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}$. In particular, Proposition 3.6 and Theorem 3.10 together with the integrability of the mapping $t\mapsto\|D_{t}\|_{2}$ on $\mathbb{R}_{0}^{+}$ yield
$\mathrm{U}_{\infty,s}a\left(\varphi\right)\mathrm{U}_{\infty,s}^{\ast}=\lim_{t\rightarrow\infty}\mathrm{U}_{t,s}a\left(\varphi\right)\mathrm{U}_{t,s}^{\ast}=\lim_{t\rightarrow\infty}\left\{a(\mathbf{u}_{t,s}\varphi)+a^{\ast}(\mathbf{v}_{t,s}\overline{\varphi})\right\}=a(\mathbf{u}_{\infty,s}\varphi)+a^{\ast}(\mathbf{v}_{\infty,s}\overline{\varphi})$
for all $\varphi\in\mathcal{F}$ and $s,t\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}$, in the strong operator topology.
Corollary 3.12 ($\mathrm{N}$–diagonalization of quadratic Hamiltonians)
Let $D$ be the solution to (191) with initial condition $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$. Assume the integrability [^22] Again, this condition is clearly satisfied under the assumptions of Theorem A.2. of the mapping $t\mapsto\|D_{t}\|_{2}$ on $\mathbb{R}_{0}^{+}$. Then, quadratic Hamiltonians defined by (35) satisfy
$\mathrm{H}_{t}=\mathrm{U}_{t,s}\mathrm{\mathrm{H}}_{s}\mathrm{U}_{t,s}^{\ast}\ ,\qquad s,t\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}\ ,$
with $\left(\mathrm{U}_{t,s}\right)_{s,t\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}}$ being the (evolution) family of unitary operators of Theorem 3.10.
Proof. Combine Theorem 3.8, Proposition 3.9 and Theorem 3.10 with elementary estimates using the triangle inequality. We omit the details. Note only that the solution $D$ to the elliptic operator-valued flow (191) with initial condition $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ satisfies $D\in C(\mathbb{R}_{0}^{+};\mathcal{L}^{2}(\mathfrak{h}))$ and $D_{t}=-D_{t}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}_{0}^{+}$, thanks to Theorem A.1. We can in particular apply Theorems 3.5 and 3.10 for this continuous family of Hilbert-Schmidt operators.
Corollaries 3.11 and 3.12 mean in particular that
$\mathrm{H}_{\infty}=\mathrm{U}_{\infty,0}\mathrm{H}_{0}\mathrm{U}_{\infty,0}^{\ast}$
is the result of the $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians via unitary Bogoliubov $\mathbf{u}$$\mathbf{v}$ transformations. In contrast with Proposition 3.9 (which needs the square integrability of the mapping $t\mapsto\|D_{t}\|_{2}$), the integrability of the mapping $t\mapsto\|D_{t}\|_{2}$ is crucial here to get always the existence of the limit operators $\mathrm{U}_{\infty,0}$ and $\mathrm{U}_{\infty,0}^{\ast}$.
3.6 Shale-Stinespring-like Condition for Generators of Bogoliubov Transformations
As explained in Section 2.6, Bach, Lieb and Solovej [15] introduced a general definition of quadratic Hamiltonians on fermionic Fock spaces. This refers to Definition 2.10. It leads to a strongly continuous one-parameter group of Bogoliubov transformations, as explained in the proof of Lemma 2.12. In this section we study such kind of groups and discuss their generators in relation with the Shale-Stinespring condition (see Theorem 2.8).
We take $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ and recall that a Bogoliubov transformation is a unitary operator $U\in\mathcal{B}(\mathcal{H})$ such that $U\mathfrak{A=A}U$, as explained in Section 2.4. We thus start by relating these properties with the generator of a group of Bogoliubov transformations:
Lemma 3.13 (Self-dual operators and groups of Bogoliubov transformations)
Take any self-adjoint operator $h=h^{\ast}$ acting on $\mathcal{H}$. Then, $\mathrm{e}^{ith}$ is a Bogoliubov transformation for any $t\in\mathbb{R}$ iff $h=-\mathfrak{A}h\mathfrak{A}$ with domain $\mathcal{D}(h)=\mathfrak{A}(\mathcal{D}(h))\subseteq\mathcal{\mathcal{H}}$.
Proof. Let $h=h^{\ast}$ acting on $\mathcal{H}$. Recall that $\mathfrak{A}$ is an involution, i.e., $\mathfrak{A}^{2}=\mathbf{1}$, and it is in particular one-to-one. So, the family $(\mathfrak{A}\mathrm{e}^{ith}\mathfrak{A})_{t\in\mathbb{R}}$ is a strongly continuous one-parameter group of unitary operators on $\mathcal{H}$. By Stone’s theorem [44, Theorem 6.2], there is a unique self-adjoint operator $g=g^{\ast}$ such that $\mathfrak{A}\mathrm{e}^{ith}\mathfrak{A}=\mathrm{e}^{itg}$ for any $t\in\mathbb{R}$. Using the antiunitarity of $\mathfrak{A}$, $\mathfrak{A}^{2}=\mathbf{1}$ and the theory of strongly continuous groups, one easily checks that $g=-\mathfrak{A}h\mathfrak{A}$ with domain $\mathcal{D}(g)=\mathfrak{A}(\mathcal{D}(h))$. Now, if $\mathrm{e}^{ith}$ is a Bogoliubov transformation for any $t\in\mathbb{R}$ then
$\mathrm{e}^{ith}=\mathfrak{A}\mathrm{e}^{ith}\mathfrak{A}=\mathrm{e}^{itg}\ \qquad t\in\mathbb{R}\ ,$
and so, $h=g=-\mathfrak{A}h\mathfrak{A}$ with domain $\mathcal{D}(h)=\mathfrak{A}(\mathcal{D}(h))$, by unicity of the generator $g$ (see again [44, Theorem 6.2]). Conversely, if $h=-\mathfrak{A}h\mathfrak{A}$ with domain $\mathcal{D}(h)=\mathfrak{A}(\mathcal{D}(h))\subseteq\mathcal{\mathcal{H}}$ then
$\mathrm{e}^{ith}=\mathrm{e}^{-it\mathfrak{A}h\mathfrak{A}}=\mathfrak{A}\mathrm{e}^{ith}\mathfrak{A}\ ,\qquad t\in\mathbb{R}\ ,$
because $ig=-i\mathfrak{A}h\mathfrak{A}$ is the (uniquely defined) generator of the strongly continuous one-parameter group $(\mathfrak{A}\mathrm{e}^{ith}\mathfrak{A})_{t\in\mathbb{R}}$ of unitary operators, as explained above. In particular, $\mathrm{e}^{ith}$ is a Bogoliubov transformation for any $t\in\mathbb{R}$.
This lemma shows that a group of Bogoliubov transformations is constructed from a special class of self-adjoint operators defined on the self-dual Hilbert space, which are named self-dual Hamiltonians [33, Definition 2.7]: A self-dual Hamiltonian is a self-adjoint operator $h=h^{\ast}$ acting on $\mathcal{H}$ and satisfying $h=-\mathfrak{A}h\mathfrak{A}$ with domain $\mathcal{D}(h)=\mathfrak{A}(\mathcal{D}(h))\subseteq\mathcal{\mathcal{H}}$.
The next step is to represent groups of Bogoliubov transformations and their generator on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ as a $2\times 2$ matrix of operator-valued coefficients. As already explained after Lemma 2.12 in Section 2.6, such a block operator matrix
$A=\left(\begin{array}[]{cc}A_{1,1}&A_{1,2}\\ A_{2,1}&A_{2,2}\end{array}\right)\ ,\qquad A_{1,1},A_{1,2},A_{2,1},A_{2,2}\in\mathcal{L}\left(\mathfrak{h}\right)\ ,$
is trivial to obtain on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ when $A\in\mathcal{B}\left(\mathcal{H}\right)$ is a bounded (linear) operator. Here, $\mathcal{L}\left(\mathfrak{h}\right)$ denotes the set of linear operators acting on $\mathfrak{h}$. However, this is not obvious anymore for unbounded operators $A$ acting on $\mathcal{H}$, as for some generators of groups of Bogoliubov transformations. Below, using Araki’s study [18], we give a general result on groups of Bogoliubov transformations and their generator as block operator matrices on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$.
Lemma 3.14 (Representation of Bogoliubov transformations and their generator)
A group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations and its generator $ih$ acting on $\mathcal{H}$, $h$ being a self-adjoint operator, can be represented on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ as follows:
  1. (i)
    Bogoliubov transformations: There are two strongly continuous families $(\mathbf{u}_{t})_{t\in\mathbb{R}},(\mathbf{v}_{t})_{t\in\mathbb{R}}\in\mathcal{B}\left(\mathfrak{h}\right)$ of bounded operators such that
    $\mathrm{e}^{ith}=\left(\begin{array}[]{cc}\mathbf{u}_{t}&\mathbf{v}_{t}\\ \overline{\mathbf{v}}_{t}&\overline{\mathbf{u}}_{t}\end{array}\right)\ ,\qquad t\in\mathbb{R}\ .$
  2. (ii)
    Generators: If $\dim\ker h\in 2\mathbb{N}\cup\{\infty\}$ then there are a dense subspace $\mathcal{Y}\subseteq\mathfrak{h}$, a symmetric operator $\Upsilon_{0}$ defined on $\mathcal{Y}$ and another operator $D_{0}=-D_{0}^{\top}$ acting on $\mathcal{CY}$ such that
    $h=\frac{1}{2}\left(\begin{array}[]{cc}\Upsilon_{0}&2D_{0}\\ -2\overline{D}_{0}&-\Upsilon_{0}^{\top}\end{array}\right)\ ,\qquad\mathcal{D}\left(h\right)=\mathcal{Y}\oplus\mathcal{CY}\ ,$
    where $\mathcal{C}$ is the complex conjugation of Remark 2.1.
Proof. Any bounded operator $U$ acting on $\mathcal{H}$ can always be represented on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ by
$U=\left(\begin{array}[]{cc}P_{\mathfrak{h}}UP_{\mathfrak{h}}&P_{\mathfrak{h}}UP_{\mathfrak{h}}^{\perp}\\ P_{\mathfrak{h}}^{\perp}UP_{\mathfrak{h}}&P_{\mathfrak{h}}^{\perp}UP_{\mathfrak{h}}^{\perp}\end{array}\right)$ (108)
with $P_{\mathfrak{h}}$ being the orthogonal projection on $\mathfrak{h}\oplus\{0\}$ and $P_{\mathfrak{h}}^{\perp}\doteq\mathbf{1}-P_{\mathfrak{h}}$. (By a slight abuse of notation, we can use in an appropriate way the identification $\mathfrak{h}\oplus\{0\}\equiv\mathfrak{h}$ and $\{0\}\oplus\mathfrak{h}\equiv\mathfrak{h}$ to see $P_{\mathfrak{h}}UP_{\mathfrak{h}}$, $P_{\mathfrak{h}}UP_{\mathfrak{h}}^{\perp}$, $P_{\mathfrak{h}}^{\perp}UP_{\mathfrak{h}}$ and $P_{\mathfrak{h}}^{\perp}UP_{\mathfrak{h}}^{\perp}$ as four operators acting on $\mathfrak{h}$.) If $U$ is a Bogoliubov transformation, i.e., $UU^{\ast}=U^{\ast}U=\mathbf{1}$ and $U\mathfrak{A=A}U$ then, using (21) together with $\overline{X}\doteq\mathcal{C}X\mathcal{C}$ (Remark 2.1) one computes that the above representation can be reduced to
$U=\left(\begin{array}[]{cc}\mathbf{u}&\mathbf{v}\\ \overline{\mathbf{v}}&\overline{\mathbf{u}}\end{array}\right)\ ,\qquad\mathbf{u},\mathbf{v}\in\mathcal{B}\left(\mathfrak{h}\right)\ .$ (109)
Applied to the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations, we get Assertion (i). It remains to represent the self-adjoint Hamiltonian $h$ leading to the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$. A priori, one can represent $h$ on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ like (108) for $U=h$, but as explained after Equation (38), there are nontrivial domain issues. To have some control on the domain of the corresponding operators, we first use Lemma 3.13 to deduce that $h$ is a self-dual Hamiltonian, i.e., $h=h^{\ast}=-\mathfrak{A}h\mathfrak{A}$ with domain $\mathcal{D}(h)=\mathfrak{A}(\mathcal{D}(h))\subseteq\mathcal{\mathcal{H}}$. Secondly, recall that basis projections are orthogonal projections $P\in\mathcal{B}(\mathcal{H})$ satisfying $\mathfrak{A}P\mathfrak{A}=P^{\bot}\doteq\mathbf{1}_{\mathcal{H}}-P$. If $\dim\ker h\in 2\mathbb{N}\cup\{\infty\}$ and $h=-\mathfrak{A}h\mathfrak{A}$, then [18, Lemma 4.8 and its proof] leads to the existence of a pair of basis projections $P$ and $P^{\bot}\doteq\mathbf{1}-P$ with range $\mathfrak{h}_{P}$ and $\mathfrak{h}_{P^{\bot}}$, respectively, such that $h$ can be represented on $\mathcal{H}=\mathfrak{h}_{P}\oplus\mathfrak{h}_{P^{\bot}}$ by
$h=\left(\begin{array}[]{cc}\mathbb{A}&0\\ 0&-\overline{\mathbb{A}}\end{array}\right)=\left(\begin{array}[]{cc}\mathbb{A}&0\\ 0&-\mathbb{A}^{\top}\end{array}\right)\ ,$
where $\mathbb{A}=\mathbb{A}^{\ast}$ is a (possibility unbounded) self-adjoint operator with domain $\mathcal{D}\left(\mathbb{A}\right)\subseteq\mathfrak{h}_{P}$. Even in the finite dimension situation, one can use the identification $\mathfrak{h}_{P}\equiv\mathfrak{A}\mathfrak{h}_{P}=\mathfrak{h}_{P^{\bot}}$ and see $-\mathbb{A}^{\top}$ as an operator acting on $\mathfrak{h}_{P^{\bot}}$. We meanwhile observe from [18, Lemma 4.8] that
$\mathcal{D}(h)=\mathcal{D}(\mathbb{A})\oplus\mathcal{D}(\mathbb{A}^{\top})\ .$ (110)
In the above matrix representation of $h$, note that we used $X^{\top}\doteq\mathcal{C}X^{\ast}\mathcal{C}$ (Remark 2.1), which yields $\mathbb{A}^{\top}=\overline{\mathbb{A}}=\mathcal{C\mathbb{A}C}$. By [18, Lemma 3.6], there is a Bogoliubov transformation $U$ such that $h$ can be represented on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ via the following product
$h=U^{\ast}\left(\begin{array}[]{cc}\mathbb{A}&0\\ 0&-\mathbb{A}^{\top}\end{array}\right)U\ .$
In this case, by [18, Proof of Lemma 3.6] and $\mathcal{D}(h)=\mathfrak{A}(\mathcal{D}(h))$, one checks from (110) the existence of one subspace $\mathcal{Y}\subseteq\mathfrak{h}$ such that
$\mathcal{D}\left(h\right)=\mathcal{Y}\oplus\mathcal{CY}\ ,$ (111)
where $\oplus$ stands (as in all the paper) for the (Hilbert) direct sums of Hilbert (sub)spaces (and not the algebraic direct sum). In fact, using (109), $h$ can be represented on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ by the following block operator matrix:
$h$ $=$ $\left(\begin{array}[]{cc}\mathbf{u}^{\ast}&\mathbf{v}^{\top}\\ \mathbf{v}^{\ast}&\mathbf{u}^{\top}\end{array}\right)\left(\begin{array}[]{cc}\mathbb{A}&0\\ 0&-\mathbb{A}^{\top}\end{array}\right)\left(\begin{array}[]{cc}\mathbf{u}&\mathbf{v}\\ \overline{\mathbf{v}}&\overline{\mathbf{u}}\end{array}\right)$ (118)
$=$ $\left(\begin{array}[]{cc}\mathbf{u}^{\ast}&\mathbf{v}^{\top}\\ \mathbf{v}^{\ast}&\mathbf{u}^{\top}\end{array}\right)\left(\begin{array}[]{cc}\mathbb{A}\mathbf{u}&\mathbb{A}\mathbf{v}\\ -\mathbb{A}^{\top}\overline{\mathbf{v}}&-\mathbb{A}^{\top}\overline{\mathbf{u}}\end{array}\right)$ (123)
$=$ $\left(\begin{array}[]{cc}\mathbf{u}^{\ast}\mathbb{A}\mathbf{u}-\mathbf{v}^{\top}\mathbb{A}^{\top}\overline{\mathbf{v}}&\mathbf{u}^{\ast}\mathbb{A}\mathbf{v}-\mathbf{v}^{\top}\mathbb{A}^{\top}\overline{\mathbf{u}}\\ \mathbf{v}^{\ast}\mathbb{A}\mathbf{u}-\mathbf{u}^{\top}\mathbb{A}^{\top}\overline{\mathbf{v}}&\mathbf{v}^{\ast}\mathbb{A}\mathbf{v}-\mathbf{u}^{\top}\mathbb{A}^{\top}\overline{\mathbf{u}}\end{array}\right)$ (126)
and the subspace $\mathcal{Y}$ of Equation (111) can be defined by
$\mathcal{Y}\doteq\left\{\varphi\in\mathfrak{h}:\mathbf{u}\varphi\in\mathcal{D}(\mathbb{A}),\ \overline{\mathbf{v}}\varphi\in\mathcal{D}(\mathbb{A}^{\top})\right\}=\left\{\varphi\in\mathfrak{h}:\mathbf{u}\varphi,\mathbf{v}\varphi\in\mathcal{D}(\mathbb{A})\right\}\subseteq\mathfrak{h}\ .$
If $\mathcal{D}\left(h\right)$ is a dense set in $\mathcal{H}$ then $\mathcal{Y}$ must [^23] By (111) and density of $\mathcal{D}\left(h\right)\subseteq\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$, for any $\varepsilon>0$ and $\varphi\in\mathfrak{h}$, there are $\psi_{1}\in\mathcal{Y}$ and $\psi_{2}\in\mathcal{CY}$ such that $\|\psi_{1}-\varphi\|_{\mathfrak{h}}\leq\|\psi_{1}-\varphi\|_{\mathfrak{h}}+\|\psi_{2}-\varphi\|_{\mathfrak{h}}=\|\left(\psi_{1},\psi_{2}\right)-\left(\varphi,0\right)\|_{\mathcal{H}}<\varepsilon$. be a dense set of $\mathfrak{h}$. Define
$\Upsilon_{0}\doteq 2\left(\mathbf{u}^{\ast}\mathbb{A}\mathbf{u}-\mathbf{v}^{\top}\mathbb{A}^{\top}\overline{\mathbf{v}}\right)=P_{\mathfrak{h}}hP_{\mathfrak{h}}\qquad\text{and}\qquad D_{0}\doteq\mathbf{u}^{\ast}\mathbb{A}\mathbf{v}-\mathbf{v}^{\top}\mathbb{A}^{\top}\overline{\mathbf{u}}=P_{\mathfrak{h}}hP_{\mathfrak{h}}^{\perp}\ ,$
which are both well-defined operators acting on $\mathfrak{h}$, with domains $\mathcal{Y}$ and $\mathcal{CY}$ respectively. Note that $\Upsilon_{0}^{\top}\mathcal{C=C}\Upsilon_{0}$ and $\Upsilon_{0}^{\top}$ is therefore defined on $\mathcal{CY}$, like $D_{0}$. Using $X^{\top}\doteq\mathcal{C}X^{\ast}\mathcal{C}$, $\overline{X}\doteq\mathcal{C}X\mathcal{C}$ and $\overline{X}^{\top}=\overline{X^{\top}}=X^{\ast}$, we observe that $D_{0}=-D_{0}^{\top}$ on the subspace $\mathcal{CY}$ and $\Upsilon_{0}$ is clearly a symmetric operator on $\mathcal{Y}$. Then, we obtain Assertion (ii).
Note that Lemma 3.14 (ii) does not assert that every self-dual Hamiltonians is compatible when $\dim\ker h\in 2\mathbb{N}\cup\{\infty\}$. Indeed, it does not necessarily imply that the densely defined symmetric operator $\Upsilon_{0}$ is essentially self-adjoint, even if $h$ is itself self-adjoint. For instance, if $h$ is self-adjoint then $\ker\left(h\pm i\right)=\left\{0\right\}$, see [44, Proposition 3.8]. In particular, using elements $\left(\varphi,0\right),\left(0,\mathcal{C}\varphi\right)\in\mathcal{D}(h)$ for $\varphi\in\mathcal{Y}$ we only deduce that
$\ker\left(\Upsilon_{0}\pm i\right)\cap\ker\left(\overline{D}_{0}\right)=\ker\left(\Upsilon_{0}^{\top}\pm i\right)\cap\ker\left(D_{0}\right)=\left\{0\right\}\ .$
Meanwhile, $\Upsilon_{0}$ is essentially self-adjoint iff $\ker\left(\Upsilon_{0}^{\ast}\pm i\right)=\left\{0\right\}$, by [44, Proposition 3.8]. Since we have a priori no special control on the kernels $\ker\left(D_{0}\right)$ and $\ker\left(\Upsilon_{0}\pm i\right)$, the essential self-adjointness of $\Upsilon_{0}$ stays unclear, even when we have the self-adjointness of $h$ at our disposal. Then, one may try to make a self-adjoint extension of the densely defined symmetric operator $\Upsilon_{0}$ , provided the deficiency indices of $\Upsilon_{0}$ are identical [^24] If this is not the case, one may use larger Hilbert space than $\mathcal{H}$, by [44, Corollary 13.11]. (cf. [44, Theorem 13.10]), as explained after Equation (41) in Section 2.6.
An updated description of available methods to extend densely defined symmetric operators can be found in [44, Part VI]. Nevertheless, such methods are rather technical and further discussions on this subject simply go beyond the main objective of the paper, beside the fact that additional conditions on $h$ should probably be imposed. In fact, even Araki in [18, Theorems 5.4 and 5.6] uses essentially self-adjoint, self-dual operators $h$ which can be written as well-defined block operator matrices and from now we consider the following assumptions on the self-dual Hamiltonians used below:
Assumption 1 (Self-dual Hamiltonians as Block operator matrices)
$h$ is a self-dual Hamiltonian with domain $\mathcal{D}\left(h\right)=\mathcal{Y}\oplus\mathcal{CY}$, $\mathcal{Y}\subseteq\mathfrak{h}$ being a dense vector subspace, which can be represented on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ by the block operator matrix
$h\doteq\frac{1}{2}\left(\begin{array}[]{cc}\Upsilon_{0}&2D_{0}\\ 2D_{0}^{\ast}&-\Upsilon_{0}^{\top}\end{array}\right)=\frac{1}{2}\left(\begin{array}[]{cc}\Upsilon_{0}&2D_{0}\\ -2\overline{D}_{0}&-\Upsilon_{0}^{\top}\end{array}\right)=\frac{1}{2}\left(\begin{array}[]{cc}\Upsilon_{0}&2D_{0}\\ -2\overline{D}_{0}&-\overline{\Upsilon}_{0}\end{array}\right)\ ,$
where $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ and $D_{0}=-D_{0}^{\top}$ are two (possibly unbounded) operators with domains $\mathcal{D}\left(\Upsilon_{0}\right)\supseteq\mathcal{Y}$ and $\mathcal{D}\left(D_{0}\right)\supseteq\mathcal{C}\mathcal{Y}$. (Note that $\mathcal{D}(\Upsilon_{0}^{\top})=\mathcal{CD}(\Upsilon_{0})$ since $\Upsilon_{0}^{\top}=\overline{\Upsilon}_{0}=\mathcal{C}\Upsilon_{0}\mathcal{C}$.)
A group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations, represented by Lemma 3.14 (i) on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$, satisfies the Shale-Stinespring condition at some time $t\in\mathbb{R}$ whenever its off-diagonal operator $\mathbf{v}_{t}$ is Hilbert-Schmidt, i.e., $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$. See Theorem 2.8. Meanwhile, if $h$ satisfies Assumption 1, with off-diagonal operator $D_{0}$ acting on the Hilbert space $\mathfrak{h}$, then we would like to understand the relations between the Hilbert-Schmidt properties of $\mathbf{v}_{t}$ and $D_{0}$. To analyze this, it is very convenient to formulate the problem by means of the theory of non-autonomous evolution equations [47, 48, 49, 50, 51].
Lemma 3.15 (Interaction picture of groups of Bogoliubov transformations)
Take any self-adjoint operator $h$ satisfying Assumption 1 and use Lemmata 3.13 and 3.14 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations. Let $c\in\mathbb{R}^{+}\cup\{\infty\}$ and assume that
$\mathrm{e}^{it\Upsilon_{0}}\mathcal{Y}\subseteq\mathcal{Y}\ ,\qquad t\in\left(-c,c\right)\ ,$ (127)
with the operator family $(\overline{D}_{0}\mathrm{e}^{it\Upsilon_{0}})_{t\in(-c,c)}$ being strongly continuous on $\mathcal{Y}$. Then, the block operator matrix
$\left(\begin{array}[]{cc}\mathbf{x}_{t,s}&\mathbf{y}_{t,s}\\ \mathbf{\tilde{y}}_{t,s}&\mathbf{\tilde{x}}_{t,s}\end{array}\right)\doteq\left(\begin{array}[]{cc}\mathrm{e}^{-it\Upsilon_{0}/2}\mathbf{u}_{t-s}\mathrm{e}^{is\Upsilon_{0}/2}&\mathrm{e}^{-it\Upsilon_{0}/2}\mathbf{v}_{t-s}\mathrm{e}^{-is\Upsilon_{0}^{\top}/2}\\ \mathrm{e}^{it\Upsilon_{0}^{\top}/2}\overline{\mathbf{v}}_{t-s}\mathrm{e}^{is\Upsilon_{0}/2}&\mathrm{e}^{it\Upsilon_{0}^{\top}/2}\overline{\mathbf{u}}_{t-s}\mathrm{e}^{-is\Upsilon_{0}^{\top}/2}\end{array}\right)\ ,\qquad s,t\in\mathbb{R}\ ,$
on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ defines a strongly continuous two-parameter family $(V_{t,s})_{s,t\in\mathbb{R}}$ of unitary operators satisfying, for any $s,t\in(-c,c)$,
$\left(\begin{array}[]{cc}\partial_{t}\mathbf{x}_{t,s}&\partial_{t}\mathbf{y}_{t,s}\\ \partial_{t}\mathbf{\tilde{y}}_{t,s}&\partial_{t}\mathbf{\tilde{x}}_{t,s}\end{array}\right)=i\left(\begin{array}[]{cc}C_{t}\mathbf{\tilde{y}}_{t,s}&C_{t}\mathbf{\tilde{x}}_{t,s}\\ C_{t}^{\ast}\mathbf{x}_{t,s}&C_{t}^{\ast}\mathbf{y}_{t,s}\end{array}\right)\quad\text{and}\quad\left(\begin{array}[]{cc}\partial_{s}\mathbf{x}_{t,s}&\partial_{s}\mathbf{y}_{t,s}\\ \partial_{s}\mathbf{\tilde{y}}_{t,s}&\partial_{s}\mathbf{\tilde{x}}_{t,s}\end{array}\right)=-i\left(\begin{array}[]{cc}\mathbf{y}_{t,s}C_{s}^{\ast}&\mathbf{x}_{t,s}C_{s}\\ \mathbf{\tilde{x}}_{t,s}C_{s}^{\ast}&\mathbf{\tilde{y}}_{t,s}C_{s}\end{array}\right)$
on the dense set $\mathcal{D}\left(h\right)\subseteq\mathcal{H}$, where
$C_{t}\doteq\mathrm{e}^{-it\Upsilon_{0}/2}D_{0}\mathrm{e}^{-it\Upsilon_{0}^{\top}/2}=-C_{t}^{\top}\ ,\qquad t\in\left(-c,c\right)\ .$ (128)
Proof. Assume all conditions of the lemma. The proof is rather elementary. For pedagogical reasons, it is divided in two steps:
Step 1: Define on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ the self-adjoint operators
$\gamma_{0}\doteq\frac{1}{2}\left(\begin{array}[]{cc}\Upsilon_{0}&0\\ 0&-\Upsilon_{0}^{\top}\end{array}\right)=\gamma_{0}^{\ast}\qquad\text{and}\qquad d_{0}\doteq\left(\begin{array}[]{cc}0&D_{0}\\ D_{0}^{\ast}&0\end{array}\right)=\left(\begin{array}[]{cc}0&D_{0}\\ -\overline{D}_{0}&0\end{array}\right)=d_{0}^{\ast}\ ,$
the domains of which equal
$\mathcal{D}\left(\gamma_{0}\right)=\mathcal{D}\left(\Upsilon_{0}\right)\oplus\mathcal{D}\left(\Upsilon_{0}^{\top}\right)\qquad\text{and}\qquad\mathcal{D}\left(d_{0}\right)=\mathcal{D}\left(\overline{D}_{0}\right)\oplus\mathcal{D}\left(D_{0}\right)\ .$
Note from the self-adjointness of $\Upsilon_{0}$ that $\mathcal{D}\left(\Upsilon_{0}^{\top}\right)=\mathcal{C}\mathcal{D}\left(\Upsilon_{0}\right)$. Compare with (111). Their sum is the self-adjoint operator $h=\gamma_{0}+d_{0}$ on the domain
$\mathcal{D}\left(h\right)=\mathcal{Y}\oplus\mathcal{CY\subseteq D}\left(\gamma_{0}\right)\cap\mathcal{D}\left(d_{0}\right)=\mathcal{D}\left(\Upsilon_{0}\right)\cap\mathcal{D}\left(\overline{D}_{0}\right)\oplus\mathcal{D}\left(\Upsilon_{0}^{\top}\right)\cap\mathcal{D}\left(D_{0}\right)\ .$
Note that the family $(\mathcal{C}\mathrm{e}^{it\Upsilon_{0}^{\top}}\mathcal{C})_{t\in\mathbb{R}}$ is a strongly continuous one-parameter group of unitary operators on $\mathcal{H}$ and we deduce from the theory of strongly continuous unitary groups (in particular Stone’s theorem [44, Theorem 6.2]) together with the antiunitarity of $\mathcal{C}$, $\mathcal{C}^{2}=\mathbf{1}$ and $X^{\top}\doteq\mathcal{C}X^{\ast}\mathcal{C}$ (Remark 2.1) that
$\mathcal{C}\mathrm{e}^{it\Upsilon_{0}^{\top}}\mathcal{C}=\mathrm{e}^{-it\mathcal{C}\Upsilon_{0}^{\top}\mathcal{C}}=\mathrm{e}^{-it\Upsilon_{0}}\ ,\qquad t\in\mathbb{R}\ .$ (129)
By Condition (127) and recalling again that $\mathcal{C}$ is an involution, it follows that
$\mathrm{e}^{it\Upsilon_{0}^{\top}}\mathcal{CY}\subseteq\mathcal{CY}\ ,\qquad t\in\left(-c,c\right)\ ,$ (130)
and therefore,
$\mathrm{e}^{it\gamma_{0}}\mathcal{D}\left(h\right)\subseteq\mathcal{D}\left(\gamma_{0}\right)\cap\mathcal{D}\left(d_{0}\right)\ ,\qquad t\in\left(-c,c\right)\ .$ (131)
Meanwhile, $\overline{X}\doteq\mathcal{C}X\mathcal{C}$,
$\mathcal{C}\mathrm{e}^{it\Upsilon_{0}}\mathcal{C}=\mathrm{e}^{-it\mathcal{C}\Upsilon_{0}\mathcal{C}}=\mathrm{e}^{-it\Upsilon_{0}^{\top}}\ ,\qquad t\in\mathbb{R}\ .$ (132)
(cf. (129)) and $(\overline{D}_{0}\mathrm{e}^{it\Upsilon_{0}})_{t\in(-c,c)}$ is by assumption strongly continuous on $\mathcal{Y}$. Hence, the collection $(D_{0}\mathrm{e}^{it\Upsilon_{0}^{\top}})_{t\in(-c,c)}$ is a well-defined strongly continuous family of operators on $\mathcal{CY}$. As a consequence $(C_{t})_{t\in(-c,c)}$ is also a strongly continuous family of operators on $\mathcal{CY}$, because $(\mathrm{e}^{it\Upsilon_{0}})_{t\in\mathbb{R}}$ is a strongly continuous group of unitary operators as $\Upsilon_{0}=\Upsilon_{0}^{\ast}$. Note additionally from $D_{0}=-D_{0}^{\top}$, $\mathcal{C}^{2}=\mathbf{1}$, (129) and (132) that
$C_{t}^{\top}\doteq\mathcal{C}C_{t}^{\ast}\mathcal{C}=(\mathcal{C}\mathrm{e}^{it\Upsilon_{0}^{\top}/2}\mathcal{C})(\mathcal{C}D_{0}^{\ast}\mathcal{C})(\mathcal{C}\mathrm{e}^{it\Upsilon_{0}/2}\mathcal{C})=\mathrm{e}^{-it\Upsilon_{0}/2}(-D_{0})\mathrm{e}^{-it\Upsilon_{0}^{\top}/2}=-C_{t}$
for all times $t\in\mathbb{R}$.
Step 2: We are now in a position to use evolution equations to prove the lemma. We compute that, for any $s,t\in(-c,c)$,
$\partial_{t}\left\{\mathrm{e}^{-it\gamma_{0}}\mathrm{e}^{i\left(t-s\right)h}\mathrm{e}^{is\gamma_{0}}\right\}=i\left(\mathrm{e}^{-it\gamma_{0}}d_{0}\mathrm{e}^{it\gamma_{0}}\right)\left(\mathrm{e}^{-it\gamma_{0}}\mathrm{e}^{i\left(t-s\right)h}\mathrm{e}^{is\gamma_{0}}\right)$
as well as
$\partial_{s}\left\{\mathrm{e}^{-it\gamma_{0}}\mathrm{e}^{i\left(t-s\right)h}\mathrm{e}^{is\gamma_{0}}\right\}=\left(\mathrm{e}^{-it\gamma_{0}}\mathrm{e}^{i\left(t-s\right)h}\mathrm{e}^{is\gamma_{0}}\right)\left(-i\mathrm{e}^{-is\gamma_{0}}d_{0}\mathrm{e}^{is\gamma_{0}}\right)$
both in the strong sense on the dense set $\mathcal{Y}\oplus\mathcal{CY}\subseteq\mathcal{H}$. In other words, the family
$V_{t,s}\doteq\mathrm{e}^{-it\gamma_{0}}\mathrm{e}^{i\left(t-s\right)h}\mathrm{e}^{is\gamma_{0}}\ ,\qquad s,t\in\mathbb{R}\ ,$ (133)
of (uniformly) bounded operators is a strongly continuous two-parameter family of unitary operators solving the non-autonomous evolution equations
$\forall s,t\in(-c,c):\qquad\partial_{t}Z_{t,s}=id_{t}Z_{t,s}\ ,\qquad\partial_{s}Z_{t,s}=-iZ_{t,s}d_{s}\ ,\qquad Z_{s,s}=\mathbf{1}\ ,$ (134)
in the strong sense on the dense set $\mathcal{Y}\oplus\mathcal{CY}\subseteq\mathcal{H}$, where
$d_{t}\doteq\mathrm{e}^{-it\gamma_{0}}d_{0}\mathrm{e}^{it\gamma_{0}}=d_{t}^{\ast}\ ,\qquad t\in(-c,c)\ .$
This formulation of the strongly continuous one-parameter group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations via $(V_{t,s})_{s,t\in\mathbb{R}}$ is advantageous because the operator family $(d_{t})_{t\in(-c,c)}$ can be represented in a simple way on the Hilbert space $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$:
$d_{t}=\left(\begin{array}[]{cc}\mathrm{e}^{-it\Upsilon_{0}/2}&0\\ 0&\mathrm{e}^{it\Upsilon_{0}^{\top}/2}\end{array}\right)\left(\begin{array}[]{cc}0&D_{0}\\ D_{0}^{\ast}&0\end{array}\right)\left(\begin{array}[]{cc}\mathrm{e}^{it\Upsilon_{0}/2}&0\\ 0&\mathrm{e}^{-it\Upsilon_{0}^{\top}/2}\end{array}\right)=\left(\begin{array}[]{cc}0&C_{t}\\ C_{t}^{\ast}&0\end{array}\right)$ (135)
for each time $t\in\mathbb{R}$, where $C_{t}$ is the operator defined by (128). Note that $(C_{t})_{t\in(-c,c)}$ is a strongly continuous family of operators on $\mathcal{CY}$ and so, $(d_{t})_{t\in(-c,c)}$ is a strongly continuous family of operators on $\mathcal{D}\left(h\right)=\mathcal{Y}\oplus\mathcal{CY}$. The lemma is then obtained by writing all operators on $\mathcal{H}$ as well as the non-autonomous evolution equations in terms of block operator matrices. For instance, one computes from Lemma 3.14 (i) that
$V_{t,s}\doteq\left(\begin{array}[]{cc}\mathbf{x}_{t,s}&\mathbf{y}_{t,s}\\ \mathbf{\tilde{y}}_{t,s}&\mathbf{\tilde{x}}_{t,s}\end{array}\right)=\left(\begin{array}[]{cc}\mathrm{e}^{-it\Upsilon_{0}/2}\mathbf{u}_{t-s}\mathrm{e}^{is\Upsilon_{0}/2}&\mathrm{e}^{-it\Upsilon_{0}/2}\mathbf{v}_{t-s}\mathrm{e}^{-is\Upsilon_{0}^{\top}/2}\\ \mathrm{e}^{it\Upsilon_{0}^{\top}/2}\overline{\mathbf{v}}_{t-s}\mathrm{e}^{is\Upsilon_{0}/2}&\mathrm{e}^{it\Upsilon_{0}^{\top}/2}\overline{\mathbf{u}}_{t-s}\mathrm{e}^{-is\Upsilon_{0}^{\top}/2}\end{array}\right)$
for any $s,t\in\mathbb{R}$.
Observe that Assumption 1 and Condition (127) imply that $\mathcal{Y}$ must be a core for the self-adjoint operator $\Upsilon_{0}$, by [44, Proposition 6.3]. In other words, $\Upsilon_{0}$ restricted to the subspace $\mathcal{Y}\subseteq\mathcal{D}\left(\Upsilon_{0}\right)$ is essentially self-adjoint. Note also that Lemma 3.15 can always be applied whenever the off-diagonal operator-valued coefficient $\overline{D}_{0}$ is relatively bounded with respect of the diagonal one $\Upsilon_{0}$ (see (47)). This is a direct consequence of the following lemma:
Lemma 3.16 (Relative boundedness of off-diagonal coefficients)
Let $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ and $D_{0}$ be two (possibly unbounded) operators with domains $\mathcal{Y}\doteq\mathcal{D}\left(\Upsilon_{0}\right)$ and $\mathcal{D}\left(\overline{D}_{0}\right)\supseteq\mathcal{D}\left(\Upsilon_{0}\right)$. If
$\left\|\overline{D}_{0}\left(\Upsilon_{0}+i\mathbf{1}\right)^{-1}\right\|_{\mathrm{op}}<\infty\ ,$ (136)
then $\mathrm{e}^{it\Upsilon_{0}}\mathcal{Y}\subseteq\mathcal{Y}$ for any $t\in\mathbb{R}$ and the operator family $(\overline{D}_{0}\mathrm{e}^{it\Upsilon_{0}})_{t\in\mathbb{R}}$ is strongly continuous on the dense vector subspace $\mathcal{Y}\subseteq\mathfrak{h}$.
Proof. As is well-known, if $\Upsilon_{0}$ is a self-adjoint operator, then $\mathcal{Y\doteq D}\left(\Upsilon_{0}\right)$ is a dense vector subspace of $\mathfrak{h}$ which is preserved by $\mathrm{e}^{it\Upsilon_{0}}$, i.e. $\mathrm{e}^{it\Upsilon_{0}}\mathcal{Y}\subseteq\mathcal{Y}$ for any $t\in\mathbb{R}$ and $\mathrm{e}^{it\Upsilon_{0}}\Upsilon_{0}=\Upsilon_{0}\mathrm{e}^{it\Upsilon_{0}}$ on $\mathcal{Y}$. See for instance [31, Chapter II, Lemma 1.3 (ii)]. Moreover, for any $\varphi\in\mathcal{Y}$ and $s,t\in\mathbb{R}$,
$\left\|\overline{D}_{0}\left(\mathrm{e}^{it\Upsilon_{0}}-\mathrm{e}^{is\Upsilon_{0}}\right)\varphi\right\|_{\mathfrak{h}}\leq\left\|\overline{D}_{0}\left(\Upsilon_{0}+i\mathbf{1}\right)^{-1}\right\|_{\mathrm{op}}\left\|\left(\mathrm{e}^{it\Upsilon_{0}}-\mathrm{e}^{is\Upsilon_{0}}\right)\left(\Upsilon_{0}+i\mathbf{1}\right)\varphi\right\|_{\mathfrak{h}}\ .$
Therefore, if (136) holds true, then the operator family $(\overline{D}_{0}\mathrm{e}^{it\Upsilon_{0}})_{t\in\mathbb{R}}$ is strongly continuous on $\mathcal{Y}$, because $(\mathrm{e}^{it\Upsilon_{0}})_{t\in\mathbb{R}}$ is a strongly continuous group of unitary operators as $\Upsilon_{0}=\Upsilon_{0}^{\ast}$.
Remark 3.17
Under the same assumptions as Lemma 3.16 by replacing $\overline{D}_{0}$ with $D_{0}$, the family $(D_{0}\mathrm{e}^{it\Upsilon_{0}^{\top}})_{t\in\mathbb{R}}$ is strongly continuous on $\mathcal{C}\mathcal{Y}$.
Condition (136) is a very natural condition to consider in order to be able to define block operator matrices as at least closable operators, like in Assumption 1. See for instance [37, 43, 46]. The best situation is when the off-diagonal operator-valued coefficient is directly bounded, i.e., $D_{0}\in\mathcal{B}(\mathfrak{h})$. In the following lemma, we give a sufficient condition in terms of groups of Bogoliubov transformations to have their generator $ih$ with bounded off-diagonal term $D_{0}\in\mathcal{B}(\mathfrak{h})$:
Lemma 3.18 (Boundedness of off-diagonal coefficients)
Under conditions of Lemma 3.15, if
$\liminf_{s\rightarrow 0^{-}}s^{-1}\left\|\mathbf{v}_{s}\right\|_{\mathrm{op}}<\infty\ ,$ (137)
then $D_{0}$ can be extended by continuity to a bounded operator (again denoted by $D_{0}\in\mathcal{B}(\mathfrak{h})$, by a slight abuse of notation).
Proof. Using Lemma 3.15 and the triangle inequality, for any $\varphi\in\mathcal{C}\mathcal{Y}$ and $s\in\left(-c,0\right)$, we have
$\left\|D_{0}\varphi\right\|_{\mathfrak{h}}\leq\left|s\right|^{-1}\int_{s}^{0}\left\|\left(D_{0}-C_{\tau}\right)\varphi\right\|_{\mathfrak{h}}\mathrm{d}\tau+\left\|s^{-1}\int_{s}^{0}C_{\tau}\varphi\mathrm{d}\tau\right\|_{\mathfrak{h}}$
while
$\mathbf{y}_{0,s}\varphi=i\int_{s}^{0}\mathbf{x}_{0,\tau}C_{\tau}\varphi\mathrm{d}\tau\ .$
It follows that, for any $\varphi\in\mathcal{C}\mathcal{Y}$ and $s\in\left(-c,0\right)$,
$\left\|D_{0}\varphi\right\|_{\mathfrak{h}}\leq\left|s\right|^{-1}\left\|\mathbf{y}_{0,s}\right\|_{\mathrm{op}}\left\|\varphi\right\|_{\mathfrak{h}}+\left|s\right|^{-1}\int_{s}^{0}\left\|\left(D_{0}-C_{\tau}\right)\varphi\right\|_{\mathfrak{h}}\mathrm{d}\tau+\left|s\right|^{-1}\int_{s}^{0}\left\|\left(\mathbf{x}_{0,\tau}-\mathbf{1}\right)C_{\tau}\varphi\right\|_{\mathfrak{h}}\mathrm{d}\tau\ .$ (138)
Recall that $(C_{t})_{t\in(-c,c)}$ is a strongly continuous family of operators on $\mathcal{CY}$ with $C_{0}=D_{0}$. Therefore, for any $\varphi\in\mathcal{C}\mathcal{Y}$,
$\lim_{s\rightarrow 0^{-}}\left|s\right|^{-1}\int_{s}^{0}\left\|\left(D_{0}-C_{\tau}\right)\varphi\right\|_{\mathfrak{h}}\mathrm{d}\tau=0\ .$ (139)
Since $\mathrm{e}^{ith}$ is a Bogoliubov transformation for all times $t\in\mathbb{R}$, one checks that
$\mathbf{u}_{t}^{\ast}\mathbf{u}_{t}+\mathbf{v}_{t}^{\top}\overline{\mathbf{v}}_{t}=\mathbf{1}\ ,$ (140)
see for instance Equation (69). It follows that $\mathbf{u}_{t}^{\ast}\mathbf{u}_{t},\mathbf{v}_{t}^{\top}\overline{\mathbf{v}}_{t}\leq\mathbf{1}$ and therefore,
$\left\|\mathbf{u}_{t}\right\|_{\mathrm{op}},\left\|\mathbf{v}_{t}\right\|_{\mathrm{op}}\leq 1\ ,\qquad t\in\mathbb{R}\ .$ (141)
In particular, through Lemma 3.15, we deduce that
$\sup_{s,t\in\mathbb{R}}\left\|\mathbf{x}_{t,s}\right\|_{\mathrm{op}}=\sup_{s,t\in\mathbb{R}}\left\|\mathrm{e}^{-it\Upsilon_{0}/2}\mathbf{u}_{t-s}\mathrm{e}^{is\Upsilon_{0}/2}\right\|_{\mathrm{op}}\leq 1$
and, by the triangle inequality,
$\left|s\right|^{-1}\int_{s}^{0}\left\|\left(\mathbf{x}_{0,\tau}-\mathbf{1}\right)C_{\tau}\varphi\right\|_{\mathfrak{h}}\mathrm{d}\tau\leq 2\left|s\right|^{-1}\int_{s}^{0}\left\|\left(C_{\tau}-D_{0}\right)\varphi\right\|_{\mathfrak{h}}\mathrm{d}\tau+\left|s\right|^{-1}\int_{s}^{0}\left\|\left(\mathbf{x}_{0,\tau}-\mathbf{1}\right)D_{0}\varphi\right\|_{\mathfrak{h}}\mathrm{d}\tau\ .$ (142)
The family $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ and $(\mathrm{e}^{it\Upsilon_{0}})_{t\in\mathbb{R}}$ are both strongly continuous group of unitary operators. In particular, $(\mathbf{x}_{0,\tau}=\mathbf{u}_{-\tau}\mathrm{e}^{i\tau\Upsilon_{0}})_{t\in\mathbb{R}}$ is a strongly continuous family of bounded operators satisfying $\mathbf{x}_{0,0}=\mathbf{1}$, while $(C_{t})_{t\in(-c,c)}$ is a strongly continuous family of operators on $\mathcal{CY}$ with $C_{0}=D_{0}$. Therefore, we infer from (142) that, for any $\varphi\in\mathcal{C}\mathcal{Y}$,
$\lim_{s\rightarrow 0^{-}}\left|s\right|^{-1}\int_{s}^{0}\left\|\left(\mathbf{x}_{0,\tau}-\mathbf{1}\right)C_{\tau}\varphi\right\|_{\mathfrak{h}}\mathrm{d}\tau=0\ .$ (143)
Now, we combine (138) with the limits (139) and (143) to arrive at
$\left\|D_{0}\varphi\right\|_{\mathfrak{h}}\leq\left\|\varphi\right\|_{\mathfrak{h}}\liminf_{s\rightarrow 0^{-}}s^{-1}\left\|\mathbf{v}_{s}\right\|_{\mathrm{op}}$ (144)
for any $\varphi\in\mathcal{C}\mathcal{Y}$. Note that $\mathcal{Y}\subseteq\mathfrak{h}$ is by assumption a dense vector subspace and $\mathcal{C}^{2}=\mathbf{1}$. Therefore, $\mathcal{C}\mathcal{Y}$ is also a dense vector space and if (137) holds true then we deduce from (144) that $D_{0}$ can be extended by continuity to a bounded operator acting on $\mathfrak{h}$.
From now, we restrict our focus from (possibly) fully unbounded cases to bounded off-diagonal operator-valued coefficients, i.e., $D_{0}\in\mathcal{B}(\mathfrak{h})$. It is worth mentioning that the diagonal element $\Upsilon_{0}$ is still taken as a possibly unbounded operator. In this case, using very standard arguments with Lemmata 3.15 and 3.16, we easily write all operator-valued coefficients of groups of Bogoliubov transformations via Dyson series involving operator-valued coefficients of their generator.
Lemma 3.19 (Coefficients of Bogoliubov transformations as Dyson series)
Take any self-adjoint operator $h$ satisfying Assumption 1 with $D_{0}\in\mathcal{B}(\mathfrak{h})$ and use Lemmata 3.13 and 3.14 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations as
$\mathrm{e}^{ith}=\left(\begin{array}[]{cc}\mathbf{u}_{t}&\mathbf{v}_{t}\\ \overline{\mathbf{v}}_{t}&\overline{\mathbf{u}}_{t}\end{array}\right)\ ,\qquad t\in\mathbb{R}\ .$
Then, for any $t\in\mathbb{R}$,
$\mathrm{e}^{-it\Upsilon_{0}/2}\mathbf{u}_{t}$ $=$ $\mathbf{1}+\sum_{p=1}^{\infty}\left(-1\right)^{p}\int_{0}^{t}\mathrm{d}\tau_{1}\cdots\int_{0}^{\tau_{2p-1}}\mathrm{d}\tau_{2p}(C_{\tau_{1}}C_{\tau_{2}}^{\ast})\cdots(C_{\tau_{2p-1}}C_{\tau_{2p}}^{\ast})\ ,$
$\mathrm{e}^{-it\Upsilon_{0}/2}\mathbf{v}_{t}$ $=$ $i\int_{0}^{t}\mathrm{d}\tau C_{\tau}+i\sum_{p=1}^{\infty}\left(-1\right)^{p}\int_{0}^{t}\mathrm{d}\tau_{1}\cdots\int_{0}^{\tau_{2p}}\mathrm{d}\tau_{2p+1}C_{\tau_{1}}\left((C_{\tau_{2}}^{\ast}C_{\tau_{3}})\cdots(C_{\tau_{2p}}^{\ast}C_{\tau_{2p+1}})\right)\ ,$
the series being absolutely summable in the Hilbert space $\mathfrak{h}$, where $\left(C_{t}\right)_{t\in\mathbb{R}}\subseteq\mathcal{B}(\mathfrak{h})$ is the strongly continuous family of bounded operators defined by (128).
Proof. Lemma 3.15 can always be applied for any bounded $D_{0}\in\mathcal{B}(\mathfrak{h})$. In this situation, the parameter $c$ of Lemma 3.15 is $c=\infty$, as one can see from Lemma 3.16. For instance, under all the assumptions of the lemma, by Equation (135), $(d_{t})_{t\in\mathbb{R}}\subseteq\mathcal{B}(\mathcal{H})$ is a strongly continuous family of bounded operators, since $(\mathrm{e}^{it\Upsilon_{0}})_{t\in\mathbb{R}}$ and $(\mathrm{e}^{it\Upsilon_{0}^{\top}})_{t\in\mathbb{R}}$ are strongly continuous group of unitary operators, $\Upsilon_{0}$ and $\Upsilon_{0}^{\top}$ being self-adjoint operators. As is well-known, we can therefore define the Dyson series
$\tilde{V}_{t,s}\doteq\mathbf{1}+\sum_{n=1}^{\infty}i^{n}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{s}^{\tau_{n-1}}\mathrm{d}\tau_{n}d_{\tau_{1}}\cdots d_{\tau_{n}}\mathrm{\ },\qquad s,t\in\mathbb{R}\ .$ (145)
This series is of course well-defined for any $\varphi\in\mathcal{H}$ since $\tilde{V}_{t,s}\varphi$ becomes an absolutely summable series in the Hilbert space $\mathcal{H}$. In fact, $(\tilde{V}_{t,s})_{s,t\in\mathbb{R}}$ is another strongly continuous two-parameter operator family solving the non-autonomous evolution equations (134) and is therefore nothing else that $(V_{t,s})_{s,t\in\mathbb{R}}$, by uniqueness [^25] Compute that, for any $s,t\in\mathbb{R}$ and $\varphi\in\mathcal{H}$, $(\tilde{V}_{t,s}-V_{t,s})\varphi=\int_{s}^{t}\partial_{\tau}\{V_{t,\tau}\tilde{V}_{\tau,s}\}\varphi\ \mathrm{d}\tau=0$. of the solution to (134). This is completly standard. Using the representation (135) of the operator family $(d_{t})_{t\in\mathbb{R}}$ and
$V_{t,s}=\left(\begin{array}[]{cc}\mathbf{x}_{t,s}&\mathbf{y}_{t,s}\\ \mathbf{\tilde{y}}_{t,s}&\mathbf{\tilde{x}}_{t,s}\end{array}\right)=\tilde{V}_{t,s}\ ,\qquad s,t\in\mathbb{R}\ ,$
on the Hilbert space $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$, we compute from direct computations of the Dyson series (145) that
$\mathbf{x}_{t,s}$ $=$ $\mathbf{1}+\sum_{p=1}^{\infty}\left(-1\right)^{p}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{s}^{\tau_{2p-1}}\mathrm{d}\tau_{2p}(C_{\tau_{1}}C_{\tau_{2}}^{\ast})\cdots(C_{\tau_{2p-1}}C_{\tau_{2p}}^{\ast})$ (146)
$\mathbf{\tilde{x}}_{t,s}$ $=$ $\mathbf{1}+\sum_{p=1}^{\infty}\left(-1\right)^{p}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{s}^{\tau_{2p-1}}\mathrm{d}\tau_{2p}(C_{\tau_{1}}^{\ast}C_{\tau_{2}})\cdots(C_{\tau_{2p-1}}^{\ast}C_{\tau_{2p}})$ (147)
$\mathbf{y}_{t,s}$ $=$ $i\int_{s}^{t}\mathrm{d}\tau C_{\tau}+i\sum_{p=1}^{\infty}\left(-1\right)^{p}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{s}^{\tau_{2p}}\mathrm{d}\tau_{2p+1}C_{\tau_{1}}\left((C_{\tau_{2}}^{\ast}C_{\tau_{3}})\cdots(C_{\tau_{2p}}^{\ast}C_{\tau_{2p+1}})\right)$ (148)
$\mathbf{\tilde{y}}_{t,s}$ $=$ $i\int_{s}^{t}\mathrm{d}\tau C_{\tau}^{\ast}+i\sum_{p=1}^{\infty}\left(-1\right)^{p}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{s}^{\tau_{2p}}\mathrm{d}\tau_{2p+1}C_{\tau_{1}}^{\ast}\left((C_{\tau_{2}}C_{\tau_{3}}^{\ast})\cdots(C_{\tau_{2p}}C_{\tau_{2p+1}}^{\ast})\right)$ (149)
for all times $s,t\in\mathbb{R}$. Observe that $\mathbf{\tilde{x}}_{t,s}=\overline{\mathbf{x}}_{t,s}$ and $\mathbf{\tilde{y}}_{t,s}=\overline{\mathbf{y}}_{t,s}$ for any $s,t\in\mathbb{R}$. Since Lemma 3.15 shows that
$\left(\begin{array}[]{cc}\mathbf{x}_{t,s}&\mathbf{y}_{t,s}\\ \mathbf{\tilde{y}}_{t,s}&\mathbf{\tilde{x}}_{t,s}\end{array}\right)\doteq\left(\begin{array}[]{cc}\mathrm{e}^{-it\Upsilon_{0}/2}\mathbf{u}_{t-s}\mathrm{e}^{is\Upsilon_{0}/2}&\mathrm{e}^{-it\Upsilon_{0}/2}\mathbf{v}_{t-s}\mathrm{e}^{-is\Upsilon_{0}^{\top}/2}\\ \mathrm{e}^{it\Upsilon_{0}^{\top}/2}\overline{\mathbf{v}}_{t-s}\mathrm{e}^{is\Upsilon_{0}/2}&\mathrm{e}^{it\Upsilon_{0}^{\top}/2}\overline{\mathbf{u}}_{t-s}\mathrm{e}^{-is\Upsilon_{0}^{\top}/2}\end{array}\right)\ ,\qquad s,t\in\mathbb{R}\ ,$
we arrive at the two equalities of the lemma from (146) and (148) for $s=0$.
Lemma 3.19 is a key result to understand the relation between the Hilbert-Schmidt property of $\mathbf{v}_{t}$ and $D_{0}$ because, at the cost of time-dependent generators, it allows us to disregard in some sense the possibly unbounded diagonal element $\Upsilon_{0}$ in Assumption 1, which only appears via the strongly continuous one-parameter group $(\mathrm{e}^{it\Upsilon_{0}})_{t\in\mathbb{R}}$ of unitary operators. Using such a result, it is straightforward to prove that Hilbert-Schmidt off-diagonal coefficients $D_{0}$ lead to the Shale-Stinespring condition at all times:
Lemma 3.20 (Sufficient condition for the Shale-Stinespring condition)
Take any self-adjoint operator $h$ satisfying Assumption 1 and use Lemmata 3.13 and 3.14 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations. If $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ then $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}$.
Proof. If $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ is Hilbert-Schmidt then one can combine Proposition 3.4 and Lemma 2.7 with Theorem 2.8 to deduce that $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for all times $t\in\mathbb{R}$, proving in this way this lemma. However, these arguments are far from being natural. It is indeed somehow artificial to use the second quantization formalism in order to analyze a property only related to operators on the (dubbed) one-particle Hilbert space $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$. As a result, we present here another (direct) proof of this lemma that can be obtained from Lemma 3.19 via an explicit upper bound: If $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ then we deduce from Lemma 3.19 together with the triangle inequality and properties of the trace that
$\left\|\mathbf{v}_{t}\right\|_{2}\leq\sum_{p=0}^{\infty}\frac{\left(\left|t\right|\left\|D_{0}\right\|_{2}\right)^{2p+1}}{(2p+1)!}=\sinh\left(\left|t\right|\left\|D_{0}\right\|_{2}\right)<\infty$
for all times $t\in\mathbb{R}$.
We now study the converse of Lemma 3.20. In other words, we would like to know whether $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for some times $t\in\mathbb{R}$ can imply $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$. This problem is much more delicate. We first give a general result dependent on the continuous operator family $(\varkappa_{t})_{t\in\mathbb{R}}\subseteq\mathcal{B}(\mathcal{B}(\mathfrak{h}))$ defined by (48), that is, for each time $t\in\mathbb{R}$,
$\varkappa_{t}\left(A\right)\doteq\int_{0}^{t}\mathrm{e}^{-i\tau\Upsilon_{0}/2}A\mathrm{e}^{-i\tau\Upsilon_{0}^{\top}/2}\mathrm{d}\tau\ ,\qquad A\in\mathcal{B}(\mathfrak{h})\ .$ (150)
Note again that the integral is well-defined for any vector $\varphi\in\mathfrak{h}$ since $(\mathrm{e}^{it\Upsilon_{0}})_{t\in\mathbb{R}}$ and $(\mathrm{e}^{it\Upsilon_{0}^{\top}})_{t\in\mathbb{R}}$ are strongly continuous groups of unitary operators.
Lemma 3.21 (Hilbert-Schmidt upper bound of twisted off-diagonal elements)
Take any self-adjoint operator $h$ acting on $\mathcal{H}=\mathfrak{h}\oplus\mathfrak{h}$ satisfying Assumption 1 and use Lemmata 3.13 and 3.14 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations. If $D_{0}\in\mathcal{B}(\mathfrak{h})$ then, for any $t\in\mathbb{R}$,
$\left\|\varkappa_{t}\left(D_{0}\right)\right\|_{2}\leq\left\|\mathbf{v}_{t}\right\|_{2}+\frac{\left|t\right|^{2}}{2}\left\|D_{0}\right\|_{\mathrm{op}}^{2}\sup_{\tau\in\left[-\left|t\right|,\left|t\right|\right]}\left\|\mathbf{v}_{\tau}\right\|_{2}\ .$
Proof. Using Lemmata 3.15 and 3.16 with $D_{0}\in\mathcal{B}(\mathfrak{h})$ and $c=\infty$ as well as the mapping (150), we obtain that
$\mathbf{y}_{t,0}=i\varkappa_{t}\left(D_{0}\right)-\int_{0}^{t}\mathrm{d}\tau_{1}C_{\tau_{1}}\int_{0}^{\tau_{1}}\mathrm{d}\tau_{2}C_{\tau_{2}}^{\ast}\mathbf{y}_{\tau_{2},0}\ ,$ (151)
from which we deduce that
$\left\|\varkappa_{t}\left(D_{0}\right)\right\|_{2}\leq\left\|\mathbf{y}_{t,0}\right\|_{2}+\int_{0}^{\left|t\right|}\mathrm{d}\tau_{1}\int_{0}^{\tau_{1}}\mathrm{d}\tau_{2}\left\|C_{\tau_{1}}C_{\tau_{2}}^{\ast}\mathbf{y}_{\tau_{2},0}\right\|_{2}\leq\left\|\mathbf{y}_{t,0}\right\|_{2}+\frac{\left|t\right|^{2}}{2}\left\|D_{0}\right\|_{\mathrm{op}}^{2}\sup_{\tau\in\left[-\left|t\right|,\left|t\right|\right]}\left\|\mathbf{y}_{\tau,0}\right\|_{2}\ ,$
where $C_{t}$ is the operator defined by (128) for any $t\in\mathbb{R}$. Using now Lemma 3.15 (with $c=\infty$), more precisely,
$\mathbf{y}_{t,0}=\mathrm{e}^{-it\Upsilon_{0}/2}\mathbf{v}_{t}\ ,\qquad t\in\mathbb{R}\ ,$
one arrives at the desired upper bound.
Corollary 3.22 (Hilbert-Schmidt continuity of twisted off-diagonal elements)
Under the conditions of Lemma 3.21, if $(\mathbf{v}_{t})_{t\in\left[-c,c\right]}\subseteq\mathcal{L}^{2}(\mathfrak{h})$ for some time $c\in\mathbb{R}^{+}$ and is continuous in $\mathcal{L}^{2}(\mathfrak{h})$ at $t=0$, i.e.,
$\lim_{t\rightarrow 0}\left\|\mathbf{v}_{t}\right\|_{2}=0\ ,$
then there is $\delta\in\mathbb{R}^{+}$ such that $(\varkappa_{t}(D_{0}))_{t\in\left[-\delta,\delta\right]}\subseteq\mathcal{L}^{2}(\mathfrak{h})$ and this operator family is continuous in $\mathcal{L}^{2}(\mathfrak{h})$ at $t=0$, i.e.,
$\lim_{t\rightarrow 0}\left\|\varkappa_{t}\left(D_{0}\right)\right\|_{2}=0\ .$
Proof. This is an obvious consequence of Lemma 3.21. In particular, since $(\mathbf{v}_{t})_{t\in\left[-c,c\right]}\subseteq\mathcal{L}^{2}(\mathfrak{h})$ is continuous in $\mathcal{L}^{2}(\mathfrak{h})$ at $t=0$ with $\mathbf{v}_{0}=0$, there is $\delta\in\mathbb{R}^{+}$ such that
$\sup_{t\in\left[-\delta,\delta\right]}\left\|\mathbf{v}_{t}\right\|_{2}<+\infty\ .$
Moreover, for the same reason, given any $\varepsilon\in\mathbb{R}^{+}$ there exists $\tilde{\delta}\in\left[-\delta,\delta\right]$ such that
$\sup_{t\in[-\tilde{\delta},\tilde{\delta}]}\left\|\mathbf{v}_{t}\right\|_{2}\leq\frac{\varepsilon}{1+0.5\left\|D_{0}\right\|_{\mathrm{op}}^{2}\delta^{2}}<+\infty\ ,$
which in turns implies that $\left\|\varkappa_{t}(D_{0})\right\|_{2}\leq\varepsilon$ for any $t\in[-\tilde{\delta},\tilde{\delta}]$.
Corollary 3.23 (Hilbert-Schmidt property of off-diagonal elements)
Under the conditions of Lemma 3.21, if $(\mathbf{v}_{t})_{t\in\left[0,c\right]}\subseteq\mathcal{L}^{2}(\mathfrak{h})$ for some time $c\in\mathbb{R}^{+}$ and
$\liminf_{t\rightarrow 0^{+}}t^{-1}\left\|\mathbf{v}_{t}\right\|_{2}<\infty\ ,$ (152)
then $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$.
Proof. By Lemma 3.21, for any $t\in\mathbb{R}^{+}$,
$\left\|t^{-1}\varkappa_{t}\left(D_{0}\right)\right\|_{2}\leq t^{-1}\left\|\mathbf{v}_{t}\right\|_{2}+\frac{t}{2}\left\|D_{0}\right\|_{\mathrm{op}}^{2}\sup_{\tau\in\left[-\left|t\right|,\left|t\right|\right]}\left\|\mathbf{v}_{\tau}\right\|_{2}\ .$ (153)
Under the conditions of Corollary 3.23, note that $(\mathbf{v}_{t})_{t\in\left[0,c\right]}\subseteq\mathcal{L}^{2}(\mathfrak{h})$ is in particular continuous in $\mathcal{L}^{2}(\mathfrak{h})$ at $t=0$ along the subsequence $(t_{n})_{n\in\mathbb{N}}$ satisfying
$\lim_{n\rightarrow\infty}t_{n}^{-1}\left\|\mathbf{v}_{t_{n}}\right\|_{2}=\liminf_{t\rightarrow 0^{+}}t^{-1}\left\|\mathbf{v}_{t}\right\|_{2}<\infty$
and there is $N\in\mathbb{R}^{+}$ such that
$\sup_{n\in\{N,N+1,\ldots,\infty\}}\left\|\mathbf{v}_{t_{n}}\right\|_{2}<+\infty\ .$ (154)
Additionally, for any $\varphi\in\mathfrak{h}$ and $t\in\mathbb{R}^{+}$,
$\left\|D_{0}\varphi\right\|_{\mathfrak{h}}^{2}\leq 2\left\|\left(D_{0}-t^{-1}\varkappa_{t}\left(D_{0}\right)\right)\varphi\right\|_{\mathfrak{h}}^{2}+2\left\|t^{-1}\varkappa_{t}\left(D_{0}\right)\varphi\right\|_{\mathfrak{h}}^{2}\ ,$
by the triangle inequality and the inequality $\left(x+y\right)^{2}\leq 2x^{2}+2y^{2}$ for $x,y\in\mathbb{R}_{0}^{+}$. Since $(\mathrm{e}^{it\Upsilon_{0}})_{t\in\mathbb{R}}$ and $(\mathrm{e}^{it\Upsilon_{0}^{\top}})_{t\in\mathbb{R}}$ are strongly continuous groups of unitary operators and $D_{0}\in\mathcal{B}(\mathfrak{h})$, we deduce from the last inequality that, for any $\varphi\in\mathfrak{h}$,
$\left\|D_{0}\varphi\right\|_{\mathfrak{h}}^{2}\leq 2\liminf_{t\rightarrow 0^{+}}\left\|t^{-1}\varkappa_{t}\left(D_{0}\right)\varphi\right\|_{\mathfrak{h}}^{2}\ .$
As a consequence, taking any orthonormal basis $\left\{\psi_{k}\right\}_{k=1}^{\infty}\subseteq\mathfrak{h}$, for any $D_{0}\in\mathcal{B}(\mathfrak{h})$, we have to evaluate the limit $k_{0}\rightarrow\infty$ of the quantity
$\sum_{k=1}^{k_{0}}\left\|D_{0}\psi_{k}\right\|_{\mathfrak{h}}^{2}\leq 2\liminf_{t\rightarrow 0^{+}}\sum_{k=1}^{k_{0}}\left\|t^{-1}\varkappa_{t}\left(D_{0}\right)\psi_{k}\right\|_{\mathfrak{h}}^{2}\leq 2\liminf_{t\rightarrow 0^{+}}\left\|t^{-1}\varkappa_{t}\left(D_{0}\right)\right\|_{2}^{2}\ .$
By (153)–(154) it follows that, for any $\varepsilon\in\mathbb{R}^{+}$,
$\left\|D_{0}\right\|_{2}\leq\sqrt{2}\liminf_{t\rightarrow 0^{+}}\left\|t^{-1}\varkappa_{t}\left(D_{0}\right)\right\|_{2}\leq\sqrt{2}\liminf_{t\rightarrow 0^{+}}t^{-1}\left\|\mathbf{v}_{t}\right\|_{2}+\varepsilon\ .$
In particular, if (152) holds true then $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$.
The continuity properties expressed in Corollary 3.22 are satisfied for all groups $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations that can be implemented in the Fock space. This is proven in Proposition 3.26. However, the sufficient condition of Corollary 3.23 to get a Hilbert-Schmidt off-diagonal element $D_{0}$ is a priori unclear for such implementable groups of Bogoliubov transformations, except if the vacuum state of the fermionic Fock space belongs to the domain of the Hamiltonian implementing the group of Bogoliubov transformations. See Proposition 3.26 and Corollary 3.27.
In particular situations, Corollary 3.23 is not necessary. In fact, heuristically, under the conditions of Corollary 3.22, we should expect in some cases that
$\varkappa_{t}\left(D_{0}\right)\doteq\int_{0}^{t}\mathrm{e}^{-i\tau\Upsilon_{0}/2}D_{0}\mathrm{e}^{-i\tau\Upsilon_{0}^{\top}/2}\mathrm{d}\tau=\int_{0}^{t}C_{\tau}\mathrm{d}\tau=tD_{0}+\mathcal{O}\left(1\right)\ ,$
in the Hilbert-Schmidt topology, as $t\rightarrow 0$. For instance, $\mathcal{O}(1)=0$ when $D_{0}\mathrm{e}^{-i\tau\Upsilon_{0}^{\top}/2}=\mathrm{e}^{i\tau\Upsilon_{0}/2}D_{0}$ for any $\tau\in[0,t]$. In the following lemma, we give the simple case for which the operator $D_{0}$ intertwines [^26] We say that an operator $A$ intertwines between two operators $B$ and $C$ when $AB=CA$. between $\Upsilon_{0}$ and $\Upsilon_{0}^{\top}$ (resp. $-\Upsilon_{0}^{\top}$).
Lemma 3.24 (Hilbert-Schmidt property – commuting case)
Take any self-adjoint operator $h$ acting on $\mathcal{H}=\mathfrak{h}\oplus\mathfrak{h}$ satisfying Assumption 1 with $D_{0}\in\mathcal{B}(\mathfrak{h})$ and use Lemmata 3.13 and 3.14 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations.
  1. (i)
    Assume $D_{0}\Upsilon_{0}=-\Upsilon_{0}^{\top}D_{0}$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for positive times $t$ in a neighborhood of 00. Then, $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}$.
  2. (ii)
    Assume $D_{0}\Upsilon_{0}=\Upsilon_{0}^{\top}D_{0}$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for positive times $t$ in a neighborhood of 00. If additionally
    $2\sup_{t\in\left[0,c\right]}\left\|\mathrm{e}^{-2it\Upsilon_{0}}-\mathbf{1}\right\|_{\mathrm{op}}<1\ ,$ (155)
    for some $c\in\mathbb{R}^{+}$ then $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}$.
Proof. We separate the proof in three short steps, the first one being a preliminary study while Step 2 and 3 prove Assertions (i) and (ii), respectively.
Step 1: Assume all conditions of the lemma as well as $\Upsilon_{0}D_{0}=\pm D_{0}\Upsilon_{0}^{\top}$. We show here that $D_{0}$ also intertwines between the groups generated by $i\Upsilon_{0}$ and ±$\pm i\Upsilon_{0}^{\top}$. Since $D_{0}\in\mathcal{B}(\mathfrak{h})$, the equality $\Upsilon_{0}D_{0}=\pm D_{0}\Upsilon_{0}^{\top}$ means in particular that
$D_{0}\mathcal{D}\left(\Upsilon_{0}^{\top}\right)\subseteq\mathcal{D}\left(\Upsilon_{0}\right)\ .$
Then, observing that
$\partial_{t}\left\{\mathrm{e}^{it\Upsilon_{0}}D_{0}\mathrm{e}^{\mp it\Upsilon_{0}^{\top}}\right\}=\mathrm{e}^{it\Upsilon_{0}}\left(\Upsilon_{0}D_{0}\mp D_{0}\Upsilon_{0}^{\top}\right)\mathrm{e}^{\mp it\Upsilon_{0}^{\top}}=0$
in the strong sense on the domain $\mathcal{D}(\Upsilon_{0}^{\top})$, we deduce that
$\mathrm{e}^{it\Upsilon_{0}}D_{0}=D_{0}\mathrm{e}^{\pm it\Upsilon_{0}^{\top}}\ ,\qquad t\in\mathbb{R}\ .$ (156)
Note that we use above the density of the domain $\mathcal{D}(\Upsilon_{0}^{\top})$ and the unitarity of operators $\mathrm{e}^{it\Upsilon_{0}^{\top}}$ for $t\in\mathbb{R}$, $\Upsilon_{0}^{\top}$ being self-adjoint.
Step 2: Assume in this step that $D_{0}\Upsilon_{0}=-\Upsilon_{0}^{\top}D_{0}$. Then, we find from Lemma 3.19 and Equality (156) that
$\mathrm{e}^{-it\Upsilon_{0}/2}\mathbf{v}_{t}=itD_{0}\left(\mathbf{1}+R\left(t\right)\right)\ ,\qquad t\in\mathbb{R}_{0}^{+}\ ,$ (157)
with
$R\left(t\right)\doteq\sum_{p=1}^{\infty}\left(-1\right)^{p}\frac{t^{2p}(D_{0}^{\ast}D_{0})^{p}}{(2p+1)!}\ ,\qquad t\in\mathbb{R}_{0}^{+}\ .$ (158)
Since $D_{0}\in\mathcal{B}(\mathfrak{h})$ and $\lim_{t\rightarrow 0}\cosh(t)-1=0$, there is $\delta\in\mathbb{R}^{+}$ such that
$\max_{t\in\left[0,\delta\right]}\left\|R\left(t\right)\right\|_{\mathrm{op}}\leq\sum_{p=1}^{\infty}\frac{\delta^{2p}\left\|D_{0}\right\|_{\mathrm{op}}^{2p}}{(2p+1)!}\leq\cosh\left(\delta\left\|D_{0}\right\|_{\mathrm{op}}\right)-1\leq\frac{1}{4}\ .$ (159)
Therefore, we infer from (157)–(159) that
$\left\|\mathbf{v}_{t}\right\|_{2}^{2}=\left\|\mathrm{e}^{-it\Upsilon_{0}/2}\mathbf{v}_{t}\right\|_{2}^{2}\geq t^{2}\left(1-2\left\|R\left(t\right)\right\|_{\mathrm{op}}\right)^{2}\left\|D_{0}\right\|_{2}^{2}\geq\frac{t^{2}}{4}\left\|D_{0}\right\|_{2}^{2}$ (160)
for small times $t\in\left[0,\delta\right]$. In particular, if $D_{0}\Upsilon_{0}=-\Upsilon_{0}^{\top}D_{0}$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for positive times $t$ in a neighborhood of 00, then $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$, which, combined with Lemma 3.20, leads to Assertion (i).
Step 3: Assume now that $\Upsilon_{0}D_{0}=D_{0}\Upsilon_{0}^{\top}$. Then, again via Lemma 3.19 and Equality (156),
$\mathbf{y}_{t,0}=it\left(\mathbf{1}+S\left(t\right)\right)D_{0}\ ,\qquad t\in\mathbb{R}^{+}\ ,$ (161)
where, for any $t\in\mathbb{R}^{+}$,
$S\left(t\right)\doteq t^{-1}\int_{0}^{t}\mathrm{d}\tau\left(\mathrm{e}^{-i\tau\Upsilon_{0}}-\mathbf{1}\right)+t^{-1}\sum_{p=1}^{\infty}\left(-1\right)^{p}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{s}^{\tau_{2p}}\mathrm{d}\tau_{2p+1}\mathrm{e}^{i\sum_{k=1}^{2p+1}\left(-1\right)^{k}\tau_{k}\Upsilon_{0}}\left(D_{0}D_{0}^{\ast}\right)^{p}\ .$ (162)
We now use the following estimate
$\left\|S\left(t\right)\right\|_{\mathrm{op}}\leq\sup_{\tau\in\left[0,t\right]}\left\|\mathrm{e}^{-i\tau\Upsilon_{0}}-\mathbf{1}\right\|_{\mathrm{op}}+\sum_{p=1}^{\infty}\frac{t^{2p}\left\|D_{0}\right\|_{\mathrm{op}}^{2p}}{(2p+1)!}\leq\sup_{\tau\in\left[0,t\right]}\left\|\mathrm{e}^{-it\Upsilon_{0}}-\mathbf{1}\right\|_{\mathrm{op}}+\cosh\left(t\left\|D_{0}\right\|_{\mathrm{op}}\right)-1$
for any $t\in\mathbb{R}^{+}$ to show the existence of $\delta\in\mathbb{R}^{+}$ such that
$2\max_{t\in\left[0,\delta\right]}\left\|S\left(t\right)\right\|_{\mathrm{op}}<1\ ,$ (163)
provided that (155) holds true. By (161)–(163), we then get
$\left\|\mathbf{v}_{t}\right\|_{2}^{2}=\left\|\mathrm{e}^{-it\Upsilon_{0}/2}\mathbf{v}_{t}\right\|_{2}^{2}\geq t^{2}\left(1-2\max_{t\in\left[0,\delta\right]}\left\|S\left(t\right)\right\|_{\mathrm{op}}\right)^{2}\left\|D_{0}\right\|_{2}^{2}\ .$
In particular, if (155) holds true as well as $D_{0}\Upsilon_{0}=\Upsilon_{0}^{\top}D_{0}$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for positive times $t$ in a neighborhood of 00, then $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$, which, combined with Lemma 3.20, leads to Assertion (ii).
Remark 3.25
$\Upsilon_{0}\in\mathcal{B}(\mathfrak{h})$ implies the norm continuity of the group $(\mathrm{e}^{it\Upsilon_{0}})_{t\in\mathbb{R}}$ of unitary operators, see, e.g., [31, Chapter I, Proposition 3.5 and Theorem 3.7]. Therefore, Lemma 3.24 (ii) clearly holds also true when $\Upsilon_{0}\in\mathcal{B}(\mathfrak{h})$, $D_{0}\Upsilon_{0}=\Upsilon_{0}^{\top}D_{0}$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for positive times $t$ in a neighborhood of 00.
We conclude this section by proving that the continuity properties given by Corollary 3.22 are satisfied for all groups $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations that can be implemented in the fermionic Fock space $\mathcal{F}$. In fact, Bruneau and Derezinski prove in [7, Theorem 4.2] that a strongly continuous one-parameter symplectic group, or a group of Bogoliubov transformations in the bosonic case [^27] It corresponds to the framework presented in section 2.4, except we consider a CCR $C^{\ast}$ algebra instead of a CAR $C^{\ast}$ algebra, and the involution $\mathfrak{A}$ is replaced by an anti-involution, i.e. $\mathfrak{A}^{2}=-\mathfrak{A}$. see [7]., is implementable on the bosonic Fock space if and only if the off-diagonal part of this group, corresponding here to $\mathbf{v}_{t}$, is not only Hilbert-Schmidt for all times but also continuous at time $t=0$ in the Hilbert-Schmidt topology. We partially show this fact for the fermionic case in Proposition 3.26 (i) while Proposition 3.26 (ii) is a new result extending the previous one.
Proposition 3.26 (Hilbert-Schmidt continuity of the Shale-Stinespring condition)
Take any self-adjoint operator $h$ satisfying Assumption 1 and use Lemmata 3.13 and 3.14 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations. Assume that $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ can be implemented in the Fock space $\mathcal{F}$ via a strongly continuous group $(\mathrm{e}^{it\mathrm{H}/2})_{t\in\mathbb{R}}$, $\mathrm{H}=\mathrm{H}^{\ast}$ being a self-adjoint operator acting on $\mathcal{F}$, as in Definition 2.10. Then the following assertions hold true:
  1. (i)
    The Shale-Stinespring condition is continuous at $t=0$:
    $\lim_{t\rightarrow 0}\left\|\mathbf{v}_{t}\right\|_{2}=\lim_{t\rightarrow 0}\left\|\mathbf{1}-\mathbf{u}_{t}^{\ast}\mathbf{u}_{t}\right\|_{1}=0\ .$
  2. (ii)
    If the vacuum state $\Psi\doteq\left(1,0,\ldots\right)$ belongs to the domain $\mathcal{D}\left(\mathrm{H}\right)$ then
    $\limsup_{t\rightarrow 0}t^{-1}\left\|\mathbf{v}_{t}\right\|_{2}\in\mathbb{R}_{0}^{+}\ .$
Proof. Fix all conditions of the proposition. The proof is done in several steps. Note that the proof of Assertion (i) is similar to the one of [7, Theorem 4.2] done in the bosonic case, but the fermionic case is a bit more complicated.
Step 1: Recall that $\mathcal{A}$ is the universal (self-dual CAR) $C^{\ast}$-algebra $\mathcal{A}$ (19) generated by fields $\mathrm{B}(\psi)$, $\psi\in\mathcal{\mathcal{H}}$, and an unit $\mathfrak{1}$, while $\mathbf{\pi}_{\mathcal{F}}$ denotes its Fock space representation, as uniquely defined by Equation (22). Since $\mathrm{e}^{ith}$ is an implementable Bogoliubov transformation for all times, as defined in [16, Equations (6.4)–(6.5)], for any $t\in\mathbb{R}$, there is a unitary operator $\mathbb{U}_{t}$ acting on the fermionic Fock space $\mathcal{F}$ and satisfying, for any $t\in\mathbb{R}$,
$\mathbb{U}_{t}\left(\mathbf{\pi}_{\mathcal{F}}\left(A\right)\right)\mathbb{U}_{t}^{\ast}=\mathbf{\pi}_{\mathcal{F}}\circ\mathbf{\chi}_{\mathrm{e}^{ith}}\left(A\right)=\mathrm{e}^{it\mathrm{H}/2}\left(\mathbf{\pi}_{\mathcal{F}}\left(A\right)\right)\mathrm{e}^{-it\mathrm{H}/2}\ ,\qquad A\in\mathcal{A}\ ,$ (164)
as well as
$\mathbb{U}_{t}\Psi=\left\{\begin{array}[]{lll}\det\left(1+K_{t}^{\ast}K_{t}\right)^{-1/4}\exp\left(\mathbb{K}_{t}\right)\Psi&\text{if}&L_{t}=0\ ,\\ \det\left(1+K_{t}^{\ast}K_{t}\right)^{-1/4}\prod_{\ell=1}^{L_{t}}a^{\ast}\left(\psi_{\ell,t}\right)\exp\left(\mathbb{K}_{t}\right)\Psi&\text{if}&L_{t}\in\mathbb{N}\ ,\end{array}\right.$ (165)
where $\Psi\doteq\left(1,0,\ldots\right)\in\mathcal{F}$ is the vacuum state of the Fock space, $a\left(\varphi\right)$ is the usual creation operator associated with $\varphi\in\mathfrak{h}$ and acting on $\mathcal{F}$, $L_{t}\doteq\dim\ker(\mathbf{u}_{t}^{\ast})<\infty$, $\{\psi_{\ell,t}\}_{\ell=1}^{L_{t}}$ is an orthonormal basis of $\ker(\mathbf{u}_{t}^{\ast})$ when $L_{t}\in\mathbb{N}$,
$K_{t}=\mathbf{v}_{t}\overline{\mathbf{u}}_{t}^{-1}\in\mathcal{L}^{2}(\mathfrak{h})\ ,\qquad t\in\mathbb{R}\ ,$ (166)
and $\mathbb{K}_{t}$ is an operator formally written from the kernel $k_{t}\left(p,q\right)$ of the Hilbert-Schmidt operator $K_{t}^{\ast}K_{t}$ as
$\mathbb{K}_{t}\doteq\frac{1}{2}\int_{\mathcal{M}}\mathfrak{a}\left(p\right)\mathfrak{a}\left(q\right)k_{t}\left(p,q\right)a^{\ast}\left(p\right)a^{\ast}\left(q\right)$ (167)
keeping in mind that $\mathfrak{h}\doteq L^{2}(\mathcal{M})$ is assumed to be realized as a space of square-integrable (complex-valued) functions on a measure space $(\mathcal{M},\mathfrak{a})$. Equations (164)–(167) requires additional imputs to be well-defined: Here, the operator $\overline{\mathbf{u}}_{t}$ is not necessarly invertible. However, because of Lemma 3.14 (i) and the fact that $\mathrm{e}^{ith}$ is always a Bogoliubov transformation, one invokes Ruijsenaars’ arguments of [16] to show in the present case that the restriction of $\overline{\mathbf{u}}_{t}$ to $\ker(\overline{\mathbf{u}}_{t})^{\bot}$ has a bounded inverse from $\ker(\mathbf{u}_{t}^{\top})^{\bot}$ to $\ker(\overline{\mathbf{u}}_{t})^{\bot}$, which is by definition extended by setting it equal to zero on $\ker(\mathbf{u}_{t}^{\top})$. The result of this extension is denoted by $\overline{\mathbf{u}}_{t}^{-1}$. This is used for instance in Equation (166) and it shows that $K_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ and $\mathbb{K}_{t}$ can be well-defined. Note that $\ker(\overline{\mathbf{u}}_{t})$ and $\ker(\mathbf{u}_{t}^{\ast})$ have always finite dimension because
$L_{t}\doteq\dim\ker(\mathbf{u}_{t}^{\ast})=\dim\ker(\mathbf{u}_{t})=\dim\ker(\overline{\mathbf{u}}_{t})=\mathrm{tr}\left(\left.(\mathbf{1}-\mathbf{u}_{t}\mathbf{u}_{t}^{\ast})\right|_{\ker(\mathbf{u}_{t}^{\ast})}\right)$
thanks to the equalities $\ker(\overline{\mathbf{u}}_{t})=\mathcal{C}\ker(\mathbf{u}_{t})$ together with [16, Equation (6.1)]. So, by using (140), one obtains that
$L_{t}\leq\mathrm{tr}(\mathbf{1}-\mathbf{u}_{t}\mathbf{u}_{t}^{\ast})=\mathrm{tr}\left(\mathbf{v}_{t}^{\top}\overline{\mathbf{v}}_{t}\right)=\mathrm{tr}\left(\mathbf{v}_{t}^{\ast}\mathbf{v}_{t}\right)=\left\|\mathbf{v}_{t}\right\|_{2}^{2}<\infty\ .$ (168)
Step 2: Note that $\mathbb{U}_{t}$ is a priori different from $\mathrm{e}^{it\mathrm{H}/2}$. In fact, (164) means that
$\mathbb{U}_{t}^{\ast}\mathrm{e}^{it\mathrm{H}/2}\in\mathbf{\pi}_{\mathcal{F}}\left(\mathcal{A}\right)^{\prime}\doteq\left\{B\in\mathcal{B}\left(\mathcal{F}\right):\forall A\in\mathcal{A}\ ,\left[B,\mathbf{\pi}_{\mathcal{F}}\left(A\right)\right]=0\right\}\ .$ (169)
The Fock space representation $\mathbf{\pi}_{\mathcal{F}}$ is irreducible, see, e.g. [32, Definition 2.3.7] and [27, Proposition 5.2.2]. It follows from Schur’s lemma [52, Proposition 9.20] that
[…]
which, combined with (169), implies in turn the existence of constants $b_{t}\in\mathbb{C}$, $t\in{\mathbb{R}}$, such that
$\mathrm{e}^{it\mathrm{H}/2}=b_{t}\mathbb{U}_{t}\ ,\qquad t\in{\mathbb{R}}\ .$ (171)
Note that, for any $t\in{\mathbb{R}}$ and every normalized vector $\varphi\in\mathcal{F}$, i.e., vectors $\varphi$ in the Fock space $\mathcal{F}$ satisfying $\left\|\varphi\right\|_{\mathcal{F}}=1$,
$\left|b_{t}\right|=\left\|b_{t}\varphi\right\|_{\mathcal{F}}=\left\|\mathbb{U}_{t}^{\ast}\mathrm{e}^{it\mathrm{H}/2}\varphi\right\|_{\mathcal{F}}=\left\|\varphi\right\|_{\mathcal{F}}=1\ .$ (172)
Since $(\mathrm{e}^{it\mathrm{H}})_{t\in\mathbb{R}}$ is a strongly continous group of unitary operators, it follows from (171)–(172) that
$\Psi=\lim_{t\rightarrow 0}\mathrm{e}^{it\mathrm{H}/2}\Psi=\lim_{t\rightarrow 0}b_{t}\mathbb{U}_{t}\Psi\ ,$ (173)
keeping in mind that $\Psi\doteq\left(1,0,\ldots\right)\in\mathcal{F}$ is the vacuum state. Assume that, for any $\delta\in\mathbb{R}^{+}$ there is $t\in\left[-\delta,\delta\right]$ such that $L_{t}\geq 1$. Then, we deduce from (165) and (172) the existence of a subsequence $(t_{n})_{n\in\mathbb{N}}\subseteq\mathbb{R}$ converging to zero such that
$\left\langle\Psi,b_{t_{n}}\mathbb{U}_{t_{n}}\Psi\right\rangle_{\mathcal{F}}=0\ ,\qquad n\in\mathbb{N}\ .$
This is not possible because of (173). Therefore, there is $\delta\in\mathbb{R}^{+}$ such that $L_{t}=0$ for any $t\in\left[-\delta,\delta\right]$. It follows from (168) that $\overline{\mathbf{u}}_{t}$ is invertible for $t\in\left[-\delta,\delta\right]$ and from (165) that
$\left\langle\Psi,\mathbb{U}_{t}\Psi\right\rangle_{\mathfrak{h}}=\det\left(1+K_{t}^{\ast}K_{t}\right)^{-1/4}>0\ ,\qquad t\in\left[-\delta,\delta\right]\ .$ (174)
In particular, by Equation (166), $\mathbf{v}_{t}=K_{t}\overline{\mathbf{u}}_{t}$ and
$\left\|\mathbf{v}_{t}\right\|_{2}^{2}=\mathrm{tr}\left(\mathbf{v}_{t}\mathbf{v}_{t}^{\ast}\right)=\mathrm{tr}\left(K_{t}\overline{\mathbf{u}}_{t}\mathbf{u}_{t}^{\top}K_{t}^{\ast}\right)\leq\left\|K_{t}\right\|_{2}^{2}$ (175)
for any $t\in\left[-\delta,\delta\right]$, $\delta\in\mathbb{R}^{+}$ being a sufficient small but fixed parameter.
Step 3: By (171)–(172), the mappings $\iota_{t}\in\mathcal{B}\left(\mathcal{B}\left(\mathcal{F}\right)\right)$, $t\in\mathbb{R}$, defined by
$\iota_{t}\left(B\right)\doteq\mathbb{U}_{t}B\mathbb{U}_{t}^{\ast}=\mathrm{e}^{it\mathrm{H}/2}B\mathrm{e}^{-it\mathrm{H}/2}\ ,\qquad t\in{\mathbb{R}},\text{ }B\in\mathcal{B}\left(\mathcal{F}\right)\ ,$ (176)
form a weak continuous one-parameter group of automorphisms of $\mathcal{B}\left(\mathcal{F}\right)$. So, applying these automorphisms to the orthogonal projection $P_{\Psi}$ on the vacuum state $\Psi\doteq\left(1,0,\ldots\right)$, we deduce that the function
$t\mapsto f\left(t\right)\doteq\mathrm{tr}\left(P_{\Psi}\iota_{t}\left(P_{\Psi}\right)\right)=\left\langle\Psi,\iota_{t}\left(P_{\Psi}\right)\Psi\right\rangle_{\mathfrak{h}}$ (177)
is continuous on $\mathbb{R}$, with $\mathrm{tr}(\cdot)$ being the usual trace for operators, while using Equation (174) and (176), we get that, for any $t\in\left[-\delta,\delta\right]$,
$f\left(t\right)=\left\langle\Psi,\mathbb{U}_{t}\Psi\right\rangle_{\mathfrak{h}}\left\langle\Psi,\mathbb{U}_{t}^{\ast}\Psi\right\rangle_{\mathfrak{h}}=\det\left(1+K_{t}^{\ast}K_{t}\right)^{-1/2}=\mathrm{e}^{-\left(1/2\right)\mathrm{tr}\ln\left(1+K_{t}^{\ast}K_{t}\right)}$
and the function
$t\mapsto g\left(t\right)\doteq\mathrm{tr}\ln\left(1+K_{t}^{\ast}K_{t}\right)=-2\ln f\left(t\right)$ (178)
is thus continuous on $\left[-\delta,\delta\right]$. In particular, since $K_{0}=0$,
$\lim_{t\rightarrow 0}\mathrm{tr}\ln\left(1+K_{t}^{\ast}K_{t}\right)=0\ .$ (179)
Using the spectral theorem, observe that, for any time $t\in\left[-\delta,\delta\right]$, the spectrum of $K_{t}^{\ast}K_{t}$ is an ordered discrete set $\left\{\lambda_{k,t}\right\}_{k=1}^{\infty}\subseteq\mathbb{R}_{0}^{+}$ of positive eigenvalues satisfying $\lambda_{k,t}\leq\lambda_{q,t}$ when $k\geq q$ as well as
$g\left(t\right)=\sum_{k=1}^{\infty}\ln\left(1+\lambda_{k,t}\right)\ .$ (180)
It is worth recalling that $K_{t}^{\ast}K_{t}$ is here trace-class, i.e., $K_{t}\in\mathcal{L}^{2}(\mathfrak{h})$. Additionally, we conclude from (179) and $\lambda_{k,t}\geq 0$ that
$\lim_{t\rightarrow 0}\sup_{k\in\mathbb{N}}\lambda_{k,t}=0$
as well as
$\lim_{t\rightarrow 0}\left\|K_{t}\right\|_{2}^{2}=\lim_{t\rightarrow 0}\sum_{k=1}^{\infty}\lambda_{k,t}\leq 2\lim_{t\rightarrow 0}\sum_{k=1}^{\infty}\ln\left(1+\lambda_{k,t}\right)=0\ ,$ (181)
using additionally the inequality $x\leq 2\ln\left(1+x\right)$ for any $x\in\left[0,1\right]$. Since $\mathrm{e}^{ith}$ is a Bogoliubov transformation for all times one has (141). In particular $\left\|\mathbf{u}_{t}\right\|_{\mathrm{op}}\leq 1$ and one can now perform the limit $t\rightarrow 0$ in (175) while using (181) to deduce that
$\lim_{t\rightarrow 0}\left\|\mathbf{v}_{t}\right\|_{2}=0\ .$ (182)
Since $\mathrm{tr}(X^{\top})=\mathrm{tr}(X)$ (use, e.g., Remark 2.1, in particular $X^{\top}\doteq\mathcal{C}X^{\ast}\mathcal{C}$), we also deduce from (140) and (182) that
$\lim_{t\rightarrow 0}\left\|\mathbf{1}-\mathbf{u}_{t}^{\ast}\mathbf{u}_{t}\right\|_{1}=0\ .$ (183)
Assertion (i) is therefore proven and we now show Assertion (ii) of the proposition.
Step 4: Observe from (176)–(177) that, for any $t\in{\mathbb{R}}$,
$f\left(t\right)=\mathrm{tr}\left(P_{\Psi}\mathrm{e}^{it\mathrm{H}/2}P_{\Psi}\mathrm{e}^{-it\mathrm{H}/2}\right)=\left\langle\Psi,\mathrm{e}^{it\mathrm{H}/2}\Psi\right\rangle_{\mathfrak{h}}\left\langle\Psi,\mathrm{e}^{-it\mathrm{H}/2}\Psi\right\rangle_{\mathfrak{h}}\ ,$
from which we deduce that
$\partial_{t}f\left(t\right)=\frac{i}{2}\left(\left\langle\Psi,\mathrm{e}^{it\mathrm{H}/2}\mathrm{H}\Psi\right\rangle_{\mathfrak{h}}\left\langle\Psi,\mathrm{e}^{-it\mathrm{H}/2}\Psi\right\rangle_{\mathfrak{h}}-\left\langle\Psi,\mathrm{e}^{it\mathrm{H}/2}\Psi\right\rangle_{\mathfrak{h}}\left\langle\Psi,\mathrm{e}^{-it\mathrm{H}/2}\mathrm{H}\Psi\right\rangle_{\mathfrak{h}}\right)$
as well as
$\partial_{t}^{2}f\left(t\right)=\frac{1}{2}\left(\left\langle\Psi,\mathrm{e}^{it\mathrm{H}/2}\mathrm{H}\Psi\right\rangle_{\mathfrak{h}}\left\langle\Psi,\mathrm{e}^{-it\mathrm{H}/2}\mathrm{H}\Psi\right\rangle_{\mathfrak{h}}-\mathfrak{Re}\left(\left\langle\mathrm{H}\Psi,\mathrm{e}^{it\mathrm{H}/2}\mathrm{H}\Psi\right\rangle_{\mathfrak{h}}\left\langle\Psi,\mathrm{e}^{-it\mathrm{H}/2}\Psi\right\rangle_{\mathfrak{h}}\right)\right)\ ,$
as soon as $\Psi\in\mathcal{D}\left(\mathrm{H}\right)$. In particular, $\partial_{t}f\left(0\right)=0$ and
$\partial_{t}^{2}f\left(0\right)=\frac{1}{2}\left(\left\langle\Psi,\mathrm{H}\Psi\right\rangle_{\mathfrak{h}}^{2}-\left\langle\mathrm{H}\Psi,\mathrm{H}\Psi\right\rangle_{\mathfrak{h}}\right)\leq 0\ .$
($-\partial_{t}^{2}f\left(0\right)$ is in fact the quantum fluctuation of the energy observable $\mathrm{H}$ in the state $\left\langle\Psi,\left(\cdot\right)\Psi\right\rangle_{\mathfrak{h}}$, which is positive by the Cauchy-Schwarz inequality.) As $f\left(t\right)>0$ at least for all $t\in\left[-\delta,\delta\right]$, it follows from (178) that, for any $t\in\left[-\delta,\delta\right]$,
$\partial_{t}g\left(t\right)=-2\frac{\partial_{t}f\left(t\right)}{f\left(t\right)}\qquad\text{and}\qquad\partial_{t}^{2}g\left(t\right)=2\frac{\left(\partial_{t}f\left(t\right)\right)^{2}}{f\left(t\right)^{2}}-2\frac{\partial_{t}^{2}f\left(t\right)}{f\left(t\right)}\ .$
In particular, since $f\left(0\right)=1$ and $\partial_{t}f\left(0\right)=0$,
$g\left(0\right)=0\ ,\quad\partial_{t}g\left(0\right)=0\quad\text{and}\quad\partial_{t}^{2}g\left(0\right)=-2\partial_{t}^{2}f\left(0\right)\ .$
Using Taylor’s theorem, we conclude that
$\lim_{t\rightarrow 0}\left|t^{-2}g\left(t\right)\right|=\partial_{t}^{2}g\left(0\right)=\left\langle\mathrm{H}\Psi,\mathrm{H}\Psi\right\rangle_{\mathfrak{h}}-\left\langle\Psi,\mathrm{H}\Psi\right\rangle_{\mathfrak{h}}^{2}\in\mathbb{R}_{0}^{+}$
when $\Psi\in\mathcal{D}\left(\mathrm{H}\right)$. By (180), it implies in this case that
$\lim_{t\rightarrow 0}\left|t^{-2}g\left(t\right)\right|=\lim_{t\rightarrow 0}\sum_{k=1}^{\infty}t^{-2}\ln\left(1+\lambda_{k,t}\right)\in\mathbb{R}_{0}^{+}\ .$
In the same way we get (181), we deduce from this last equality that
$\limsup_{t\rightarrow 0}t^{-1}\left\|K_{t}\right\|_{2}\in\mathbb{R}_{0}^{+}\ .$
Combined with (141) and (175), this last inequality yields Assertion (ii).
Corollary 3.27 (Hilbert-Schmidt property of twisted off-diagonal elements)
Assume all conditions of Proposition 3.26.
  1. (i)
    If $D_{0}$ can be extended by continuity to a bounded operator, again denoted $D_{0}\in\mathcal{B}(\mathfrak{h})$, then there is $\delta\in\mathbb{R}^{+}$ such that $(\varkappa_{t}(D_{0}))_{t\in\left[-\delta,\delta\right]}\subseteq\mathcal{L}^{2}(\mathfrak{h})$ and this operator family is continuous in $\mathcal{L}^{2}(\mathfrak{h})$ at $t=0$, i.e.,
    $\lim_{t\rightarrow 0}\left\|\varkappa_{t}\left(D_{0}\right)\right\|_{2}=0\ .$
  2. (ii)
    If for some $c\in\mathbb{R}^{+}\cup\{\infty\}$,
    $\mathrm{e}^{it\Upsilon_{0}}\mathcal{Y}\subseteq\mathcal{Y}\ ,\qquad t\in\left(-c,c\right)\ ,$
    with the operator family $(\overline{D}_{0}\mathrm{e}^{it\Upsilon_{0}})_{t\in(-c,c)}$ being strongly continuous on $\mathcal{Y}$ and $\Psi\in\mathcal{D}\left(\mathrm{H}\right)$, then $D_{0}$ can be extended by continuity to a Hilbert-Schmidt operator, again denoted $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$.
Proof. Combine Proposition 3.26 with Lemma 3.18 as well as Corollaries 3.22 and 3.23.
3.7 Paradigmatic Example in Condensed Matter Physics
Hamiltonians which are quadratic in terms of annihilation and creation operators appear very often in theoretical physics. Here, we shortly present the paradigmatic example of the BCS theory of superconductivity, proposed in the late 1950s (1957) to explain conventional type I superconductors and named after Bardeen, Cooper and Schrieffer [2, 3, 4]. Indeed, this theory implies the first use of a fermionic quadratic Hamiltonian in Physics, even if it is very elementary as compared to the general quadratic Hamiltonians studied here.
An important model (of the lattice version) of the BCS theory is given by the so-called (reduced) BCS Hamiltonian defined as follows: In solid-state physics, quantum systems are usually done on crystalline structures. Having in mind a traditional view of crystals as objects with a regular spatial order, we consider fermions inside a cubic box $\Lambda_{L}\doteq\{\mathbb{Z}\cap\left[-L,L\right]\}^{d}$ of volume $|\Lambda_{L}|$ for $L\in\mathbb{N}$, where $d\in\mathbb{N}$. It means that the separable one-particle Hilbert space is $\mathfrak{h}=\ell^{2}(\Lambda_{L}\times\mathrm{S})$, where $\mathrm{S}$ is some finite set of spins that is fixed once and for all by $\mathrm{S}\doteq\left\{\downarrow,\uparrow\right\}$. The orthonormal basis of this Hilbert space we use is defined by the functions
$\varphi_{\left(k,\mathrm{s}\right)}\left(x,\mathrm{t}\right)\doteq\frac{1}{\left|\Lambda_{L}\right|^{1/2}}\mathrm{e}^{-ik\cdot x}\delta_{\mathrm{s},\mathrm{t}}\ ,\qquad x\in\Lambda_{L},\ \mathrm{t}\in\mathrm{S}\ ,$ (184)
for any $k\in\Lambda_{L}^{\ast}$ and $\mathrm{s}\in\mathrm{S}$, where
$\Lambda_{L}^{\ast}\doteq\frac{2\pi}{(2L+1)}\Lambda_{L}\subseteq\left[-\pi,\pi\right]^{d}$
is the reciprocal lattice of quasi-momenta (periodic boundary conditions) associated with $\Lambda_{L}$. The orthonormal basis taken in this example is not real. Reality of orthonormal bases is a condition used in Section 2. This assumption is however not necessary and is only used for convenience, in order to slightly simplify our computations by avoiding various conjugate functions in our expressions. See Equation (56). One could anyway use the real and imaginary parts of each function (184) to define the real orthonormal basis, and thus obtain a new form of the Hamiltonian presented below. But we stick to the complex basis to perform the computations since it is the natural (particularly in Physics) and original framework to describe the BCS theory.
The reduced BCS Hamiltonian is a self-adjoint operator acting on the fermionic Fock space $\mathcal{F}$ (3) and defined by
$\mathrm{H}_{L}^{BCS}\doteq\sum\limits_{k\in\Lambda_{L}^{\ast},\ \mathrm{s}\in\mathrm{S}}\left(\varepsilon_{k}-\kappa\right)a_{(k,\mathrm{s})}^{\ast}a_{(k,\mathrm{s})}-\frac{1}{\left|\Lambda_{L}\right|}\sum_{k,q\in\Lambda_{L}^{\ast}}\gamma_{k,q}a_{(k,\uparrow)}^{\ast}a_{(-k,\downarrow)}^{\ast}a_{(-q,\downarrow)}a_{(q,\uparrow)}\ ,$ (185)
where $\gamma_{k,q}$ is a positive function and $\kappa\in\mathbb{R}$ is the chemical potential, while $\varepsilon_{k}=\varepsilon_{-k}\in\mathbb{R}_{0}^{+}$ is the kinetic energy [^28] In the lattice $\mathbb{Z}^{d}$, usually, $\varepsilon_{k}=2\left(d-\left(\cos(k_{1})+\cdots+\cos(k_{d})\right)\right)$ is the one-particle energy spectrum in the quasi-momenta $k=(k_{1},\ldots,k_{d})\in\Lambda_{L}^{\ast}$ of free fermions. of a fermion in the quasi-momenta $k\in\Lambda_{L}^{\ast}$. In physics, one usually takes
$\gamma_{k,q}\doteq\left\{\begin{array}[]{l}\gamma\geq 0\\ 0\end{array}\begin{array}[]{l}\mathrm{for\ }\left|k-q\right|\leq\mathrm{C}\\ \mathrm{for\ }\left|k-q\right|>\mathrm{C}\end{array}\right.$
with $\mathrm{C}\in\left(0,\infty\right]$. The simple choice $\mathrm{C}=\infty$, i.e., $\gamma_{k,q}=\gamma>0$ in (185), is still physically very interesting since, even when $\varepsilon_{k}=0$, the BCS Hamiltonian qualitatively displays most of basic properties of real conventional type I superconductors. See, e.g. [53, Chapter VII, Section 4] or [54] for a more recent mathematically rigorous study.
In fact, the thermodynamic and dynamical behaviors of the quantum system described by the (mean-field) model (185) for $\gamma_{k,q}=\gamma>0$ can be rigorously [^29] See [26, Section 6.6.2] (static case) and [55] (dynamical case) for a first overview or [56, 57, 58] for the full theory of mean-field models for lattice fermions and quantum spin systems. obtained in infinite volume $L\rightarrow\infty$ by using the approximating model
$\mathrm{H}_{L}^{BCS}(c)=\sum\limits_{k\in\Lambda_{L}^{\ast},\ \mathrm{s}\in\mathrm{S}}\left(\varepsilon_{k}-\kappa\right)a_{(k,\mathrm{s})}^{\ast}a_{(k,\mathrm{s})}-\gamma\sum_{k\in\Lambda_{L}^{\ast}}\left(ca_{(k,\uparrow)}^{\ast}a_{(-k,\downarrow)}^{\ast}+\overline{c}a_{(k,\downarrow)}a_{(-k,\uparrow)}\right)+\gamma\left|\Lambda_{L}\right|\left|c\right|^{2}\mathbf{1}$ (186)
for any complex number $c\in\mathbb{C}$. Note that the above Hamiltonian can be written in the real orthonormal basis made of the real and imaginary parts of $\varphi_{\left(k,\mathrm{s}\right)}$ for $k\in\Lambda_{L}^{\ast}$ and $\mathrm{s}\in\{\uparrow,\downarrow\}$, since the mapping $\varphi\mapsto a\left(\varphi\right)$ is antilinear. It is a quadratic model in terms of annihilation and creation operators that can be diagonalized and written as
$H_{L}^{\mathrm{BCS}}(c)=\sum_{k\in\Lambda_{L}^{\ast},\ \mathrm{s}\in\mathrm{S}}\sqrt{\left(\varepsilon_{k}-\kappa\right)^{2}+\gamma^{2}|c|^{2}}\ b_{(k,\mathrm{s})}^{\ast}b_{(k,\mathrm{s})}+E\mathbf{1}$ (187)
for $\0c\in\mathbb{C}\backslash\{0\}$, where the ground state energy $E$ is given by
$E\doteq\gamma|\Lambda_{L}||c|^{2}+\sum_{k\in\Lambda_{L}^{\ast}}\left(\varepsilon_{k}-\kappa-\sqrt{\left(\varepsilon_{k}-\kappa\right)^{2}+\gamma^{2}|c|^{2}}\right)\ .$ (188)
This is done via a Bogoliubov $\ast$–automorphism uniquely defined by the algebraic relation
$\begin{cases}b_{(k,\uparrow)}=u_{k}a_{(k,\uparrow)}-v_{k}a_{(-k,\downarrow)}^{\ast}\ ,\\ b_{(-k,\downarrow)}^{\ast}=v_{k}^{\ast}a_{(k,\uparrow)}+u_{k}^{\ast}a_{(-k,\downarrow)}^{\ast}\ ,\end{cases}$ (189)
where $u_{k}$ and $v_{k}$ are the complex numbers
$u_{k}=\frac{\mathrm{e}^{i\arg c}}{\sqrt{2}}\sqrt{1+\frac{\varepsilon_{k}-\kappa}{\sqrt{\left(\varepsilon_{k}-\kappa\right)^{2}+\gamma^{2}|c|^{2}}}}\quad\text{and}\quad v_{k}=\frac{\mathrm{e}^{i\arg c}}{\sqrt{2}}\sqrt{1-\frac{\varepsilon_{k}-\kappa}{\sqrt{\left(\varepsilon_{k}-\kappa\right)^{2}+\gamma^{2}|c|^{2}}}}$
for each $k\in\Lambda_{L}^{\ast}$ and $\0c\in\mathbb{C}\backslash\{0\}$. To get (187), it suffices to inverse the system (189) and substitute the corresponding equalities into (186). Such kind of algebraic manipulation was originally invented for bosons by Bogoliubov [1] in 1947 and it is a key ingredient in the Bogoliubov theory of superfluidity for Helium 4. For fermions, such quadratic Hamiltonians are at the origin of the BCS theory of superconductivity done in 1957 [2, 3, 4] and their diagonalization has been performed in 1958 [5, 59] via a fermionic version of Bogoliubov’s transformation. In modern words, the above algebraic relations are nothing else than the implementation of a Bogoliubov transformation, as given by Theorem 2.8. Compare for instance (189) with Theorem 2.8.
In our framework, it refers to applying Theorem 2.4 and Equation (16) for $E_{0}=\gamma|\Lambda_{L}||c|^{2}$ and operators $\Upsilon_{0}$,$D_{0}$ acting on the $4$-dimensional Hilbert space with orthonormal basis
$\{\varphi_{(-k,\downarrow)},\varphi_{(-k,\uparrow)},\varphi_{(k,\downarrow)},\varphi_{(k,\uparrow)}\}\ ,$
in each fiber indexed by the set $\left\{k,-k\right\}$ for $k\in\Lambda_{L}^{\ast}$. In this orthonormal basis, $\Upsilon_{0}$ and $D_{0}$ take the following matrix forms:
$\Upsilon_{0}=\left(\begin{array}[]{cccc}\varepsilon_{k}-\kappa&0&0&0\\ 0&\varepsilon_{k}-\kappa&0&0\\ 0&0&\varepsilon_{k}-\kappa&0\\ 0&0&0&\varepsilon_{k}-\kappa\end{array}\right)\quad\text{and}\quad D_{0}=\frac{1}{2}\left(\begin{array}[]{cccc}0&0&0&\gamma c\\ 0&0&\gamma c&0\\ 0&-\gamma c&0&0\\ -\gamma c&0&0&0\end{array}\right)$ (190)
for $\varepsilon_{k}\in\mathbb{R}_{0}^{+}$, $\kappa\in\mathbb{R}$, $\gamma\in\mathbb{R}^{+}$ and $c\in\mathbb{C}$. See (4) and (186). Note in particular that $D_{0}=-D_{0}^{\top}$, $\Upsilon_{0}=\Upsilon_{0}^{\top}=\Upsilon_{0}^{\ast}$ and $\Upsilon_{0}D_{0}=D_{0}\Upsilon_{0}^{\top}$. Additionally, explicit computations show that (13) is equivalent in this elementary case to find $\mu\in\mathbb{R}^{+}$ such that
$\kappa-\varepsilon_{k}<\mu\leq\sqrt{\left(\varepsilon_{k}-\kappa\right)^{2}+\left|c\right|^{2}\gamma^{2}}\ .$
Avoiding the trivial case $c=0$ for which $D_{0}=0$, such $\mu$ always exists for $\varepsilon_{k}\in\mathbb{R}_{0}^{+}$, $\kappa\in\mathbb{R}$, $\gamma\in\mathbb{R}^{+}$ and $\0c\in\mathbb{C}\backslash\{0\}$. As a consequence, one can satisfy all the assumptions of Theorem 2.4 in this case and use Equations (16) and (17) to deduce in each fiber $\left\{k,-k\right\}$ that
[…]
as well as
$-8\int_{0}^{\infty}\left\|D_{\tau}\right\|_{2}^{2}\mathrm{d}\tau=\frac{1}{2}\mathrm{tr}\left(\Upsilon_{0}-\sqrt{\Upsilon_{0}^{2}+4D_{0}D_{0}^{\ast}}\right)=2\left(\varepsilon_{k}-\kappa-\sqrt{\left(\varepsilon_{k}-\kappa\right)^{2}+\gamma^{2}|c|^{2}}\right)$
for any $\varepsilon_{k}\in\mathbb{R}_{0}^{+}$, $\kappa\in\mathbb{R}$, $\gamma\in\mathbb{R}^{+}$ and $\0c\in\mathbb{C}\backslash\{0\}$. To recover the results (187)–(188) it then suffices to sum up over the set $\left\{k,-k\right\}_{k\in\Lambda_{L}^{\ast}}$, keeping in mind the explicit form of the Hamiltonian H in Theorem 2.4.
Last but not least, observe that the asssumptions of Berezin’s theorem (Theorem 2.3) do not always hold true even in this elementary situation because the condition $\Upsilon_{0}\geq\alpha\mathbf{1}$ for a strictly positive $\alpha\in\mathbb{R}^{+}$ in Theorem 2.3 is broken as soon as $\kappa\geq\varepsilon_{k}$.
Appendix A Appendix: The Elliptic Operator-Valued Flow
In this appendix, we gather only the important results of [24] associated with its application to the $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians.
Recall first that the Banach space of bounded operators acting on $\mathfrak{h}$ is denoted by $\mathcal{B}(\mathfrak{h})$, while $\mathcal{L}^{1}(\mathfrak{h})$ and $\mathcal{L}^{2}(\mathfrak{h})$ are the spaces of trace-class and Hilbert-Schmidt operators, respectively. In the paper [24] we study for lower semibounded self-adjoint operators $\Upsilon_{0}$ and (non-zero) bounded operators $D_{0}\in\mathcal{B}(\mathfrak{h})$ the non-linear system of differential equations
$\left\{\begin{array}[]{llll}\partial_{t}\Delta_{t}\varphi=16D_{t}D_{t}^{\ast}\varphi\ ,&\Delta_{t=0}\doteq 0\ ,&t\in\mathbb{R}_{0}^{+}\ ,&\varphi\in\mathfrak{h}\ ,\\ \partial_{t}D_{t}\varphi=-2\left(\left(\Delta_{t}+\Upsilon_{0}\right)D_{t}+D_{t}\left(\Delta_{t}+\Upsilon_{0}\right)^{\top}\right)\varphi\ ,&D_{t=0}\doteq D_{0}\ ,&t\in\mathbb{R}^{+}\ ,&\varphi\in\mathcal{D}(\Upsilon_{0}^{\top})\ ,\end{array}\right.$ (191)
either on the Banach space $\mathcal{B}(\mathfrak{h})^{2}$ or in $\mathcal{L}^{1}(\mathfrak{h})\times\mathcal{L}^{2}(\mathfrak{h})$. Here, $\mathcal{D}(\Upsilon_{0}^{\top})\subseteq\mathfrak{h}$ is the (dense) domain of the self-adjoint operator $\Upsilon_{0}^{\top}$. For the application to the $\mathrm{N}$–diagonalization of quadratic Hamiltonians, we only need the results of [24] in $\mathcal{L}^{2}(\mathfrak{h})$. In this case, the existence of a global solution to the non-linear system of differential equations (191) is given by [24, Theorems 1–2], which is reproduced below:
Theorem A.1 (Well-posedness of the flow – Hilbert-Schmidt topology)
Assume $\Upsilon_{0}=\Upsilon_{0}^{\ast}\geq-\mu\mathbf{1}$ with $\mu\in\mathbb{R}$ and $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ ($D_{0}\neq 0$). Then, there exists a unique solution $(\Delta,D)\in C(\mathbb{R}_{0}^{+};\mathcal{L}^{1}(\mathfrak{h})\times\mathcal{L}^{2}(\mathfrak{h}))$ to
$\left\{\begin{array}[]{llllll}\partial_{t}\Delta_{t}=16D_{t}D_{t}^{\ast}&,&\quad\Delta_{t=0}\doteq 0&,&\quad t\in\mathbb{R}_{0}^{+}&,\\ \partial_{t}D_{t}=-2\left(\left(\Delta_{t}+\Upsilon_{0}\right)D_{t}+D_{t}\left(\Delta_{t}+\Upsilon_{0}\right)^{\top}\right)&,&\quad D_{t=0}\doteq D_{0}&,&\quad t\in\mathbb{R}^{+}&,\end{array}\right.$
in $\mathcal{L}^{1}(\mathfrak{h})\times\mathcal{L}^{2}(\mathfrak{h})$, i.e.,in the trace class topology for $\Delta$, and in the Hilbert-Schmidt topology for $D$. Additionally, if $D_{0}^{\top}=\pm D_{0}$ then $D_{t}^{\top}=\pm D_{t}$ for any $t\in\mathbb{R}_{0}^{+}$.
The assumption $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ can be replaced with $D_{0}\in\mathcal{L}^{2p}(\mathfrak{h})$ for some $p\in[1,\infty]$, where $\mathcal{L}^{q}(\mathfrak{h})$ is the $L^{q}$–spaces constructed from Schatten norms. In this case, the operator-valued flow is a well-posed system of differential equations on $\mathcal{L}^{p}\left(\mathfrak{h}\right)\times\mathcal{L}^{2p}\left(\mathfrak{h}\right)$, at least when $\Upsilon_{0}\in\mathcal{B}(\mathfrak{h})$. See [24, Theorem 14]. However, the case $p=1$ is the only relevant one in view of the application of the flow given here.
The convergence of the flow is a difficult task also solved in [24] under sufficient, albeit still general, conditions. This is done from geometrical characteristics of the flow. We refer here to its ellipticity properties which can first be seen via its constants of motion: For any self-adjoint operator $\Upsilon_{0}$ that is bounded and every non-zero Hilbert-Schmidt operator $D_{0}=\pm D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$, one checks that
$\mathrm{tr}\left(\Upsilon_{t}^{2}+4D_{t}D_{t}^{\ast}-\Upsilon_{s}^{2}-4D_{s}D_{s}^{\ast}\right)=0\ ,\qquad s,t\in\mathbb{R}_{0}^{+}\ ,$ (192)
where $\Upsilon_{t}\doteq\Upsilon_{0}+\Delta_{t}$ for any $t\in\mathbb{R}_{0}^{+}$, while $\mathrm{tr}(\cdot)$ denotes the usual trace for operators acting the separable Hilbert space $\mathfrak{h}$. See [24, Equation (23)]. Moreover, if $\Upsilon_{0}D_{0}=D_{0}\Upsilon_{0}^{\top}$ on the domain $\mathcal{D}(\Upsilon_{0}^{\top})\subseteq\mathfrak{h}$ of $\Upsilon_{0}^{\top}$ together with $\Upsilon_{0}\geq-\left(\mu-\varepsilon\right)\mathbf{1}$ for some $\mu\in\mathbb{R}$ and $\varepsilon\in\mathbb{R}^{+}$ and
$\max\left\{\|(\Upsilon_{0}+\mu\mathbf{1)}D_{0}(\Upsilon_{0}^{\top}+\mu\mathbf{1)}^{-1}\|_{\mathrm{op}},\|(\Upsilon_{0}^{\top}+\mu\mathbf{1)}D_{0}^{\ast}(\Upsilon_{0}+\mu\mathbf{1)}^{-1}\|_{\mathrm{op}}\right\}\leq\mathrm{C}$ (193)
for some strictly positive constant $\mathrm{C}\in\mathbb{R}^{+}$, then one verifies that, for any $t,s\in\mathbb{R}_{0}^{+}$,
$\Upsilon_{t}D_{t}=D_{t}\Upsilon_{t}^{\top}\qquad\text{and}\qquad\Upsilon_{t}^{2}+4D_{t}D_{t}^{\ast}=\Upsilon_{s}^{2}+4D_{s}D_{s}^{\ast}$ (194)
on the domains $\mathcal{D}(\Upsilon_{0}^{\top})=\mathcal{D}(\Upsilon_{t}^{\top})$ and $\mathcal{D}(\Upsilon_{0}^{2})$, respectively. See [24, Equation (25)]. The last constant of motion can be viewed as the equation of an ellipse in the “commutative” situation.
In the general cases, the natural elipticity properties of the operator-valued flow are written in terms of the operators
$\mathfrak{D}_{t}\doteq\Upsilon_{t}-\mu\mathbf{1}+4D_{t}\left(\Upsilon_{t}^{\top}+\mu\mathbf{1}\right)^{-1}D_{t}^{\ast}\ ,\qquad t\in\mathbb{R}_{0}^{+}\ ,$ (195)
with $\mu\in\mathbb{R}$ such that [^30] Note that $\Upsilon_{0}\geq-\left(\mu-\varepsilon\right)\mathbf{1}$ iff $\Upsilon_{0}^{\top}\geq-\left(\mu-\varepsilon\right)\mathbf{1}$. $\Upsilon_{0}\geq-\left(\mu-\varepsilon\right)\mathbf{1}$ for some $\varepsilon\in\mathbb{R}^{+}$. For instance, in [24, Theorem 3] one proves that the real-valued function
$\zeta\left(t\right)=\inf\sigma\left(\mathfrak{D}_{t}\right)\in\left[-2\mu,\infty\right)\ ,\qquad t\in\mathbb{R}_{0}^{+}\ ,$
$\sigma(\mathfrak{D}_{t})$ being the spectrum of $\mathfrak{D}_{t}$, is upper semicontinuous [^31] For bounded operators $\Upsilon_{0}\in\mathcal{B}\left(\mathfrak{h}\right)$, $\zeta$ is always continuous on $\mathbb{R}_{0}^{+}$. at $t=0$ and continuous on $\mathbb{R}^{+}$ as well as decreasing with $\zeta\left(\mathbb{R}_{0}^{+}\right)\subseteq[0,\zeta\left(0\right)]$ whenever $\zeta\left(0\right)\in\mathbb{R}_{0}^{+}$, while it is continuously increasing with $\zeta\left(\mathbb{R}_{0}^{+}\right)\subseteq[\zeta\left(0\right),0)$ whenever $\zeta\left(0\right)\in\mathbb{R}^{-}$.
Properties of the operator family $(\mathfrak{D}_{t})_{t\in\mathbb{R}_{0}^{+}}$ proven in [24] are important to obtain a good convergence of the operator-valued flow, with possibly exponential decay. This refers to [24, Theorem 4], which corresponds to the following theorem in the Hilbert-Schmidt case:
Theorem A.2 (Asymptotics of the operator-valued flow)
Take $D_{0}=\pm D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ and $\Upsilon_{0}=\Upsilon_{0}^{\ast}$, both acting on $\mathfrak{h}$. Assume that
$\Upsilon_{0}\geq-\left(\mu-\varepsilon\right)\mathbf{1}\qquad\text{and}\qquad\mathfrak{D}_{0}\doteq\Upsilon_{0}+4D_{0}\left(\Upsilon_{0}^{\top}+\mu\mathbf{1}\right)^{-1}D_{0}^{\ast}\geq\mu\mathbf{1}$
for some $\0\mu\in\mathbb{R}\backslash\{0\}$ and $\varepsilon\in\mathbb{R}^{+}$. Then, as $t\rightarrow\infty$, $D$ exponentially converges in the Hilbert-Schmidt topology (i.e., in $\mathcal{L}^{2}(\mathfrak{h})$) to zero, while $\Delta$ exponentially converges in the trace norm topology (i.e., in $\mathcal{L}^{1}(\mathfrak{h})$) to an operator $\Delta_{\infty}\in\mathcal{L}^{1}(\mathfrak{h})$ satisfying
$\Upsilon_{\infty}\doteq\Upsilon_{0}+\Delta_{\infty}\geq\left|\mu\right|\mathbf{1}\ .$
Compare the assumptions of this theorem with the ones of Theorem 2.4.
Note that [24, Theorem 4] gives more general results beyond the Hilbert-Schmidt topology. For more details as well as additional discussions on the above assumptions, see [24, Section 2.3].
Acknowledgments: This work is supported by the Basque Government through the grant IT1615-22 and the BERC 2022-2025 program, by the grant PID2020-112948GB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. We thank Walter de Siquiera Pedra and Sascha Lill for valuable comments.